### v2

# Creativity in AI: Progresses and Challenges

Abstract.
Creativity is the ability to produce novel, useful, and surprising ideas, and has been widely studied as a crucial aspect of human cognition. Machine creativity on the other hand has been a long-standing challenge. With the rise of advanced generative AI, there has been renewed interest and debate regarding AI’s creative capabilities. Therefore, it is imperative to revisit the state of creativity in AI and identify key progresses and remaining challenges. In this work, we survey leading works studying the creative capabilities of AI systems, focusing on creative problem-solving, linguistic, artistic, and scientific creativity. Our review suggests that while the latest AI models are largely capable of producing linguistically and artistically creative outputs such as poems, images, and musical pieces, they struggle with tasks that require creative problem-solving, abstract thinking and compositionality and their generations suffer from a lack of diversity, originality, long-range incoherence and hallucinations. We also discuss key questions concerning copyright and authorship issues with generative models. Furthermore, we highlight the need for a comprehensive evaluation of creativity that is process-driven and considers several dimensions of creativity. Finally, we propose future research directions to improve the creativity of AI outputs, drawing inspiration from cognitive science and psychology.
1. Introduction
Computers can’t create anything. For creation requires, minimally, originating something. But computers originate nothing; they merely do that which we order them, via programs, to do.
Ada Lovelace
Creativity, the ability to produce novel, useful, and surprising ideas (Boden, 2004), is one of the major hallmarks of human intelligence (Guilford, 1967). Since the invention of the first known general-purpose mechanical computer (known as Analytical Engine) designed by Babbage (Babbage, 1837), the question of whether machines can truly think or create anything new has intrigued the scientific community (Wang et al., 2024b; Turing, 1950; Newell et al., 1959). Ada Lovelace, recognized as the first programmer by many, famously stated that the Analytical Engine has no pretensions to originate anything (Lovelace, 1843) and Alan Turing, who laid the foundations of computer science, asserted that machines can never take us by surprise (Turing, 1950). Nevertheless, alongside the development of personal computers and advancements in Artificial Intelligence (AI), several symbolic-based and stochastic approaches were developed to endow machines with story generation (Meehan, 1977; Lebowitz, 1983; y Pérez and Sharples, 2001; Turner, 1994), poetry writing (Masterman, 1971; Racter, 1984) and music composition skills (Brooks et al., 1957; Hiller and Isaacson, 1958). However, these early approaches could not generalize beyond a set of limited domains (Ji et al., 2020; Yao et al., 2019; Colton et al., 2012).
Fast-forward to now, the advent of the Transformer architecture (Vaswani et al., 2017) and the development of large language models (LLMs) (Zhao et al., 2023c) in the past decade ushered a new age of intelligent systems with remarkable generative, reasoning, coding, mathematical and multimodal capabilities (Gemini, 2024; Bubeck et al., 2023; Wei et al., 2022). Transformer-based models can now produce long stories in various domains (Yang et al., 2022; Yao et al., 2019), write poems about diverse topics (Chakrabarty et al., 2022a; Ormazabal et al., 2022a), compose high-fidelity songs (Dhariwal et al., 2020), generate impressive high-quality images and videos (Brooks et al., 2024; Betker et al., 2020) and discover new scientific knowledge (Jumper et al., 2021).
While these remarkable achievements can be seen as signs of the presence of creative capacity in transformer-based language models, it should be noted that these models rely on an astronomically large number of parameters and are trained on massive amounts of public and private data (Brown et al., 2020a). Hence, it is not entirely clear whether the seemingly extraordinary outputs of these models are the result of a truly creative inner process and robust generalization or the result of powerful interpolation and strong memorization skills (McCoy et al., 2023; Hupkes et al., 2022; Carlini et al., 2022; Bender et al., 2021; Marcus, 2020; Bender and Koller, 2020). Recent works have shown that language models fail at real-world commonsense reasoning and compositionality tasks (Ismayilzada et al., 2023; Dziri et al., 2023), occasionally copy large amounts of text from their training data (McCoy et al., 2021), significantly lag behind experts in creative writing (Chakrabarty et al., 2023a), produce less diverse content (Anderson et al., 2024a; Padmakumar and He, 2023), struggle with creative problem-solving (Tian et al., 2023) and abstract reasoning (Gendron et al., 2023; Mitchell et al., 2023) and suffer from factual inconsistency and hallucination issues (Banerjee et al., 2024; Elazar et al., 2021). While previous works have reviewed developments on some aspects of AI creativity (Amin and Burghardt, 2020; Lai and Nissim, 2024; Oliveira, 2017; Elzohbi and Zhao, 2023; Franceschelli and Musolesi, 2021b; Rowe and Partridge, 1993), a more holistic and broader review of the field is necessary to understand its rapid advancements.
In this work, we provide the general AI audience with a timely summary of the state of the creative capabilities of the latest AI systems. While creativity is a broad concept that can be explored in a wide range of areas, in this survey, we focus on four main areas where machine creativity has been extensively investigated: Linguistic Creativity (§3), Creative Problem-Solving (§4), Artistic Creativity (§5) and Scientific Creativity (§6). These areas capture four crucial pillars of creative thinking in humans: linguistic creativity enables us to manipulate language in novel ways for effective communication of ideas; creative problem-solving helps us find efficient solutions by thinking out-of-the-box; artistic creativity allows for the expression of emotions, ideas, and aesthetics through various media; and scientific creativity drives innovation by enabling the formulation of new hypotheses, theories, and discoveries. Together, these areas represent distinct yet interconnected facets of creativity, providing a comprehensive framework for studying how machines can emulate or assist human-like creative processes across different domains. For each area, we survey representative tasks, resources, methods, and major findings and present a taxonomy of these works in Figure 2. Our review indicates that although the latest AI models are generally proficient in generating linguistically and artistically creative outputs, such as poems, images, and music, they face challenges with tasks demanding creative problem-solving, abstract reasoning, and compositionality. Their outputs often lack diversity and originality, exhibit long-term incoherence, and are prone to hallucinations. We also briefly discuss the emerging challenges brought by generative models concerning copyright and authorship of artworks (§7). Finally, in the last section (§8), we argue for a comprehensive evaluation of creativity in AI that considers several dimensions of creativity and the creative process at its core. Furthermore, we discuss future research directions to enhance the creativity of AI systems, potentially drawing ideas from cognitive science and psychology.
Our goal in this survey is to provide a high-level yet comprehensive overview of the state of creativity in AI. We expect our survey to provide researchers working on machine creativity with comprehensive background knowledge and encourage them to explore new avenues for developing intelligent systems that can do creative generation.
2. Creativity
2.1. Definitions
While creativity as a concept seems intuitively easy to understand on the surface, there is still no consensus on what constitutes true creativity. This is primarily due to the subjective nature of creativity, as what is deemed novel and of quality can vary significantly across cultures, disciplines, and periods. Aleinikov et al. (2000) lists more than proposed definitions, and the number keeps growing. Despite the lack of global consensus, there is one definition of creativity that has seen wide adoption by many philosophers and psychologists and has been dubbed as the “standard definition” (Runco and Jaeger, 2012; Barron, 1955; Stein, 1953). According to this definition, creativity requires novelty (a.k.a originality, uniqueness, etc.) and value (a.k.a utility, effectiveness, usefulness, appropriateness, relevance, meaningfulness, etc.).
The novelty criterion is typically self-explanatory to the point that people equate it to creativity in everyday life. This is generally known as intrapersonal or personal creativity (often denoted as P-creativity), i.e. the product is novel within the frame of a person’s life (Boden, 2004; Stein, 1953; Weisberg, 1986). Researchers distinguish it from the interpersonal or historical creativity (often denoted as H-creativity), i.e. the product is novel with respect to the entire history of people such as Einstein’s general relativity theory.
Many theorists have argued that novelty is insufficient for creativity, and value dimension is needed to filter out original nonsense, such as something generated by a truly random process. While value is generally understood as something inherently “good” for the respective audience, there appears to be such a thing as malevolent or “dark” creativity. For instance, one can be creative in producing torture instruments or in committing terrorist atrocities (Gaut, 2010). Therefore, the interpretation of the value of a product as being “effective” towards its intended end, regardless of whether that end is morally good or bad, has been suggested as a better alternative (Livingston, 2018). However, we should note that evaluating utility or value still requires an outside judgment which is subjective, can be faulty or biased and can change over time and across cultures. This is especially apparent in arts as there are many great artists (Bach, Van Gogh etc.) whose “value” have only been recognized longtime after their death. Moreover, sometimes novelty is itself the value created by the artist because no one has done it before, particularly, in visual arts. Hence, some researchers have recently argued to drop the value criterion altogether from the definition of creativity (Brandt, 2021; Weisberg, 2015b).
Despite its wide adoption, the sufficiency of the novelty and value conditions for creativity has also been challenged (Gaut, 2010; Arnheim, 2001; Weisberg, 2015a). It has been argued that agency, a capacity to have beliefs, desires or intentional states, is a required attribute of creativity. For example, Gaut (2010) mentions the tectonic movement of the earth’s crust that can produce valuable (financially and aesthetically) and sometimes original (new variation) diamonds, but we would hardly call tectonic movements creative. However, mere agency without intentionality is also insufficient. Gaut (2010) illustrates this with an example where a person walking in a studio accidentally knocks over a set of paints, which spill onto a canvas and happen to create a beautiful and original painting. Weisberg (2015b) has gone even further to suggest that creativity is simply intentional novelty.
While the creative process should be intentionally initiated, others have argued that the creative process should involve an element of spontaneity (Kronfeldner, 2009). This allows the creative product to induce a surprise in the audience since the output of the process is not foreseen from the beginning. Being ignorant of the end at the outset of the creative process opens the room for creativity as opposed to a mechanical routine or algorithm that, by definition, is exact and excludes any type of spontaneous modifications. To illustrate this contrast Perkins (2001) makes a distinction between reasonable problems (i.e. that can be reasoned out step-by-step such as anagrams) and unreasonable problems (i.e. that are hard to describe with a step-by-step thinking).
The element of surprise has been further developed by (Boden, 2004) into a widely recognized third dimension of the “standard definition” of creativity. This new definition can be seen as an elaborated version of the three criteria (i.e. new, useful and non-obvious) used by the United States Patent Office to determine whether an invention can come under patent protection111http://www.uspto.gov/inventors/patents.jsp (Simonton, 2012). In this survey, we will also take this extended definition as our working definition throughout the paper.
2.2. Types
While creativity manifests itself in various forms across domains, even within a particular domain, different types of creativity can be distinguished based on the difficulty level of the inherent creative process involved. Four-C model of creativity is one such popular theory that differentiates between four types of creativity corresponding to four levels of difficulty involved in producing creative artifacts (Kaufman and Beghetto, 2009). The first major type of creativity is known as little-c creativity which is what we find in everyday life as solutions to minor problems. Examples might include combining unusual ingredients to make a new type of meal or using a hand-held vacuum cleaner on the ceiling to remove flies. Almost everyone possesses this type of creativity in one way or another. The second main type of creativity in this model is the Big-C creativity that includes major works of scientific, technological, social, or artistic importance. Examples could be Darwin’s theory of evolution, the invention of the printing press, or Leonardo Da Vinci’s painting of the Mona Lisa. In addition to these two major categories, Kaufman and Beghetto (2009) also defines two minor categories of creativity. First is the mini-c creativity for very small-scale cases of creativity such as young children’s drawing or their creative experiments with Lego pieces. Second is the Pro-C creativity which is proposed for work produced by professional but non-prominent practitioners such as professional musicians or artists who generate novel work, but do not make historical contributions.
Recently, there has been a suggestion to differentiate three types of creativity corresponding to three major levels of innovation that can be achieved (Hassabis, 2018). First can be achieved through interpolation where a prototypical creative artifact is produced by averaging all the artifacts of the same class seen before. While it is an original product that did not exist before, it still relies heavily on the other existing products. An example would be to come up with a novel winning strategy in chess that is a combination of existing different strategies. Consequently, a second type of innovation can be achieved through extrapolation where a creative artifact extends the boundaries of what has been seen before, but is still of the same class. A completely new chess move that is not related to any existing moves can be seen as an example of extrapolative creativity. Finally, the highest level of creativity can be termed as invention where a creative artifact introduces a novel class of its own. Inventing chess itself or any major scientific invention is a perfect example of this type of creativity.
2.3. Evaluation
Evaluating creativity remains a challenging task in artificial intelligence due to its inherently subjective nature (Lamb et al., 2018). Interestingly, some research work even argued against the quantitative evaluation of creativity, suggesting it is either too domain-specific to be measured effectively (Baer, 2012), or that creativity is an inherently human trait that cannot be accurately modeled computationally (Boden, 1991; Minsky, 1982). However, an overwhelming majority of the scientific community favors the possibility of computational modeling and evaluation of creativity (Veale and Cardoso, 2019). Hence, numerous evaluation methods have been proposed in the past (Lamb et al., 2018). However, most of the proposed metrics are either formal frameworks that are hard to implement in practice or manual psychometric creativity tests that require costly human involvement (Kim, 2006) or automated metrics that are too domain-specific (França et al., 2016). We refer the reader to Franceschelli and Musolesi (2021b) and Lamb et al. (2018) for more details on formal evaluation frameworks, and here we briefly summarize some of the relevant manual and automated metrics for creativity.
2.3.1. Manual Evaluation
Since creative products vary greatly in their forms and are hard to characterize with objective measures, the simplest and most common way to evaluate them is to ask other humans to manually rate them based on some criteria associated with creativity, which differs from task to task (Lamb et al., 2015). For example, in story generation, humans are typically asked to rate a generated story on aspects such as interestingness, coherence, relevance, humanlikeness and etc. (Goldfarb-Tarrant et al., 2020; Rashkin et al., 2020; Yang et al., 2022). In other tasks where the goal is to produce multiple responses such as common psychometric creativity tests Alternative Uses Task (AUT) (Guilford, 1967) and Torrance Tests of Creative Thinking (TTCT) (Torrance, 1974), evaluation is centered around four dimensions of creativity: fluency (the total number of meaningful, and relevant ideas generated in response to the stimulus), flexibility (the number of different categories of relevant responses), originality (the uniqueness or rarity of responses) and elaboration (the amount of detail in the responses).
While it is common and straightforward to conduct human evaluation with ordinary humans, some have argued that people who are not experts on a kind of creative artifact might not be good judges of those artifacts (Mirowski et al., 2022; Gervás, 2019; Veale, 2015; Lamb et al., 2018, 2015). This typically results in poor interrater reliability and even when they agree, their judgments do not correlate well with expert judgment (Lamb et al., 2018). Therefore, it is generally recommended to employ Consensual Assessment Technique (Amabile, 1983), an evaluation method that relies on the collective judgment of experts in a given field.
2.3.2. Automated Evaluation
While creativity is generally evaluated by humans, several attempts have also been made to devise automated measures of it (França et al., 2016; Cook and Colton, 2015; Jordanous et al., 2015; Maher and Fisher, 2012). These measures often target a specific dimension of creativity. Below, we review some automated measures for three dimensions of creativity: novelty, value, and surprise.
Novelty
It is typically defined as the measure of how different an artifact is from other known artifacts in its class (Maher, 2010). Then a distance metric is established to quantify this difference based on the attributes of the artifact and the task space. For example, in the text generation task, a notion of semantic distance is commonly employed as a distance measure (Johnson et al., 2022; Beaty and Johnson, 2020; Prabhakaran et al., 2013; Dunbar and Forster, 2009; Harbison and Haarmann, 2014). More specifically, the text is embedded into a vector in semantic space and some distance or dissimilarity metric (e.g. typically 1-cosine_similarity) is used to compute how much semantically different is one text from another. However, the granularity of the text can differ from task to task. For example, in the story generation task, Karampiperis et al. (2014) defines the novelty of a story as the average semantic distance between the dominant terms included in the textual representation of the story, compared to the average semantic distance of the dominant terms in all stories where distance is measured based on the embeddings of terms.
Novelty can also be characterized by the degree an artifact differs from the previously produced works that one has already seen (Elgammal and Saleh, 2015; Gunkle and Berlyne, 1975). This definition has inspired the development of Creative Adversarial Networks (CANs) (Elgammal et al., 2017) similar to the popular Generative Adversarial Networks (GANs) (Goodfellow et al., 2014a). In CANs, the generator tries to fool the discriminator into thinking its generation is “art” and at the same time, the style of its generation is nothing known to the discriminator. Consequently, the score assigned by the discriminator (more specifically, 1-score) can be used to measure the novelty of the generated artifact as suggested by Franceschelli and Musolesi (2022).
Value
This dimension is generally the hardest to evaluate as it depends on the subjective utility or performance of the artifact which is typically judged by domain experts of that artifact and can radically change across domains (Maher, 2010). In visual arts, this might correspond to “beauty”, whereas in science to “logical correctness”. Therefore, a metric appropriate for its domain should be employed. For example, in open-ended story generation, a minimally useful story can be defined as a relevant, coherent, and meaningful story. In this sense, automated metrics measuring the overall quality of a story can be leveraged (Xie et al., 2023a; Guan and Huang, 2020; Chen et al., 2022b), however, it is often challenging to measure coherence (Zhao et al., 2023b; Laban et al., 2021). Another more general evaluation of utility has been suggested by Franceschelli and Musolesi (2022) based on the discriminator score in GANs.
Surprise
Also known as unexpectedness, surprise measures the artifact’s degree of deviation from what is expected (Maher, 2010). Therefore, automatic metrics for surprise tend to be information-theoretic (Kuznetsova et al., 2013; Bunescu and Uduehi, 2022) and estimate the violation of expectation based on uncertainty reduction (Frank, 2010; Hale, 2006). However, semantic distance-based measures of surprise have also been suggested. For example, in the story generation task, Karampiperis et al. (2014) conceptualizes surprise as the average semantic distances between the consecutive fragments of a given story. Recent work has also suggested an automated measure based on the Bayesian theory of surprise (Franceschelli and Musolesi, 2022; Baldi and Itti, 2010).
forked edges,
for tree=
grow=east,
reversed=true,
anchor=base west,
parent anchor=east,
child anchor=west,
base=left,
font=,
rectangle,
draw=hidden-draw,
rounded corners,
align=left,
minimum width=4em,
edge+=darkgray, line width=1pt,
s sep=10pt,
inner xsep=2pt,
inner ysep=3pt,
ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center,
where level=1text width=4.2em,font=,,
where level=2text width=6.4em,font=,,
where level=3text width=5.5em,font=,,
where level=4text width=6.1em,font=,,
Creativity in AI, ver
Linguistic
Creativity (§3)
Humor (§3.1)
Datasets
Humorous headlines (Horvitz et al., 2024; Hossain et al., 2020b, 2019),
Puns (Miller et al., 2017; Yang et al., 2015),
Jokes (Tang et al., 2022; Meaney et al., 2021; Jain et al., 2024; Zhong et al., 2023; Zhang et al., 2019b; Weller and Seppi, 2019),
(Hasan et al., 2019; Mihalcea and Strapparava, 2005)
, lc-leaf, text width=40em
Methods
Humorous content (Ravi et al., 2024; Hessel et al., 2023b; Xie et al., 2023b; Arora et al., 2022; Peyrard et al., 2021; Kayatani et al., 2021),
(Annamoradnejad and Zoghi, 2020; Ziser et al., 2020; Amin and Burghardt, 2020; Chen and Soo, 2018; Liu et al., 2018b; Bertero and Fung, 2016),
(Radev et al., 2016; Yang et al., 2015; Zhang and Liu, 2014; Stock and Strapparava, 2005; Taylor and Mazlack, 2004; Raskin and Attardo, 1994)
(Hasan et al., 2019),
Pun (Mittal et al., 2022; He et al., 2019; Yu et al., 2018),
Sarcasm (Chakrabarty et al., 2022c, 2020a),
Joke (Jentzsch and Kersting, 2023; Xie et al., 2021; Meaney et al., 2021; Weller and Seppi, 2019; Ren and Yang, 2017; Shahaf et al., 2015),
, lc-leaf, text width=42em
Figurative
Language (§3.2)
Datasets
Metaphor (Chakrabarty et al., 2023e, 2022c; Stowe et al., 2021, 2020; Mohammad et al., 2016; Mohler et al., 2016),
Simile (Chakrabarty et al., 2021a, 2022c),
Idiom (Chakrabarty et al., 2021a, 2022c)
, lc-leaf, text width=40em
Methods
Metaphor (Chakrabarty et al., 2023e, 2021b; Stowe et al., 2021, 2020),
Simile (He et al., 2023; Chakrabarty et al., 2021a, 2020b),
Idiom (Chakrabarty et al., 2021a),
Hyperbole (Tian et al., 2021)
, lc-leaf, text width=40em
Lexical Innovation
(§3.3)
Novel compound (Coil and Shwartz, 2023; Dhar and van der Plas, 2019; Vecchi et al., 2017; Kuznetsova et al., 2013; Hendrickx et al., 2013; Butnariu et al., 2009),
Neologism (Lencione et al., 2022; Das and Ghosh, 2017),
Morphological Productivity (Ismayilzada et al., 2024; Anh et al., 2024; Weissweiler et al., 2023)
, lc-leaf, text width=48em
Creative
Problem
Solving (§4)
Convergent
Thinking (§4.1)
NYT Connections (Samadarshi et al., 2024),
MacGyver (Tian et al., 2023),
The Only Connect Wall (Naeini et al., 2023),
RiddleSense (Lin et al., 2021)
, cps-leaf, text width=48em
Divergent
Thinking (§4.2)
BrainTeaser (Jiang et al., 2023),
UnCommonsense (Zhao et al., 2023a),
MacGyver (Tian et al., 2023),
DAT (Chen and Ding, 2023),
AUT (Hubert et al., 2024a; Koivisto and Grassini, 2023; Góes et al., 2023b; Stevenson et al., 2022),
TTCT (Zhao et al., 2024; Guzik et al., 2023; Góes et al., 2023b)
, cps-leaf, text width=48em
Abstraction (§4.3.1)
ConceptARC (Mitchell et al., 2023; Moskvichev et al., 2023; Odouard and Mitchell, 2022),
ARC (Opielka et al., 2024; Xu et al., 2023; Gendron et al., 2023; Mirchandani et al., 2023; Odouard and Mitchell, 2022; Xu et al., 2022; Zhang et al., 2021; Chollet, 2019),
RAVEN (Ahrabian et al., 2024; Gendron et al., 2023; Odouard and Mitchell, 2022; Zhang et al., 2019a; Santoro et al., 2018),
Kandinsky Patterns (Lorello et al., 2024; Holzinger et al., 2019)
Serial Reproduction (Kumar et al., 2024),
Bongard Problems (Bongard, 1970)
, cps-leaf, text width=48em
Analogy (§4.3.2)
Word Analogies (Marquer et al., 2022; Ushio et al., 2021; Gladkova et al., 2016; Mikolov et al., 2013),
Scientific Analogies (Yuan et al., 2023; Czinczoll et al., 2022),
Visual Analogies (Opielka et al., 2024; Bitton et al., 2022; Zhang et al., 2019a),
Story Analogies (Jiayang et al., 2023; Sourati et al., 2023; Sultan and Shahaf, 2022; Nagarajah et al., 2022),
Sentence Analogies (Wijesiriwardene et al., 2023; Zhu and de Melo, 2020),
Analogical Reasoning (Musker et al., 2024; Lewis and Mitchell, 2024; Petersen and van der Plas, 2023),
(Yasunaga et al., 2023; Hu et al., 2023; Webb et al., 2022)
, cps-leaf, text width=48em
Artistic
Creativity (§5)
Story
Generation (§5.1)
Datasets
StoryWars (Du and Chilton, 2023),
Visual WritingPrompts (Hong et al., 2023),
STORIUM (Akoury et al., 2020),
RolePlayingGuild (Louis and Sutton, 2018),
WritingPrompts (Fan et al., 2018)
, ac-leaf, text width=40em
Methods
Re3 (Yang et al., 2022),
Creativity Support (Chakrabarty et al., 2023c),
TTCW (Chakrabarty et al., 2023b),
Dramatron (Mirowski et al., 2022),
PEER (Schick et al., 2022),
WordCraft (Yuan et al., 2022),
TaleBrush (Chung et al., 2022),
PlotMachines (Rashkin et al., 2020),
Aristotelian Rescoring (Goldfarb-Tarrant et al., 2020),
Story Centaur (Swanson et al., 2021),
Plug and Play (Pascual et al., 2021; Dathathri et al., 2019),
Plan-And-Write (Yao et al., 2019),
Reward Shaping (Tambwekar et al., 2018),
Hierarchical (Fan et al., 2018),
UNIVERSE (Lebowitz, 1984),
TALE-SPIN (Meehan, 1977)
COINS (Paul and Frank, 2021),
, ac-leaf, text width=41em
Poetry (§5.2)
Datasets
Gutenberg (Jacobs, 2018),
CoPoet (Chakrabarty et al., 2022b),
Acrostic Poems (Agarwal and Kann, 2020),
UniM- and MultiM-Poem (Liu et al., 2018a)
, ac-leaf, text width=40em
Methods
Heuristic-based (Oliveira, 2012; Colton et al., 2012; Manurung et al., 2012; Manurung, 2004; Manurung et al., 2000),
(Racter, 1984; Masterman, 1971; Milic, 1970),
Neural-networks (Liu et al., 2018a; Lau et al., 2018; Ghazvininejad et al., 2016; Zhang and Lapata, 2014),
Language models (Popescu-Belis et al., 2023; Belouadi and Eger, 2022; Chakrabarty et al., 2022b; Ormazabal et al., 2022b),
(Tian and Peng, 2022; Van de Cruys, 2020; Agarwal and Kann, 2020)
, ac-leaf, text width=42em
Visual Creativity
(§5.3)
Image Editing
Style transfer (Karras et al., 2019; Abdal et al., 2019; Karras et al., 2018; Johnson et al., 2016; Dumoulin et al., 2016; Gatys et al., 2015),
Super-resolution (Ledig et al., 2016; Dong et al., 2014), Colorization (Zhang et al., 2016a), In-painting (Pathak et al., 2016),
Generic Image-to-Image (Huang et al., 2018a; Isola et al., 2016)
, ac-leaf, text width=40em
Text-to-Image
Creativity-VLA (Shen et al., 2024), InstructDiffusion (Geng et al., 2023), DALL-E (Betker et al., 2020; Ramesh et al., 2022, 2021),
Imagen (Saharia et al., 2022), Make-A-Scene (Gafni et al., 2022), StableDiffusion (Rombach et al., 2021),
GLIDE (Nichol et al., 2021), StackGAN (Zhang et al., 2016b), CreativeSynth (Huang et al., 2024a),
Dream Art (Ruiz et al., 2022; Huang et al., 2022), InstructPix2Pix (Brooks et al., 2022), StyleCLIP (Patashnik et al., 2021),
(Chakrabarty et al., 2023e; Hertz et al., 2022; Reed et al., 2016b, a; Mansimov et al., 2015; Yan et al., 2015)
, ac-leaf, text width=40em
Video Generation
Sora (Brooks et al., 2024), W.A.L.T (Gupta et al., 2023), VideoPoet (Kondratyuk et al., 2023),
VideoFusion (Luo et al., 2023),
Video Diffusion (Ho et al., 2022b), Imagen (Ho et al., 2022a),
Make-A-Video (Singer et al., 2022),
ViViT (Arnab et al., 2021), VideoGPT (Yan et al., 2021),
MoCoGAN (Tulyakov et al., 2017), (Vondrick et al., 2016) (Xing et al., 2023)
, ac-leaf, text width=40em
Musical Creativity
(§5.4)
MusicGen (Copet et al., 2023), MusicLM (Agostinelli et al., 2023), Jukebox (Dhariwal et al., 2020),
Pop Music Transformer (Huang and Yang, 2020),
MuseNet (Payne, 2019), LakhNES (Donahue et al., 2019),
Music Transformer (Huang et al., 2018b), MusicVAE (Roberts et al., 2018),
MidiNet (Yang et al., 2017),
MuseGAN (Dong et al., 2017) EMI (Cope, 1996), GenJam (Biles et al., 1994)
(Lavrenko and Pickens, 2003; Eck and Schmidhuber, 2002; Farbood and Schöner, 2001; Todd, 1989; Ames, 1987; Brooks et al., 1957; Hiller and Isaacson, 1958)
, ac-leaf, text width=48em
Scientific
Creativity (§6)
Equation discovery
Symbolic Regression
(Chen et al., 2022a; Garcon et al., 2021), AI Feynman (Udrescu et al., 2020),
(Cranmer et al., 2020; Petersen and Landajuela, 2019; Schmidt and Lipson, 2009),
(Todorovski, 1997; Koza, 1994; Dzeroski and Todorovski, 1993; Rzevski et al., 1987), BACON (Langley, 1977)
, sc-leaf, text width=48em
Knowledge discovery
conjectures (Raayoni et al., 2021; Wu and Tegmark, 2019; Chen et al., 2016; Buchberger et al., 2006; Fajtlowicz, 1988),
theorem-proving (Hubert et al., 2024b; Trinh et al., 2024), concepts (Hakuk and Reich, 2020; Iten et al., 2020; Lenat and Brown, 1984),
molecular structures (Abramson et al., 2024; Zambaldi et al., 2024; Jumper et al., 2021; Lindsay, 1980)
, sc-leaf, text width=45em
Scientific process
automation
literature review (Skarlinski et al., 2024), idea generation (Si et al., 2024; Baek et al., 2024; Wang et al., 2024a; Girotra et al., 2023; Wang et al., 2023),
hypothesis generation (Wang et al., 2024a; Ghafarollahi and Buehler, 2024; Majumder et al., 2024; Qi et al., 2023; Yang et al., 2023; Sybrandt et al., 2020),
paper writing (Altmäe et al., 2023; Wang et al., 2019),
AI Scientist (Lu et al., 2024a; Ifargan et al., 2024; Li et al., 2024b; Liu et al., 2024; Boiko et al., 2023; King et al., 2009)
, sc-leaf, text width=50em
3. Linguistic Creativity
The creative aspect of language in linguistics has been discussed since the early days (Chomsky, 1965). Chomsky, in this paper, attributes creativity mainly to the essential property of language to provide means to express many thoughts indefinitely. However, several linguists since Chomsky have argued against using this characterization since it does not align with the everyday definition of creativity (Bergs, 2019; Sampson, 2017; Zawada, 2006). Chomsky’s theory of grammar might generate an infinite number of sentences; it, however, relies on a fixed set of rules, while creativity requires deviation from rules. In this sense, Sampson (2017) suggests distinguishing between F-creativity (fixed) and E-creativity (extending), where F-creativity refers to the Chomskian interpretation of linguistic creativity (a.k.a productivity in morphology) and E-creativity corresponds to the real linguistic innovation such as metaphors, jokes, neologisms, etc. Some recent works have explored the F-creativity of large language models and found that this task is challenging in general and even harder in more morphological complex languages (Ismayilzada et al., 2024; Anh et al., 2024; Weissweiler et al., 2023). Most past works however have focused on studying the E-creativity of AI systems which we review in the following sections.
3.1. Humor
Humor is one of the most common ways humans creatively use language to express their ideas and feelings. Early works to model humor focused on hand-crafted linguistic templates and wordplay (Stock and Strapparava, 2005; Taylor and Mazlack, 2004; Raskin and Attardo, 1994). Subsequent works have leveraged language’s lexical and syntactic properties as humor-specific features for humor detection (Liu et al., 2018b; Yang et al., 2015; Zhang and Liu, 2014). The growing interest in computational humor in recent years has resulted in several shared tasks organized by the NLP community (Meaney et al., 2021; Hossain et al., 2020a; Van Hee et al., 2018; Castro et al., 2018; Potash et al., 2017; Miller et al., 2017). Latest works have developed methods based on neural networks and language models to generate and detect humorous content (Ravi et al., 2024; Arora et al., 2022; Peyrard et al., 2021; Annamoradnejad and Zoghi, 2020; Ziser et al., 2020; Amin and Burghardt, 2020; Hossain et al., 2019; Chen and Soo, 2018; Bertero and Fung, 2016), jokes (Horvitz et al., 2024; Tang et al., 2022; Xie et al., 2021; Weller and Seppi, 2019; Ren and Yang, 2017), puns (He et al., 2019; Mittal et al., 2022; Yu et al., 2018), and sarcasm (Chakrabarty et al., 2022c, 2020a). Several datasets have also been proposed to benchmark the humor capacity of LLMs in several languages including English (Horvitz et al., 2024; Jain et al., 2024; Tang et al., 2022; Meaney et al., 2021; Hossain et al., 2020b, 2019; Miller et al., 2017), Chinese (Zhang et al., 2019b), Italian (Buscaldi and Rosso, 2007), Spanish (Castro et al., 2017), Dutch (Winters and Delobelle, 2020) and Russian (Blinov et al., 2019). Computational humor has also been explored in multimodal settings involving images, audio, and video in addition to text (Hessel et al., 2023a; Xie et al., 2023b; Hasan et al., 2019; Bertero and Fung, 2016; Radev et al., 2016; Shahaf et al., 2015). While the latest methods particularly LLMs show an impressive ability to generate and detect humorous content, recent work has also shown that these models still fail to reliably understand humor (Hessel et al., 2023a; Góes et al., 2023a; Kocoń et al., 2023; Borji, 2023a) and generated jokes typically lack diversity (Jentzsch and Kersting, 2023) which has been attributed to training on less diverse humor datasets (Baranov et al., 2023). Creative training frameworks have also been developed to improve the humor generation capabilities of LLMs (Zhong et al., 2023). We refer the reader to Amin and Burghardt (2020) for an in-depth survey on computational humor.
3.2. Figurative Language
Figurative language is a term in language studies encompassing various figures of speech like hyperbole, similes and metaphor (Veale et al., 2016; Roberts and Kreuz, 1994; Paul, 1970). These elements can be used to achieve a range of communicative goals. Figurative language generation involves transforming a text into a specific figure of speech while maintaining the original meaning (Lai and Nissim, 2024). Generating figurative language requires an understanding of abstract concepts, commonsense reasoning, and an ability to make analogies and deviate from literal meaning. Recent works have shown that language models with injected commonsense knowledge can generate textual and visual metaphors (Chakrabarty et al., 2023f, 2021b), similes (Chakrabarty et al., 2020b; He et al., 2023), idioms (Chakrabarty et al., 2021a) and hyperboles (Tian et al., 2021). Chakrabarty et al. (2023a) reveals that metaphors generated by large language models are often incoherent or cliched. Chakrabarty et al. (2023a) highlights the following example of such a metaphor generated by an LLM: “- However, she managed to laugh louder and louder until her laughter transformed into an embrace of the sun’s atmosphere.” We refer the reader to Lai and Nissim (2024) and Abulaish et al. (2020) for an in-depth survey on the automatic generation and detection of figurative language.
3.3. Lexical Innovation
Understanding and generating novel words or word compounds is a challenging linguistic task that often requires creativity, commonsense knowledge, and an ability to generalize over seen concepts (Costello and Keane, 2000; Wisniewski, 1997). Similar noun compounds might have different meanings based on our common understanding. For example, knowing that “chocolate croissant” means a “croissant filled with chocolate” does not necessarily imply that “chocolate bunny” would mean “a bunny filled with chocolate”, but rather a bar of “chocolate in the shape of a bunny”. Several works have evaluated and analyzed language models on the task of interpreting and predicting the emergence of these noun compounds and found that models generally show a moderate performance (Coil and Shwartz, 2023; Dhar and van der Plas, 2019; Kuznetsova et al., 2013). Other works have successfully trained neural networks to generate neologisms (i.e. newly coined words or phrases) (Das and Ghosh, 2017; Lencione et al., 2022). On the other hand, previous works have also shown that large language models can fail at linguistic generalization tasks such as morphologically deriving new words from nonce roots (Ismayilzada et al., 2024; Weissweiler et al., 2023) and can occasionally duplicate large amounts of text from its training data (McCoy et al., 2021). Similarly, a recent work explores the linguistic creativity of both large language models and humans by reconstructing their text output from the existing text snippets on the web and finds that the seemingly remarkable creativity of model outputs may be in large part attributable to the remarkable creativity of human-written texts on the web (Lu et al., 2024b).
4. Creative Problem-Solving
Creative problem-solving is the mental process of searching and coming up with creative solutions to a given problem (Duncker and Lees, 1948). It is a challenging task for machines as it not only requires creativity but also commonsense reasoning, and compositional generalization (Davidson et al., 2022). In addition, creatively solving a problem is usually characterized by two kinds of thinking, namely, convergent and divergent thinking, and involves deep abstraction and analogy-making abilities.
4.1. Convergent Thinking
Convergent thinking models creativity in terms of an ability to produce a single optimal solution for a given problem (Guilford, 1967). This type of creativity requires one to be able to associate seemingly remote ideas and converge to a unified solution. To evaluate this type of thinking in humans, psychologists have come up with several creativity tests such as Remote Associates Test (RAT) (Mednick, 1962) and insight problems (Webb et al., 2017). For example, the goal in RAT is to connect several unrelated words with one concept, e.g. words “broken”, “clear” and “eye” can be connected with the word “glass”.
Language models have recently been evaluated on problems that require convergent thinking. Lin et al. (2021) tests language models on solving riddles that require creativity and commonsense and finds a significant gap between model and human performance. Naeini et al. (2023) uses the popular British quiz show Only Connect’s Connecting Wall that mimics RAT formulation with built-in, deliberate red herrings (i.e. misleading stimuli or distractors) and evaluates large language models such as GPT-4 on these problems. They report poor model performance and show that models are highly susceptible to distractors in the input and manifest a form of fixation effect (a.k.a functional fixedness or Einstellung effect). This type of cognitive bias forces the model to fixate on its past knowledge and prevents it from thinking “out-of-the-box”. The same effect is also found when models are evaluated on everyday problems involving unconventional use of objects (Tian et al., 2023). Very recently, large language models such as GPT-4o have been evaluated on the popular New York Times game Connections and have been found to struggle with associating encyclopedic and linguistic knowledge at an abstract level (Samadarshi et al., 2024). Another study has revealed that language models also fall short of demonstrating human-like creativity in code generation (Lu et al., 2024c).
4.2. Divergent Thinking
Divergent thinking requires one to conceptualize multiple, often seemingly disconnected ideas (Guilford, 1967). It essentially plays the opposite role to convergent thinking and therefore, the goal is to start with a unified idea and diverge from this idea into the space of all ideas to find the ones that are relevant to the task at hand. Psychologists have also devised creativity tests to evaluate humans’ divergent thinking abilities, such as Alternate Uses Test (AUT) (Guilford, 1967) and Torrance Tests of Creative Thinking (TTCT). AUT tests creativity based on whether the participant can come up with unusual (creative) uses for an everyday object and the results are typically evaluated either manually or using semantic distance. For example, a “brick” can be used as a “paperweight” or “to break a window” and “coffee cup” can be used as “small bowl”, or “a hat for an elf” etc. TTCT consists of several verbal and non-verbal tasks such as imagining impossibilities or the consequences of actions. Works evaluating GPT-3 (Brown et al., 2020b) and GPT-4 (OpenAI, 2023) on these tests report near-human performance results (Hubert et al., 2024a; Zhao et al., 2024; Guzik et al., 2023; Koivisto and Grassini, 2023; Góes et al., 2023b; Stevenson et al., 2022). Other tests that highly correlate with human creativity measured by AUT have also been proposed such as the task of naming unrelated words (Olson et al., 2021). Chen and Ding (2023) uses this test to evaluate the creativity of large language models and finds that models outperform humans.
While language models perform strongly on AUT-like divergent thinking tasks, they, however, struggle when these tasks require some form of lateral thinking or “thinking out-of-the-box”. For example, recent works have found that defying default commonsense associations and modeling unexpected or unlikely situations are challenging for large language models (Jiang et al., 2023; Tian et al., 2023; Zhao et al., 2023a). Figure 4(a) illustrates a creative problem-solving example from Tian et al. (2023) that involves unconventional use of everyday objects.
4.3. Abstraction and Analogy-Making
Conceptual abstraction and analogy-making lie at the core of human cognition and intelligence (Chollet, 2019; Mitchell, 2021; Hofstadter, 2001). These are abilities that enable humans to generalize to new domains, invent novel concepts, and make useful and often surprising connections between concepts. In other words, abstraction and analogy-making serve as foundational building blocks for creative thinking.
4.3.1. Abstraction
As the cornerstone of human intelligence, abstraction, and abstract reasoning are typically evaluated using visual IQ tests in humans. Popular examples of these tests are RAVEN progressive matrices (Raven, 1938), Bongard problems (Bongard, 1970) and recently introduced Kandinsky Patterns (Holzinger et al., 2019), Abstraction and Reasoning Corpus (ARC) (Chollet, 2019) and its variations (Moskvichev et al., 2023). These tests require the participants to identify and complete an abstract visual pattern based on given examples. Although several attempts have been made to solve these tasks using both symbolic-based and neural network-driven approaches (Lorello et al., 2024; Mirchandani et al., 2023; Hu et al., 2023; Xu et al., 2022; Santoro et al., 2018), modern AI systems still struggle with solving RAVEN-like (Ahrabian et al., 2024; Gendron et al., 2023; Odouard and Mitchell, 2022; Zhang et al., 2019a) and ARC-like tasks (Moskvichev et al., 2023; Mitchell et al., 2023; Xu et al., 2023; Odouard and Mitchell, 2022; Zhang et al., 2021). Analysis of abstraction via a serial reproduction task (Langlois et al., 2021) where participants are asked to produce a textual stimulus for the next participant upon observing a visual stimulus and vice versa, has suggested that GPT-4 unlike humans relies heavily on linguistic representations even in vision-only paradigm (Kumar et al., 2024). Figure 5 illustrates an example from the ARC task (Chollet, 2019). The problems in this corpus are quite hard to solve to the extent that this task has been recognized as the de facto benchmark for measuring progress towards Artificial General Intelligence (AGI) and a public competition with a grand prize of $ has recently been launched222https://arcprize.org/. At the time of writing this paper, the highest score is far from from the passing threshold of (human-level).
4.3.2. Analogy-Making
In its basic form, analogy-making is the ability to identify a relation between two concepts and apply it to a new concept. For example, Paris is to France as Tokyo is to Japan (i.e. capital:country relation). Early approaches to computational analogy-making were symbolic-based and required extensive hand-coded input i.e. structured representations of both the entities and their relations (Turney, 2008; Gentner, 1983; Falkenhainer et al., 1989). Later, word embedding models based on neural networks were shown to exhibit analogy-making abilities at the word level (Jacob et al., 2023; Petersen and van der Plas, 2023; Marquer et al., 2022; Ushio et al., 2021; Gladkova et al., 2016; Mikolov et al., 2013). However, word-level analogies are generally simple and do not reflect the typical analogical reasoning humans perform in everyday life about complex situations. While some works have argued for emergent analogical reasoning abilities of large language models (Webb et al., 2022; Yasunaga et al., 2023; Hu et al., 2023), other works have shown that these models lack the robustness and generality exhibited by humans when it comes to long text analogies (Wijesiriwardene et al., 2023; Zhu and de Melo, 2020), scientific analogies (Yuan et al., 2023; Czinczoll et al., 2022), story analogies (Jiayang et al., 2023; Sourati et al., 2023; Sultan and Shahaf, 2022; Nagarajah et al., 2022), visual analogies (Opielka et al., 2024; Bitton et al., 2022; Zhang et al., 2019a) and complex analogical reasoning (Musker et al., 2024; Lewis and Mitchell, 2024). Figure 4(b) illustrates an example from the analogical reasoning over narratives benchmark (Sourati et al., 2023).
5. Artistic Creativity
Artistic creativity is the ability to produce original, imaginative, and expressive works in various art forms, such as creative writing, poetry, visual arts, music, dance, theater, and more. In this section, we will focus on the advancements made in AI to produce creative stories, poetry, visual, and musical content automatically and also point out the remaining challenges.
5.1. Story Generation
Storytelling is at the heart of human communication, a powerful tool for connecting and conveying ideas effectively (Suzuki et al., 2018). It requires creativity, particularly when crafting an engaging and compelling narrative. Early approaches to this task focused on algorithmic planning based on character traits and social and physical constraints (Meehan, 1977; Lebowitz, 1984). With the advent of powerful neural networks, the focus shifted to machine learning-based data-driven approaches (Du and Chilton, 2023; Hong et al., 2023; Akoury et al., 2020; Louis and Sutton, 2018; Fan et al., 2018). While these networks are trained on large datasets of stories and prompted to directly generate a new story, often producing locally coherent narratives, they suffer from long-term coherence, irrelevance to premise, and repetitive text problems (Yao et al., 2019). Latest approaches have addressed these problems by using content planning and recursive prompting techniques where a high-level plan of the story is first generated, followed by iterative prompting that aims to generate the story in multiple steps based on the plan (Yao et al., 2019; Goldfarb-Tarrant et al., 2020; Yang et al., 2022). Since language models are designed for open-ended text generation, controlling the attributes of its generations (e.g. topic, characters) is another major challenge (Dathathri et al., 2019). While several methods have been developed towards controllable text generation (Dathathri et al., 2019; Pascual et al., 2021; Paul and Frank, 2021; Tambwekar et al., 2018; Chung et al., 2022; Rashkin et al., 2020), language models still struggle with following constraints (Sun et al., 2023). In addition, long-term factual inconsistency and hallucinations still remain as major issues in language model generated texts (Banerjee et al., 2024; Elazar et al., 2021; Tam et al., 2022; Zhang et al., 2023).
Language models have also been evaluated on their ability to produce and judge creative content as a professional writer (Chakrabarty et al., 2023b; Marco et al., 2024). Chakrabarty et al. (2023b) generates short stories from LLMs based on the plots of popular fictional stories published in the New York Times and conducts a fine-grained expert assessment of both model-generated and original stories. Their study shows that LLMs significantly lag behind seasoned writers in producing inherently creative content. Studies also demonstrate that LLMs are unreliable evaluators of creativity (Chhun et al., 2024; Chakrabarty et al., 2023b). Additionally, (Tian et al., 2024) finds that LLM-generated stories are positively homogenous and typically lack suspense and tension.
To complement the shortcomings of LLMs in creative content generation, recently several works have developed frameworks to use these models as creative assistants for humans and these collaborative systems have shown strong performance across domains and editing tasks (Yuan et al., 2022; Schick et al., 2022; Mirowski et al., 2022; Chakrabarty et al., 2023c; Swanson et al., 2021). However, recent works have also demonstrated that the output of human and language model collaboration lacks lexical and idea diversity (Padmakumar and He, 2023; Anderson et al., 2024b). Particularly, adapting language models with human feedback (Ouyang et al., 2022) has been found to be a main contributing factor in diversity reduction (Mohammadi, 2024; Padmakumar and He, 2023; Bai et al., 2022).
5.2. Poetry
Poetry is a form of literary expression that uses rhythmic and often condensed creative language to evoke emotions, convey ideas, or tell stories. Early approaches to poetry generation have been based on hand-crafted templates, heuristics, and linguistic features of the target language which were limited in their expressivity (Oliveira, 2012; Colton et al., 2012; Manurung et al., 2012; Manurung, 2004; Manurung et al., 2000; Racter, 1984; Masterman, 1971; Milic, 1970). However, recent statistical approaches using (recurrent) neural networks (Lau et al., 2018; Ghazvininejad et al., 2016; Zhang and Lapata, 2014) and language models (Popescu-Belis et al., 2023; Belouadi and Eger, 2022; Chakrabarty et al., 2022b; Ormazabal et al., 2022b; Tian and Peng, 2022; Van de Cruys, 2020; Agarwal and Kann, 2020) have been shown to generate high-quality poems. While these generations almost always follow natural poetic style with appropriate rhyme and meter, they typically fail to express a poetically deep meaning (Chakrabarty et al., 2023d). Figure 3 illustrates the qualitative difference between human and machine-generated poems. We refer the reader to (Oliveira, 2017; Elzohbi and Zhao, 2023) for an in-depth survey on automatic poetry generation.
5.3. Visual Creativity
Humans have been producing visual content to convey emotions, concepts, and narratives since ancient times, from cave paintings and hieroglyphics to classical and Renaissance art masterpieces. For centuries, visual creativity was primarily the domain of professional artists, however, the invention of photography in the 19th century and the traditional image editing software, such as Adobe Photoshop in the past few decades enabled ordinary individuals to produce visually creative outputs without the need for formal artistic training. The advancements of AI have further transformed the landscape of visual creativity, pushing the boundaries of what can be created and who can create it. Early works employed Generative Adversarial Networks (GAN) (Goodfellow et al., 2014b) and Convolutional Neural Networks (CNN) (LeCun et al., 1989) to model images (Li and Wand, 2016; van den Oord et al., 2016b, a; Radford et al., 2015) and generate images by applying specific transformations such as style transfer (Karras et al., 2019; Abdal et al., 2019; Karras et al., 2018; Johnson et al., 2016; Dumoulin et al., 2016; Gatys et al., 2015), super-resolution (Ledig et al., 2016; Dong et al., 2014), colorization (Zhang et al., 2016a) and inpainting (Pathak et al., 2016) or learning a generic mapping between two images (Richardson et al., 2020; Huang et al., 2018a; Isola et al., 2016) or conditioning on text (Zhang et al., 2016b; Reed et al., 2016b, a; Mansimov et al., 2015; Yan et al., 2015; Mirza and Osindero, 2014). In recent years, the development of Transformer architecture (Vaswani et al., 2017) and Diffusion models (Ho et al., 2020) has further pushed AI-driven art to new heights. Trained on large amounts of multimodal data, these models are capable of generating from arbitrary instructions not only high-quality images (Shen et al., 2024; Geng et al., 2023; Saharia et al., 2022; Ramesh et al., 2022; Gafni et al., 2022; Brooks et al., 2022; Chakrabarty et al., 2023e; Patashnik et al., 2021; Ruiz et al., 2022; Huang et al., 2022; Hertz et al., 2022; Rombach et al., 2021; Nichol et al., 2021; Ramesh et al., 2021) but also short photo-realistic videos (Brooks et al., 2024; Xing et al., 2023; Gupta et al., 2023; Kondratyuk et al., 2023; Luo et al., 2023; Ho et al., 2022b, a; Singer et al., 2022; Arnab et al., 2021; Yan et al., 2021; Tulyakov et al., 2017; Vondrick et al., 2016).
Despite their impressive quality, AI systems still exhibit trivial errors in their generations. Recent work has shown that these models struggle to effectively compose objects with different attributes and relationships (Murphy et al., 2024; Huang et al., 2023; Zarei et al., 2024; Marcus et al., 2022; Thrush et al., 2022; Leivada et al., 2022; Conwell and Ullman, 2022), fails to reliably capture common syntactic processes such as negation, word order, comparatives etc. (Leivada et al., 2022; Murphy et al., 2024; Marcus et al., 2022), struggles with representing numbers and texts in images (Marcus et al., 2022; Borji, 2023b), often fall short when it comes to accurately depicting the intricate details of human extremities such as hands and fingers (Borji, 2023b) and lacks robust commonsense reasoning ability (Marcus et al., 2022; Thrush et al., 2022; Borji, 2023b; Rassin et al., 2022). Similarly, video generation models often suffer from a lack of reliable spatial reasoning, appearance inconsistency, temporal inalignment, body deformation and occlusion issues (Lei et al., 2024; Brooks et al., 2024).
5.4. Musical Creativity
Music is another major artistic medium that allows individuals to express emotions, ideas, and cultural narratives through sound, often transcending language barriers to connect people across diverse backgrounds and experiences. Automatic music generation using computers has also a long history dating back to the 1950s (Ji et al., 2023). Early attempts at music generation employed rule-based methods (Hiller and Isaacson, 1958), stochastic models (typically Hidden Markov Models) (Farbood and Schöner, 2001; Ames, 1987; Brooks et al., 1957), evolutionary algorithms (Lavrenko and Pickens, 2003; Cope, 1996; Biles et al., 1994) and recurrent neural networks (Eck and Schmidhuber, 2002; Todd, 1989). However, these methods suffered from long-range incoherence and produced only short pieces often with low music quality (Ji et al., 2020).
With the advent of powerful deep generative models, it became possible to capture the long-term structure of polyphonic music. Recent years have seen models that can compose multi-instrument polyphonic pieces using variational auto-encoders (Roberts et al., 2018; Kingma and Welling, 2013), generative adversarial networks (Yang et al., 2017; Dong et al., 2017; Goodfellow et al., 2014a; Yu et al., 2016) and transformers (Yuan et al., 2024; Deng et al., 2024; Qu et al., 2024; Copet et al., 2023; Agostinelli et al., 2023; Dhariwal et al., 2020; Huang and Yang, 2020; Donahue et al., 2019; Payne, 2019; Huang et al., 2018b).
While music generated by these systems often seems quite impressive, an automatic objective evaluation of music composition remains a challenge because of its subjective and complex nature (Yang and Lerch, 2018). It is not yet entirely clear whether the AI-generated pieces are truly novel as past work has found that deep learning-based music generation models gradually copy increasingly distinctive chunks from pieces in the training set (Yin et al., 2021). Recent studies also show that humans exhibit a preference for human compositions over AI compositions and they report something “off” about the latter such as a lack of sense of coherence or consistency, odd note choices, unnecessary complexity, repetition, uninterestingness, and failure to come to a resolution (Sarmento et al., 2024).
6. Scientific Creativity
Scientific creativity refers to the ability to generate novel ideas, approaches, or solutions within the realm of science, often leading to new discoveries, theories, or technologies. Automating the process of scientific discovery (Kramer et al., 2023; Savage, 2012; Waltz and Buchanan, 2009) has long been a focus of AI research dating back to the 1970s when early attempts mainly targeted automated equation discovery and symbolic regression and were based on methods such as heuristic search and genetic programming (Schmidt and Lipson, 2009; Todorovski, 1997; Koza, 1994; Dzeroski and Todorovski, 1993; Rzevski et al., 1987; Langley, 1977). Recent methods, however, often employ Bayesian statistics (Guimerà et al., 2020) and neural networks (Chen et al., 2022a; Garcon et al., 2021; Udrescu et al., 2020; Cranmer et al., 2020; Petersen and Landajuela, 2019).
Another line of work has focused on automating the discovery of other scientific knowledge such as generating new mathematical conjectures or theories (Raayoni et al., 2021; Wu and Tegmark, 2019; Chen et al., 2016; Buchberger et al., 2006; Fajtlowicz, 1988), automatically proving theorems (Hubert et al., 2024b; Trinh et al., 2024), discovering new concepts (Hakuk and Reich, 2020; Iten et al., 2020; Lenat and Brown, 1984) and predicting molecular structures (Abramson et al., 2024; Zambaldi et al., 2024; Jumper et al., 2021; Lindsay, 1980) among others. A notable example in this area is the recent AlphaFold model (Jumper et al., 2021) that can predict millions of intricate 3D protein structures which has the potential to significantly accelerate research in biology.
While the above works have mainly targeted one aspect of the scientific process, namely the automatic discovery of particular scientific knowledge, there have also been attempts to partially or fully automate the entire process itself. The scientific process typically starts with a scientific question or an idea that is then used to formulate a hypothesis, followed up with designing and running experiments and analyzing results to test the validity of the hypothesis and ends with communicating the findings to the scientific community (Kramer et al., 2023). Recently, the field has seen a surge in the development of frameworks using neural networks and especially, large language models to automate several steps of the scientific process such as literature review (Skarlinski et al., 2024), idea generation (Si et al., 2024; Baek et al., 2024; Wang et al., 2024a; Girotra et al., 2023; Wang et al., 2023), hypothesis generation (Wang et al., 2024a; Ghafarollahi and Buehler, 2024; Majumder et al., 2024; Qi et al., 2023; Yang et al., 2023; Sybrandt et al., 2020) and paper writing (Altmäe et al., 2023; Wang et al., 2019). Yet other works have gone further to introduce a so-called “AI Scientist” that automates almost the entire scientific process from the idea generation to experiment execution to even paper writing (Lu et al., 2024a; Ifargan et al., 2024; Li et al., 2024b; Liu et al., 2024; Boiko et al., 2023; King et al., 2009).
While the latest advancements in the automation of scientific creativity are remarkable, these results should be taken with a grain of salt. Most of the recent end-to-end automation frameworks are powered by LLMs, hence, they face the same challenges and issues we discussed in the previous sections such as hallucinations, lack of content diversity, novelty, and robust reasoning capabilities. For example, one of the aforementioned large-scale idea generation studies (Si et al., 2024) finds that out of LLM-generated ideas only are unique. Their qualitative analysis also reveals some common failure modes such as vague implementation details, misuse of datasets, inappropriate baselines, unrealistic assumptions, and overall poorly-motivated ideas. Similarly, another study benchmarking the machine learning experimentation capabilities of LLMs reports hallucinations and poor planning as some of the major issues with these models (Huang et al., 2024b).
Another important aspect of scientific discovery is the explainability (Li et al., 2021) which helps humans prevent or better prepare for a possible future technological singularity (Good, 1965; Ulam, 1958). However, current LLMs are largely black-box AI systems, and allowing them to make discoveries that are incomprehensible to humans may lead to a scenario where human knowledge is left far behind the machine’s knowledge resulting in machines that humans can’t control (Good, 1965).
7. Creativity and Copyright
Our brief survey into the methods used to produce creative outputs showed that the predominant approach is currently the generative deep learning techniques, especially LLMs. These models typically have billions of adjustable parameters (Brown et al., 2020b) and are trained on massive amounts of public and private data (Raffel et al., 2023). Consequently, these models have been found to exhibit strong memorization skills (Carlini et al., 2022, 2018) such that they can sometimes copy large passages (Chang et al., 2023; McCoy et al., 2021) or replicate images from their training data (Somepalli et al., 2022). While this could be of little concern when the duplicated content is public and generic, however, the training datasets of popular LLMs are often undisclosed and can include private and copyrighted data leading to concerns about copyright infringement and privacy violation (Franceschelli and Musolesi, 2021a). Although several approaches have been developed to detect (Duarte et al., 2024; Shi et al., 2024; Li et al., 2024a; Oren et al., 2023; Carlini et al., 2021) and prevent (Hans et al., 2024; Zhao et al., 2022; Kandpal et al., 2022; Ippolito et al., 2022) unintended memorization in LLMs, major questions concerning the use of copyrighted material for training and authorship of the machine-generated content remain unresolved (Abbott and Rothman, 2023; Franceschelli and Musolesi, 2021a). Recent lawsuit between The New York Times and OpenAI (Grynbaum and Mac, 2023) and the class action333A class action is a type of civil lawsuit brought on behalf of many similarly situated people who have been harmed in the same way by the same entity. against Stable Diffusion, Midjourney, and DeviantArt (Brittain, 2023) have further highlighted the urgency of the matter and the need for clear legal frameworks that address the complex issues surrounding intellectual property rights, ethical use, and the boundaries of fair use in AI development.
More specifically, two key questions concerning copyright and authorship are of interest and here we briefly discuss them with respect to machine-generated artworks. We refer the reader to Franceschelli and Musolesi (2021a) and Abbott and Rothman (2023) for a detailed discussion of these questions.
Is it copyright infringement to use protected works for the training of generative models?
To answer this question, we will review the implications of the existing relevant laws from the US and EU. Under the US Law Code, reproduction of a copyrighted work can be allowed if the use can be considered a fair use of the work (Netanel, 2011). Analyzing the criteria used to determine fair use, Franceschelli and Musolesi (2021a) concludes that it is not straightforward to assess this for generative deep models and if these models do not add any form of novelty to their output. Their outputs may not qualify for fair use, which can potentially derail the progress in AI (Sobel, 2017).
Under the EU law, on the other hand, the use of lawfully accessible protected work for training is permitted as long as 1) the rightsholder of the used data has not reserved the right to withhold its data from being reproduced and 2) the accessed data is retained only for the time required for the purposes of scientific research (Franceschelli and Musolesi, 2021a). However, Franceschelli and Musolesi (2021a) also notes while the second criterion is reasonably easy to satisfy, the first criterion is hard to verify in practice because nowadays, models are being trained on large amounts of data published on the internet for which there is no centralized repository allowing to filter reservation-free works. Finally, whether providers of such a repository or the developers of the models should be forced to perform this check is unclear.
Who is the author (if someone) or who will own the copyright on the generated work?
To answer this question, first, we have to make a distinction between the AI-assisted and AI-generated content. If the generative model is merely used as a tool to assist a human to produce a creative artwork, then the human will be considered the author and own the copyright. However, it becomes tricky to determine the authorship and the copyright status of the work that is generated mostly by AI with little human involvement (e.g. human as prompter). First, let’s consider the authorship issue. Some have argued that for an author to exist there has to be a message that the author wants to convey through their work, but since no one can reliably predict the output of a generative model, no author exists (Ginsburg, 2018). However, if we suppose that there is an author, then there are mainly three contenders in question: 1) the person who developed the AI model (developers) 2) the person who used the AI model to produce creative work (users), and finally 3) the AI model itself. Since the existing laws in most countries only attribute copyright to a human, but not to a machine, the main tension is around deciding whether to attribute the authorship (also the copyright) to users or developers (Abbott and Rothman, 2023; dos Santos and Machado, 2020; Deltorn and Macrez, 2018; Guadamuz, 2017). Some have argued that the criterion to determine authorship should center around the incentives to create and promote the work, not the ideation and creation of the work itself (Miller, 1993) and since the users of the generative models are best positioned to do so, they should be assigned the authorship (Samuelson, 1986; Denicola, 2016; Franceschelli and Musolesi, 2021a). Another argument supporting this assignment is by ruling out the developer as the author since they just create the potentiality for the creation of the output, but not its actuality (Franceschelli and Musolesi, 2021a; Samuelson, 1986). Using the analogy proposed by Ralston (2005), it would be similar to claim a knife manufacturer is more responsible for murder than the person who wielded the knife or assigning copyright to the teacher of the painter rather than the painter himself/herself (Franceschelli and Musolesi, 2021a). Finally, arguments in favor of AI authorship have also been made recently suggesting that this will promote transparency, efficient allocations of rights, and even counterintuitively protect human authors (Abbott and Rothman, 2023).
8. Future Directions
In the previous sections, our brief exploration into the creativity of modern AI systems revealed that these systems exhibit some capacity for producing linguistically and artistically creative outputs and thinking creatively. However, true human-like creative abilities seem to be still out of reach, as indicated by challenges with tasks demanding creative problem-solving (Jiang et al., 2023; Tian et al., 2023), abstract reasoning (Mitchell et al., 2023; Gendron et al., 2023), and compositionality (Huang et al., 2023; Murphy et al., 2024). Some studies also highlighted major issues in machine outputs, such as lack of originality (Lu et al., 2024b; Chakrabarty et al., 2023b), diversity (Anderson et al., 2024b; Padmakumar and He, 2023) and incoherence (Tam et al., 2022; Sarmento et al., 2024). From the Four-C model perspective, these models seem to manifest only mini-c or little-c type of creativity while Pro-C and Big-C creativity remain elusive. Similarly, current AI models exhibit strong interpolation and moderate extrapolation capabilities. However, they are still far from truly inventing a completely new type of creative artefact. In this section, we discuss potential research directions that can help us better measure and improve the creative abilities of AI systems.
8.1. Evaluating Creativity
8.1.1. Creative Process
Cognitive scientists and psychologists have proposed theoretical frameworks to evaluate creativity such as characterizing it based on input, process and output (Jordanous, 2012; Pease et al., 2002; Ritchie, 2007) or four Ps: person, product, process and press (Jordanous, 2016; Rhodes, 1961). A common thread across all these theories is their emphasis on evaluating the process aspect of creativity. However, most works in AI, including the ones we reviewed before, focus on evaluating and analyzing creativity from the output or product perspective. Creative process, on the other hand, is an equally (or perhaps more) important aspect of creativity that can tell us how creativity “arises” in the first place and what the key ingredients involved (Colton, 2008). For example, in computational creativity, one popular theory by Boden (2004) defines the creative process in terms of manipulations over a conceptual space. This theory divides creativity into three types: combinatorial that makes unfamiliar connections between familiar concepts (e.g. creating hybrid fictional creatures such as pegasus, sphinx, or mermaid), exploratory that involves an open-ended search in a conceptual space (e.g. a novel chess move) and transformational that requires a fundamental transformation of the existing conceptual space (e.g. non-Euclidean geometry444https://en.wikipedia.org/wiki/Non-Euclidean_geometry). Another popular theory by Wallas (1926) explains the creative process in four stages akin to how scientists develop their ideas: preparation stage where the problem at hand is investigated in all directions, information is gathered and analyzed, incubation stage where you step back from the problem and let your unconscious work through it in the background, inspiration stage where a creative insight is typically realized (an “Aha!” or “Eureka!” moment) and finally, verification stage where you test, evaluate and build further on your creative idea to make it perhaps useful.
While AI systems produce seemingly creative outputs, the nature of the creative process they employ (if any) remains unknown. Only very recently attempts have been made to study the creative process of machines (Nath et al., 2024) which analyzes the creative process of language models and humans to solve AUT task using response pathways (persistent vs. flexible) (Baas et al., 2013; Nijstad et al., 2010) and finds that while humans are able to follow a mixture of pathways, models are biased towards either one of them pointing to a limited capacity. Hence, analyzing the creative process of machines is an emerging and exciting area for which much work remains to be explored. We believe a strong collaboration between the computational creativity (Veale and Cardoso, 2019; Colton and Wiggins, 2012) and NLP communities drawing ideas from past research on studying human creative process and techniques from research on (mechanistic) interpretability555https://www.neelnanda.io/mechanistic-interpretability (Saphra and Wiegreffe, 2024; Bereska and Gavves, 2024) could lead to a better understanding of the creative capacity of AI systems.
8.1.2. Dimensions of Creativity
As we discussed earlier, there are many dimensions of creativity, but most works generally focus on evaluation of the novelty and usefulness dimensions. However, surprise, agency and spontaneity dimensions are also equally important. Humans typically communicate an emotion or a deeper meaning through creative products and their creative process is characterized by spontaneous “Aha” or “Eureka” moments coupled with deliberate decisions made at each step of the way 666https://www.newyorker.com/culture/the-weekend-essay/why-ai-isnt-going-to-make-art. However, current AI systems lack agency and are typically trained to generate the most likely output leaving no room for any intentional or spontaneous action (Franceschelli and Musolesi, 2023). Therefore, a holistic evaluation of machine creativity should involve consideration of all these different dimensions that characterize human creativity.
8.2. Improving Creativity
Recent years have seen a surge in human-AI creative collaboration (Vinchon et al., 2023) popularized by the introduction of chat-based products such as ChatGPT777https://chat.openai.com and Gemini888https://gemini.google.com. However, the poor creative capacity of current AI systems necessitates the innovation of new techniques to improve the creativity of their outputs. In this section, we discuss several possible directions to take.
8.2.1. Creative Architectures
As we argued before, current AI architectures optimized for the most likely outcome might have fundamental limitations to exhibit true human-like creativity. In fact, by definition, current AI models are optimized to model the training distribution while creating something new requires the model to diverge from its learned distribution. Therefore, innovating at the architecture level to endow machines with mechanisms to actively diverge from the training data and a capacity for agency and spontaneity might be a necessary step towards robust creativity. An emerging new research area called active divergence attempts to optimize models for creativity using methods such as novelty search, divergent fine-tuning, and objective functions targeting different dimensions of creativity (Broad et al., 2021; Elgammal et al., 2017; Guimaraes et al., 2017; Bunescu and Uduehi, 2019).
8.2.2. Creative Prompt Engineering
Natural language-based interaction with the current AI systems has created an intuitive playground to elicit more capabilities from these systems (Qiao et al., 2022). These so-called prompt engineering techniques have also been shown to enhance the creativity of large language models (Nair et al., 2024; Mehrotra et al., 2024; Tian et al., 2023; Summers-Stay et al., 2023). We can draw ideas from psychology that has shown techniques such as brainstorming (Osborn, 1957), competence injection (Liu and Xu, 2020) and threatening situations (Riley and Gabora, 2019) stimulate creativity of humans. Hence, designing prompts inspired by these methods is a promising direction to get the most out of future AI systems.
8.2.3. Creative Decoding
An important component in natural language generation is the decoding strategy which is a significant contributor to the quality of the generation (Meister et al., 2022). Past work has shown that simple greedy decoding results in repetitive and uninteresting generations (Li et al., 2023) and numerous powerful decoding algorithms have been developed to address these problems (Holtzman et al., 2019; Fan et al., 2018; Meister et al., 2023). These decoding strategies mainly target generating human-like text and do not directly target creativity. A popular approach is to increase the randomness of the output by increasing the temperature parameter, however, recent work shows that this parameter is weakly correlated with the novelty of the output (Peeperkorn et al., 2024). A potential direction could be to devise new creative decoding algorithms that go beyond the temperature parameter by injecting semantic planning or intentionality (Franceschelli and Musolesi, 2024) and employing information-theoretic measures of novelty, utility, and surprise (Bunescu and Uduehi, 2022; Kuznetsova et al., 2013; Heinen and Johnson, 2017).
9. Conclusion
In conclusion, while the rapid advancements in AI, particularly through state-of-the-art models such as large language models, diffusion models, etc., have demonstrated impressive capabilities in generating creative outputs, the question of genuine machine creativity remains unresolved. This survey has explored key areas of linguistic creativity, creative problem-solving, and artistic and scientific creativity, providing a comprehensive overview of the state of AI creativity. We also discussed pressing copyright and authorship issues with generative artworks, highlighted major challenges facing current AI systems and proposed potential research directions on how to evaluate and improve the creativity of these systems. We believe our suggestions can help future research to determine if machines can achieve a human-like creative process, ultimately enriching our understanding of artificial intelligence and its capabilities.
Acknowledgements.
We gratefully acknowledge the support of Swiss National Science Foundation (grant 205121_207437: C - LING).References
- (1)
- Abbott and Rothman (2023) Ryan Abbott and Elizabeth Rothman. 2023. Disrupting Creativity: Copyright Law in the Age of Generative Artificial Intelligence. Elsevier (2023).
- Abdal et al. (2019) Rameen Abdal, Yipeng Qin, and Peter Wonka. 2019. Image2StyleGAN: How to Embed Images Into the StyleGAN Latent Space? 2019 IEEE/CVF International Conference on Computer Vision (ICCV) (2019), 4431–4440. https://api.semanticscholar.org/CorpusID:102350964
- Abramson et al. (2024) Josh Abramson, Jonas Adler, Jack Dunger, Richard Evans, Tim Green, Alexander Pritzel, Olaf Ronneberger, Lindsay Willmore, Andrew J. Ballard, Joshua Bambrick, Sebastian W. Bodenstein, David A. Evans, Chia-Chun Hung, Michael O’Neill, David Reiman, Kathryn Tunyasuvunakool, Zachary Wu, Akvilė Žemgulytė, Eirini Arvaniti, Charles Beattie, Ottavia Bertolli, Alex Bridgland, Alexey Cherepanov, Miles Congreve, Alexander I. Cowen-Rivers, Andrew Cowie, Michael Figurnov, Fabian B. Fuchs, Hannah Gladman, Rishub Jain, Yousuf A. Khan, Caroline M. R. Low, Kuba Perlin, Anna Potapenko, Pascal Savy, Sukhdeep Singh, Adrian Stecula, Ashok Thillaisundaram, Catherine Tong, Sergei Yakneen, Ellen D. Zhong, Michal Zielinski, Augustin Žídek, Victor Bapst, Pushmeet Kohli, Max Jaderberg, Demis Hassabis, and John M. Jumper. 2024. Accurate structure prediction of biomolecular interactions with AlphaFold 3. Nature 630, 8016 (01 Jun 2024), 493–500. https://doi.org/10.1038/s41586-024-07487-w
- Abulaish et al. (2020) Muhammad Abulaish, Ashraf Kamal, and Mohammed J. Zaki. 2020. A Survey of Figurative Language and Its Computational Detection in Online Social Networks. ACM Trans. Web 14, 1, Article 3 (Feb. 2020), 52 pages. https://doi.org/10.1145/3375547
- Agarwal and Kann (2020) Rajat Agarwal and Katharina Kann. 2020. Acrostic Poem Generation. arXiv:2010.02239 [cs.CL] https://arxiv.org/abs/2010.02239
- Agostinelli et al. (2023) Andrea Agostinelli, Timo I. Denk, Zalán Borsos, Jesse Engel, Mauro Verzetti, Antoine Caillon, Qingqing Huang, Aren Jansen, Adam Roberts, Marco Tagliasacchi, Matthew Sharifi, Neil Zeghidour, and Christian Havnø Frank. 2023. MusicLM: Generating Music From Text. ArXiv abs/2301.11325 (2023). https://api.semanticscholar.org/CorpusID:256274504
- Ahrabian et al. (2024) Kian Ahrabian, Zhivar Sourati, Kexuan Sun, Jiarui Zhang, Yifan Jiang, Fred Morstatter, and Jay Pujara. 2024. The Curious Case of Nonverbal Abstract Reasoning with Multi-Modal Large Language Models. ArXiv abs/2401.12117 (2024). https://api.semanticscholar.org/CorpusID:267068363
- Akoury et al. (2020) Nader Akoury, Shufan Wang, Josh Whiting, Stephen Hood, Nanyun Peng, and Mohit Iyyer. 2020. STORIUM: A Dataset and Evaluation Platform for Machine-in-the-Loop Story Generation. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, Online, 6470–6484. https://doi.org/10.18653/v1/2020.emnlp-main.525
- Aleinikov et al. (2000) A.G. Aleinikov, S. Kackmeister, and R. Koenig. 2000. Creating Creativity: 101 Definitions (what Webster Never Told You). Alden B. Dow Creativity Center Press. https://books.google.ch/books?id=M2QpAAAACAAJ
- Altmäe et al. (2023) Signe Altmäe, Alberto Sola-Leyva, and Andres Salumets. 2023. Artificial intelligence in scientific writing: a friend or a foe? Reproductive BioMedicine Online 47 (04 2023). https://doi.org/10.1016/j.rbmo.2023.04.009
- Amabile (1983) T. M. Amabile. 1983. The social psychology of creativity: A componential conceptualization. Journal of Personality and Social Psychology 45 (1983), 357–376. https://api.semanticscholar.org/CorpusID:145719543
- Ames (1987) Charles Ames. 1987. Automated composition in retrospect: 1956-1986. Leonardo (1987), 169–185.
- Amin and Burghardt (2020) Miriam Amin and Manuel Burghardt. 2020. A Survey on Approaches to Computational Humor Generation. In Proceedings of the 4th Joint SIGHUM Workshop on Computational Linguistics for Cultural Heritage, Social Sciences, Humanities and Literature, Stefania DeGaetano, Anna Kazantseva, Nils Reiter, and Stan Szpakowicz (Eds.). International Committee on Computational Linguistics, Online, 29–41. https://aclanthology.org/2020.latechclfl-1.4
- Anderson et al. (2024a) Barrett R Anderson, Jash Hemant Shah, and Max Kreminski. 2024a. Homogenization Effects of Large Language Models on Human Creative Ideation. In Proceedings of the 16th Conference on Creativity and Cognition (Chicago, IL, USA). Association for Computing Machinery, New York, NY, USA, 413–425. https://doi.org/10.1145/3635636.3656204
- Anderson et al. (2024b) Barrett R Anderson, Jash Hemant Shah, and Max Kreminski. 2024b. Homogenization Effects of Large Language Models on Human Creative Ideation. Proceedings of the 16th Conference on Creativity & Cognition (2024). https://api.semanticscholar.org/CorpusID:267406608
- Anh et al. (2024) Dang Anh, Limor Raviv, and Lukas Galke. 2024. Morphology Matters: Probing the Cross-linguistic Morphological Generalization Abilities of Large Language Models through a Wug Test. In Proceedings of the Workshop on Cognitive Modeling and Computational Linguistics, Tatsuki Kuribayashi, Giulia Rambelli, Ece Takmaz, Philipp Wicke, and Yohei Oseki (Eds.). Association for Computational Linguistics, Bangkok, Thailand, 177–188. https://doi.org/10.18653/v1/2024.cmcl-1.15
- Annamoradnejad and Zoghi (2020) Issa Annamoradnejad and Gohar Zoghi. 2020. ColBERT: Using BERT sentence embedding in parallel neural networks for computational humor. Expert Syst. Appl. 249 (2020), 123685. https://api.semanticscholar.org/CorpusID:254125774
- Arnab et al. (2021) Anurag Arnab, Mostafa Dehghani, Georg Heigold, Chen Sun, Mario Lucic, and Cordelia Schmid. 2021. ViViT: A Video Vision Transformer. 2021 IEEE/CVF International Conference on Computer Vision (ICCV) (2021), 6816–6826. https://api.semanticscholar.org/CorpusID:232417054
- Arnheim (2001) Rudolf Arnheim. 2001. What it Means to be Creative. British Journal of Aesthetics 41 (01 2001). https://doi.org/10.1093/bjaesthetics/41.1.24
- Arora et al. (2022) Aseem Arora, Gaël Dias, Adam Jatowt, and Asif Ekbal. 2022. Transfer Learning for Humor Detection by Twin Masked Yellow Muppets. In Proceedings of the 2nd Conference of the Asia-Pacific Chapter of the Association for Computational Linguistics and the 12th International Joint Conference on Natural Language Processing (Volume 2: Short Papers), Yulan He, Heng Ji, Sujian Li, Yang Liu, and Chua-Hui Chang (Eds.). Association for Computational Linguistics, Online only, 1–7. https://aclanthology.org/2022.aacl-short.1
- Baas et al. (2013) Matthijs Baas, Marieke Roskes, Daniel Sligte, Bernard Nijstad, and Carsten De Dreu. 2013. Personality and Creativity: The Dual Pathway to Creativity Model and a Research Agenda. Social and Personality Psychology Compass 7 (10 2013). https://doi.org/10.1111/spc3.12062
- Babbage (1837) Charles Babbage. 1837. Analytical engine. https://api.semanticscholar.org/CorpusID:63432444
- Baek et al. (2024) Jinheon Baek, Sujay Kumar Jauhar, Silviu Cucerzan, and Sung Ju Hwang. 2024. ResearchAgent: Iterative Research Idea Generation over Scientific Literature with Large Language Models. ArXiv abs/2404.07738 (2024). https://api.semanticscholar.org/CorpusID:269042844
- Baer (2012) John Baer. 2012. Domain Specificity and the Limits of Creativity Theory. Journal of Creative Behavior 46 (2012), 16–29. https://api.semanticscholar.org/CorpusID:45983030
- Bai et al. (2022) Yuntao Bai, Andy Jones, Kamal Ndousse, Amanda Askell, Anna Chen, Nova Dassarma, Dawn Drain, Stanislav Fort, Deep Ganguli, Tom Henighan, Nicholas Joseph, Saurav Kadavath, John Kernion, Tom Conerly, Sheer El-Showk, Nelson Elhage, Zac Hatfield-Dodds, Danny Hernandez, Tristan Hume, Scott Johnston, Shauna Kravec, Liane Lovitt, Neel Nanda, Catherine Olsson, Dario Amodei, Tom B. Brown, Jack Clark, Sam McCandlish, Christopher Olah, Benjamin Mann, and Jared Kaplan. 2022. Training a Helpful and Harmless Assistant with Reinforcement Learning from Human Feedback. ArXiv abs/2204.05862 (2022). https://api.semanticscholar.org/CorpusID:248118878
- Baldi and Itti (2010) Pierre Baldi and Laurent Itti. 2010. Of bits and wows: A Bayesian theory of surprise with applications to attention. Neural networks : the official journal of the International Neural Network Society 23 5 (2010), 649–66. https://api.semanticscholar.org/CorpusID:2070585
- Banerjee et al. (2024) Sourav Banerjee, Ayushi Agarwal, and Saloni Singla. 2024. LLMs Will Always Hallucinate, and We Need to Live With This. arXiv:2409.05746 [stat.ML] https://arxiv.org/abs/2409.05746
- Baranov et al. (2023) Alexander Baranov, Vladimir Kniazhevsky, and Pavel Braslavski. 2023. You Told Me That Joke Twice: A Systematic Investigation of Transferability and Robustness of Humor Detection Models. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 13701–13715. https://doi.org/10.18653/v1/2023.emnlp-main.845
- Barron (1955) Frank Barron. 1955. The disposition toward originality. The Journal of Abnormal and Social Psychology 51, 3 (1955), 478.
- Beaty and Johnson (2020) Roger E. Beaty and Dan Richard Johnson. 2020. Automating creativity assessment with SemDis: An open platform for computing semantic distance. Behavior Research Methods 53 (2020), 757 – 780. https://api.semanticscholar.org/CorpusID:221403286
- Belouadi and Eger (2022) Jonas Belouadi and Steffen Eger. 2022. ByGPT5: End-to-End Style-conditioned Poetry Generation with Token-free Language Models. ArXiv abs/2212.10474 (2022). https://api.semanticscholar.org/CorpusID:254877406
- Bender et al. (2021) Emily M. Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret Shmitchell. 2021. On the Dangers of Stochastic Parrots: Can Language Models Be Too Big? Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency (2021). https://api.semanticscholar.org/CorpusID:262580630
- Bender and Koller (2020) Emily M. Bender and Alexander Koller. 2020. Climbing towards NLU: On Meaning, Form, and Understanding in the Age of Data. In Annual Meeting of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:211029226
- Bereska and Gavves (2024) Leonard Bereska and Efstratios Gavves. 2024. Mechanistic Interpretability for AI Safety - A Review. ArXiv abs/2404.14082 (2024). https://api.semanticscholar.org/CorpusID:269293418
- Bergs (2019) Alexander Bergs. 2019. What, If Anything, Is Linguistic Creativity? Gestalt Theory 41, 2 (2019), 173–183. https://doi.org/doi:10.2478/gth-2019-0017
- Bertero and Fung (2016) Dario Bertero and Pascale Fung. 2016. Deep Learning of Audio and Language Features for Humor Prediction. In Proceedings of the Tenth International Conference on Language Resources and Evaluation (LREC’16), Nicoletta Calzolari, Khalid Choukri, Thierry Declerck, Sara Goggi, Marko Grobelnik, Bente Maegaard, Joseph Mariani, Helene Mazo, Asuncion Moreno, Jan Odijk, and Stelios Piperidis (Eds.). European Language Resources Association (ELRA), Portorož, Slovenia, 496–501. https://aclanthology.org/L16-1079
- Betker et al. (2020) James Betker, Gabriel Goh, Li Jing, TimBrooks, Jianfeng Wang, Linjie Li, Long Ouyang, Juntang Zhuang, Joyce Lee, Yufei Guo, Wesam Manassra, Prafulla Dhariwal, Casey Chu, Yunxin Jiao, and Aditya Ramesh. 2020. Improving Image Generation with Better Captions. https://api.semanticscholar.org/CorpusID:264403242
- Biles et al. (1994) John Biles et al. 1994. GenJam: A genetic algorithm for generating jazz solos. In ICMC, Vol. 94. Ann Arbor, MI, 131–137.
- Bitton et al. (2022) Yonatan Bitton, Ron Yosef, Eli Strugo, Dafna Shahaf, Roy Schwartz, and Gabriel Stanovsky. 2022. VASR: Visual Analogies of Situation Recognition. arXiv:2212.04542 [cs.CV] https://arxiv.org/abs/2212.04542
- Blinov et al. (2019) Vladislav Blinov, Valeria Bolotova-Baranova, and Pavel Braslavski. 2019. Large dataset and language model fun-tuning for humor recognition. In Proceedings of the 57th annual meeting of the association for computational linguistics. 4027–4032.
- Boden (2004) M.A. Boden. 2004. The Creative Mind: Myths and Mechanisms. Routledge. https://books.google.ch/books?id=6Zkm4dz32Y4C
- Boden (1991) Margaret A. Boden. 1991. The creative mind : myths & mechanisms. https://api.semanticscholar.org/CorpusID:143261160
- Boiko et al. (2023) Daniil A. Boiko, Robert MacKnight, and Gabe Gomes. 2023. Emergent autonomous scientific research capabilities of large language models. arXiv:2304.05332 [physics.chem-ph] https://arxiv.org/abs/2304.05332
- Bongard (1970) M. M. Bongard. 1970. Pattern recognition. https://api.semanticscholar.org/CorpusID:267827262
- Borji (2023a) Ali Borji. 2023a. A Categorical Archive of ChatGPT Failures. arXiv:2302.03494 [cs.CL] https://arxiv.org/abs/2302.03494
- Borji (2023b) Ali Borji. 2023b. Qualitative Failures of Image Generation Models and Their Application in Detecting Deepfakes. ArXiv abs/2304.06470 (2023). https://api.semanticscholar.org/CorpusID:257826680
- Brandt (2021) Anthony Brandt. 2021. Defining Creativity: A View from the Arts. Creativity Research Journal 33, 2 (2021), 81–95. https://doi.org/10.1080/10400419.2020.1855905 arXiv:https://doi.org/10.1080/10400419.2020.1855905
- Brittain (2023) B. Brittain. 2023. Artists take new shot at Stability, Midjourney in updated copyright lawsuit. https://www.reuters.com/legal/litigation/artists-take-new-shot-stability-midjourney-updated-copyright-lawsuit-2023-11-30/
- Broad et al. (2021) Terence Broad, Sebastian Berns, Simon Colton, and Mick Grierson. 2021. Active Divergence with Generative Deep Learning - A Survey and Taxonomy. ArXiv abs/2107.05599 (2021). https://api.semanticscholar.org/CorpusID:235794880
- Brooks et al. (1957) Frederick P Brooks, AL Hopkins, Peter G Neumann, and William V Wright. 1957. An experiment in musical composition. IRE Transactions on Electronic Computers 3 (1957), 175–182.
- Brooks et al. (2022) Tim Brooks, Aleksander Holynski, and Alexei A. Efros. 2022. InstructPix2Pix: Learning to Follow Image Editing Instructions. 2023 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2022), 18392–18402. https://api.semanticscholar.org/CorpusID:253581213
- Brooks et al. (2024) Tim Brooks, Bill Peebles, Connor Holmes, Will DePue, Yufei Guo, Li Jing, David Schnurr, Joe Taylor, Troy Luhman, Eric Luhman, Clarence Ng, Ricky Wang, and Aditya Ramesh. 2024. Video generation models as world simulators. (2024). https://openai.com/research/video-generation-models-as-world-simulators
- Brown et al. (2020a) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeff Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020a. Language Models are Few-Shot Learners. ArXiv abs/2005.14165 (2020). https://api.semanticscholar.org/CorpusID:218971783
- Brown et al. (2020b) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020b. Language Models are Few-Shot Learners. arXiv:2005.14165 [cs.CL]
- Bubeck et al. (2023) Sébastien Bubeck, Varun Chandrasekaran, Ronen Eldan, Johannes Gehrke, Eric Horvitz, Ece Kamar, Peter Lee, Yin Tat Lee, Yuanzhi Li, Scott Lundberg, Harsha Nori, Hamid Palangi, Marco Tulio Ribeiro, and Yi Zhang. 2023. Sparks of Artificial General Intelligence: Early experiments with GPT-4. arXiv:2303.12712 [cs.CL]
- Buchberger et al. (2006) Bruno Buchberger, Adrian Crǎciun, Tudor Jebelean, Laura Kovács, Temur Kutsia, Koji Nakagawa, Florina Piroi, Nikolaj Popov, Judit Robu, Markus Rosenkranz, and Wolfgang Windsteiger. 2006. Theorema: Towards computer-aided mathematical theory exploration. Journal of Applied Logic 4, 4 (2006), 470–504. https://doi.org/10.1016/j.jal.2005.10.006 Towards Computer Aided Mathematics.
- Bunescu and Uduehi (2019) Razvan C. Bunescu and Oseremen O. Uduehi. 2019. Learning to Surprise: A Composer-Audience Architecture. In International Conference on Innovative Computing and Cloud Computing. https://api.semanticscholar.org/CorpusID:218606180
- Bunescu and Uduehi (2022) Razvan C. Bunescu and Oseremen O. Uduehi. 2022. Distribution-Based Measures of Surprise for Creative Language: Experiments with Humor and Metaphor. In Proceedings of the 3rd Workshop on Figurative Language Processing (FLP), Debanjan Ghosh, Beata Beigman Klebanov, Smaranda Muresan, Anna Feldman, Soujanya Poria, and Tuhin Chakrabarty (Eds.). Association for Computational Linguistics, Abu Dhabi, United Arab Emirates (Hybrid), 68–78. https://doi.org/10.18653/v1/2022.flp-1.10
- Buscaldi and Rosso (2007) Davide Buscaldi and Paolo Rosso. 2007. Some experiments in humour recognition using the italian wikiquote collection. In International workshop on fuzzy logic and applications. Springer, 464–468.
- Butnariu et al. (2009) Cristina Butnariu, Su Nam Kim, Preslav Nakov, Diarmuid Ó Séaghdha, Stan Szpakowicz, and Tony Veale. 2009. SemEval-2010 Task 9: The Interpretation of Noun Compounds Using Paraphrasing Verbs and Prepositions. In Proceedings of the Workshop on Semantic Evaluations: Recent Achievements and Future Directions (SEW-2009), Eneko Agirre, Lluís Màrquez, and Richard Wicentowski (Eds.). Association for Computational Linguistics, Boulder, Colorado, 100–105. https://aclanthology.org/W09-2416
- Carlini et al. (2022) Nicholas Carlini, Daphne Ippolito, Matthew Jagielski, Katherine Lee, Florian Tramèr, and Chiyuan Zhang. 2022. Quantifying Memorization Across Neural Language Models. ArXiv abs/2202.07646 (2022). https://api.semanticscholar.org/CorpusID:246863735
- Carlini et al. (2018) Nicholas Carlini, Chang Liu, Úlfar Erlingsson, Jernej Kos, and Dawn Xiaodong Song. 2018. The Secret Sharer: Evaluating and Testing Unintended Memorization in Neural Networks. In USENIX Security Symposium. https://api.semanticscholar.org/CorpusID:170076423
- Carlini et al. (2021) Nicholas Carlini, Florian Tramer, Eric Wallace, Matthew Jagielski, Ariel Herbert-Voss, Katherine Lee, Adam Roberts, Tom Brown, Dawn Song, Ulfar Erlingsson, Alina Oprea, and Colin Raffel. 2021. Extracting Training Data from Large Language Models. arXiv:2012.07805 [cs.CR] https://arxiv.org/abs/2012.07805
- Castro et al. (2018) Santiago Castro, Luis Chiruzzo, and Aiala Rosá. 2018. Overview of the HAHA Task: Humor Analysis Based on Human Annotation at IberEval 2018. In IberEval@SEPLN. https://api.semanticscholar.org/CorpusID:51940157
- Castro et al. (2017) Santiago Castro, Luis Chiruzzo, Aiala Rosá, Diego Garat, and Guillermo Moncecchi. 2017. A crowd-annotated spanish corpus for humor analysis. arXiv preprint arXiv:1710.00477 (2017).
- Chakrabarty et al. (2021a) Tuhin Chakrabarty, Yejin Choi, and Vered Shwartz. 2021a. It’s not Rocket Science: Interpreting Figurative Language in Narratives. Transactions of the Association for Computational Linguistics 10 (2021), 589–606. https://api.semanticscholar.org/CorpusID:237371845
- Chakrabarty et al. (2020a) Tuhin Chakrabarty, Debanjan Ghosh, Smaranda Muresan, and Nanyun Peng. 2020a. R^3: Reverse, Retrieve, and Rank for Sarcasm Generation with Commonsense Knowledge. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, Dan Jurafsky, Joyce Chai, Natalie Schluter, and Joel Tetreault (Eds.). Association for Computational Linguistics, Online, 7976–7986. https://doi.org/10.18653/v1/2020.acl-main.711
- Chakrabarty et al. (2023a) Tuhin Chakrabarty, Philippe Laban, Divyansh Agarwal, Smaranda Muresan, and Chien-Sheng Wu. 2023a. Art or Artifice? Large Language Models and the False Promise of Creativity. Proceedings of the CHI Conference on Human Factors in Computing Systems (2023). https://api.semanticscholar.org/CorpusID:262826094
- Chakrabarty et al. (2023b) Tuhin Chakrabarty, Philippe Laban, Divyansh Agarwal, Smaranda Muresan, and Chien-Sheng Wu. 2023b. Art or Artifice? Large Language Models and the False Promise of Creativity. arXiv:2309.14556 [cs.CL]
- Chakrabarty et al. (2020b) Tuhin Chakrabarty, Smaranda Muresan, and Nanyun Peng. 2020b. Generating similes effortlessly like a Pro: A Style Transfer Approach for Simile Generation. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), Bonnie Webber, Trevor Cohn, Yulan He, and Yang Liu (Eds.). Association for Computational Linguistics, Online, 6455–6469. https://doi.org/10.18653/v1/2020.emnlp-main.524
- Chakrabarty et al. (2023c) Tuhin Chakrabarty, Vishakh Padmakumar, Faeze Brahman, and Smaranda Muresan. 2023c. Creativity Support in the Age of Large Language Models: An Empirical Study Involving Emerging Writers. ArXiv abs/2309.12570 (2023). https://api.semanticscholar.org/CorpusID:262217523
- Chakrabarty et al. (2022a) Tuhin Chakrabarty, Vishakh Padmakumar, and Hengxing He. 2022a. Help me write a Poem - Instruction Tuning as a Vehicle for Collaborative Poetry Writing. ArXiv abs/2210.13669 (2022). https://api.semanticscholar.org/CorpusID:253107865
- Chakrabarty et al. (2022b) Tuhin Chakrabarty, Vishakh Padmakumar, and He He. 2022b. Help me write a Poem - Instruction Tuning as a Vehicle for Collaborative Poetry Writing. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, Yoav Goldberg, Zornitsa Kozareva, and Yue Zhang (Eds.). Association for Computational Linguistics, Abu Dhabi, United Arab Emirates, 6848–6863. https://doi.org/10.18653/v1/2022.emnlp-main.460
- Chakrabarty et al. (2023d) Tuhin Chakrabarty, Vishakh Padmakumar, He He, and Nanyun Peng. 2023d. Creative Natural Language Generation. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing: Tutorial Abstracts, Qi Zhang and Hassan Sajjad (Eds.). Association for Computational Linguistics, Singapore, 34–40. https://doi.org/10.18653/v1/2023.emnlp-tutorial.6
- Chakrabarty et al. (2022c) Tuhin Chakrabarty, Arkadiy Saakyan, Debanjan Ghosh, and Smaranda Muresan. 2022c. FLUTE: Figurative Language Understanding through Textual Explanations. In Conference on Empirical Methods in Natural Language Processing. https://api.semanticscholar.org/CorpusID:252780978
- Chakrabarty et al. (2023e) Tuhin Chakrabarty, Arkadiy Saakyan, Olivia Winn, Artemis Panagopoulou, Yue Yang, Marianna Apidianaki, and Smaranda Muresan. 2023e. I Spy a Metaphor: Large Language Models and Diffusion Models Co-Create Visual Metaphors. In Annual Meeting of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:258865878
- Chakrabarty et al. (2023f) Tuhin Chakrabarty, Arkadiy Saakyan, Olivia Winn, Artemis Panagopoulou, Yue Yang, Marianna Apidianaki, and Smaranda Muresan. 2023f. I Spy a Metaphor: Large Language Models and Diffusion Models Co-Create Visual Metaphors. In Findings of the Association for Computational Linguistics: ACL 2023, Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 7370–7388. https://doi.org/10.18653/v1/2023.findings-acl.465
- Chakrabarty et al. (2021b) Tuhin Chakrabarty, Xurui Zhang, Smaranda Muresan, and Nanyun Peng. 2021b. MERMAID: Metaphor Generation with Symbolism and Discriminative Decoding. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Kristina Toutanova, Anna Rumshisky, Luke Zettlemoyer, Dilek Hakkani-Tur, Iz Beltagy, Steven Bethard, Ryan Cotterell, Tanmoy Chakraborty, and Yichao Zhou (Eds.). Association for Computational Linguistics, Online, 4250–4261. https://doi.org/10.18653/v1/2021.naacl-main.336
- Chang et al. (2023) Kent K. Chang, Mackenzie Cramer, Sandeep Soni, and David Bamman. 2023. Speak, Memory: An Archaeology of Books Known to ChatGPT/GPT-4. arXiv:2305.00118 [cs.CL] https://arxiv.org/abs/2305.00118
- Chen et al. (2022a) Boyuan Chen, Kuang Huang, Sunand Raghupathi, Ishaan Preetam Chandratreya, Qi Du, and Hod Lipson. 2022a. Automated discovery of fundamental variables hidden in experimental data. Nature Computational Science 2 (2022), 433 – 442. https://api.semanticscholar.org/CorpusID:251087119
- Chen and Ding (2023) Honghua Chen and Nai Ding. 2023. Probing the Creativity of Large Language Models: Can models produce divergent semantic association?. In Conference on Empirical Methods in Natural Language Processing. https://api.semanticscholar.org/CorpusID:264172665
- Chen et al. (2022b) Hong Chen, Duc Minh Vo, Hiroya Takamura, Yusuke Miyao, and Hideki Nakayama. 2022b. StoryER: Automatic Story Evaluation via Ranking, Rating and Reasoning. ArXiv abs/2210.08459 (2022). https://api.semanticscholar.org/CorpusID:252918409
- Chen and Soo (2018) Peng-Yu Chen and V. Soo. 2018. Humor Recognition Using Deep Learning. In North American Chapter of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:44158782
- Chen et al. (2016) William YC Chen, Qing-Hu Hou, and Doron Zeilberger. 2016. Automated discovery and proof of congruence theorems for partial sums of combinatorial sequences. Journal of Difference Equations and Applications 22, 6 (2016), 780–788.
- Chhun et al. (2024) Cyril Chhun, Fabian M. Suchanek, and Chloé Clavel. 2024. Do Language Models Enjoy Their Own Stories? Prompting Large Language Models for Automatic Story Evaluation. Transactions of the Association for Computational Linguistics 12 (2024), 1122–1142. https://api.semanticscholar.org/CorpusID:269982125
- Chollet (2019) François Chollet. 2019. On the Measure of Intelligence. ArXiv abs/1911.01547 (2019). https://api.semanticscholar.org/CorpusID:207870692
- Chomsky (1965) Noam Chomsky. 1965. Aspects of the Theory of Syntax (50 ed.). The MIT Press. http://www.jstor.org/stable/j.ctt17kk81z
- Chung et al. (2022) John Joon Young Chung, Wooseok Kim, Kang Min Yoo, Hwaran Lee, Eytan Adar, and Minsuk Chang. 2022. TaleBrush: Sketching Stories with Generative Pretrained Language Models. Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (2022). https://api.semanticscholar.org/CorpusID:247625751
- Coil and Shwartz (2023) Albert Coil and Vered Shwartz. 2023. From chocolate bunny to chocolate crocodile: Do Language Models Understand Noun Compounds?. In Findings of the Association for Computational Linguistics: ACL 2023, Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 2698–2710. https://doi.org/10.18653/v1/2023.findings-acl.169
- Colton (2008) Simon Colton. 2008. Creativity Versus the Perception of Creativity in Computational Systems. In AAAI Spring Symposium: Creative Intelligent Systems. https://api.semanticscholar.org/CorpusID:2876679
- Colton et al. (2012) Simon Colton, Jacob Goodwin, and Tony Veale. 2012. Full-FACE Poetry Generation. In International Conference on Innovative Computing and Cloud Computing. https://api.semanticscholar.org/CorpusID:2218980
- Colton and Wiggins (2012) Simon Colton and Geraint A. Wiggins. 2012. Computational Creativity: The Final Frontier?. In European Conference on Artificial Intelligence. https://api.semanticscholar.org/CorpusID:5880786
- Conwell and Ullman (2022) Colin Conwell and Tomer David Ullman. 2022. Testing Relational Understanding in Text-Guided Image Generation. ArXiv abs/2208.00005 (2022). https://api.semanticscholar.org/CorpusID:251224307
- Cook and Colton (2015) Michael Cook and Simon Colton. 2015. Generating Code For Expressing Simple Preferences: Moving On From Hardcoding And Randomness.. In ICCC. 8–16.
- Cope (1996) David Cope. 1996. Experiments in musical intelligence. Vol. 12. AR editions Madison, WI.
- Copet et al. (2023) Jade Copet, Felix Kreuk, Itai Gat, Tal Remez, David Kant, Gabriel Synnaeve, Yossi Adi, and Alexandre D’efossez. 2023. Simple and Controllable Music Generation. ArXiv abs/2306.05284 (2023). https://api.semanticscholar.org/CorpusID:259108357
- Costello and Keane (2000) Fintan J Costello and Mark T Keane. 2000. Efficient creativity: Constraint-guided conceptual combination. Cognitive Science 24, 2 (2000), 299–349.
- Cranmer et al. (2020) M. Cranmer, Alvaro Sanchez-Gonzalez, Peter W. Battaglia, Rui Xu, Kyle Cranmer, David N. Spergel, and Shirley Ho. 2020. Discovering Symbolic Models from Deep Learning with Inductive Biases. ArXiv abs/2006.11287 (2020). https://api.semanticscholar.org/CorpusID:219966125
- Czinczoll et al. (2022) Tamara Czinczoll, Helen Yannakoudakis, Pushkar Mishra, and Ekaterina Shutova. 2022. Scientific and Creative Analogies in Pretrained Language Models. In Findings of the Association for Computational Linguistics: EMNLP 2022, Yoav Goldberg, Zornitsa Kozareva, and Yue Zhang (Eds.). Association for Computational Linguistics, Abu Dhabi, United Arab Emirates, 2094–2100. https://doi.org/10.18653/v1/2022.findings-emnlp.153
- Das and Ghosh (2017) Kollol Das and Shaona Ghosh. 2017. Neuramanteau: A Neural Network Ensemble Model for Lexical Blends. In Proceedings of the Eighth International Joint Conference on Natural Language Processing (Volume 1: Long Papers), Greg Kondrak and Taro Watanabe (Eds.). Asian Federation of Natural Language Processing, Taipei, Taiwan, 576–583. https://aclanthology.org/I17-1058
- Dathathri et al. (2019) Sumanth Dathathri, Andrea Madotto, Janice Lan, Jane Hung, Eric Frank, Piero Molino, Jason Yosinski, and Rosanne Liu. 2019. Plug and Play Language Models: A Simple Approach to Controlled Text Generation. ArXiv abs/1912.02164 (2019). https://api.semanticscholar.org/CorpusID:208617790
- Davidson et al. (2022) Guy Davidson, Todd M Gureckis, and Brenden M Lake. 2022. Creativity, Compositionality, and Common Sense in Human Goal Generation. https://doi.org/10.31234/osf.io/byzs5
- Deltorn and Macrez (2018) Jean-Marc Deltorn and Franck Macrez. 2018. Authorship in the Age of Machine learning and Artificial Intelligence. Legal Perspectives in Information Systems eJournal (2018). https://api.semanticscholar.org/CorpusID:69806575
- Deng et al. (2024) Qixin Deng, Qikai Yang, Ruibin Yuan, Yipeng Huang, Yi Wang, Xubo Liu, Zeyue Tian, Jiahao Pan, Ge Zhang, Hanfeng Lin, et al. 2024. ComposerX: Multi-Agent Symbolic Music Composition with LLMs. arXiv preprint arXiv:2404.18081 (2024).
- Denicola (2016) Robert C. Denicola. 2016. Ex Machina: Copyright Protection for Computer-Generated Works. Innovation Law & Policy eJournal (2016). https://api.semanticscholar.org/CorpusID:88486949
- Dhar and van der Plas (2019) Prajit Dhar and Lonneke van der Plas. 2019. Learning to Predict Novel Noun-Noun Compounds. arXiv:1906.03634 [cs.CL]
- Dhariwal et al. (2020) Prafulla Dhariwal, Heewoo Jun, Christine Payne, Jong Wook Kim, Alec Radford, and Ilya Sutskever. 2020. Jukebox: A Generative Model for Music. ArXiv abs/2005.00341 (2020). https://api.semanticscholar.org/CorpusID:218470180
- Donahue et al. (2019) Chris Donahue, Huanru Henry Mao, Yiting Li, G. Cottrell, and Julian McAuley. 2019. LakhNES: Improving Multi-instrumental Music Generation with Cross-domain Pre-training. ArXiv abs/1907.04868 (2019). https://api.semanticscholar.org/CorpusID:195886341
- Dong et al. (2014) Chao Dong, Chen Change Loy, Kaiming He, and Xiaoou Tang. 2014. Learning a Deep Convolutional Network for Image Super-Resolution. In European Conference on Computer Vision. https://api.semanticscholar.org/CorpusID:18874645
- Dong et al. (2017) Hao-Wen Dong, Wen-Yi Hsiao, Li-Chia Yang, and Yi-Hsuan Yang. 2017. MuseGAN: Multi-track Sequential Generative Adversarial Networks for Symbolic Music Generation and Accompaniment. In AAAI Conference on Artificial Intelligence. https://api.semanticscholar.org/CorpusID:19098155
- dos Santos and Machado (2020) Cláudio Lisboa dos Santos and Ângela Rocha Machado. 2020. Intellectual Property on Works of Art Made by Artificial Intelligence. International Journal of Advanced Engineering Research and Science (2020). https://api.semanticscholar.org/CorpusID:234541178
- Du and Chilton (2023) Yulun Du and Lydia Chilton. 2023. StoryWars: A Dataset and Instruction Tuning Baselines for Collaborative Story Understanding and Generation. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 3044–3062. https://doi.org/10.18653/v1/2023.acl-long.171
- Duarte et al. (2024) André V. Duarte, Xuandong Zhao, Arlindo L. Oliveira, and Lei Li. 2024. DE-COP: Detecting Copyrighted Content in Language Models Training Data. arXiv:2402.09910 [cs.CL] https://arxiv.org/abs/2402.09910
- Dumoulin et al. (2016) Vincent Dumoulin, Jonathon Shlens, and Manjunath Kudlur. 2016. A Learned Representation For Artistic Style. ArXiv abs/1610.07629 (2016). https://api.semanticscholar.org/CorpusID:5687613
- Dunbar and Forster (2009) Kevin Dunbar and Eve Forster. 2009. Creativity Evaluation through Latent Semantic Analysis. https://api.semanticscholar.org/CorpusID:18209155
- Duncker and Lees (1948) K. Duncker and L.S. Lees. 1948. On Problem-solving. American Psychological Ass. https://books.google.ch/books?id=g888tAEACAAJ
- Dzeroski and Todorovski (1993) Saso Dzeroski and Ljupco Todorovski. 1993. Discovering Dynamics. In International Conference on Machine Learning. https://api.semanticscholar.org/CorpusID:17202248
- Dziri et al. (2023) Nouha Dziri, Ximing Lu, Melanie Sclar, Xiang Lorraine Li, Liwei Jian, Bill Yuchen Lin, Peter West, Chandra Bhagavatula, Ronan Le Bras, Jena D. Hwang, Soumya Sanyal, Sean Welleck, Xiang Ren, Allyson Ettinger, Zaid Harchaoui, and Yejin Choi. 2023. Faith and Fate: Limits of Transformers on Compositionality. ArXiv abs/2305.18654 (2023). https://api.semanticscholar.org/CorpusID:258967391
- Eck and Schmidhuber (2002) Douglas Eck and Juergen Schmidhuber. 2002. Finding temporal structure in music: Blues improvisation with LSTM recurrent networks. In Proceedings of the 12th IEEE workshop on neural networks for signal processing. IEEE, 747–756.
- Elazar et al. (2021) Yanai Elazar, Nora Kassner, Shauli Ravfogel, Abhilasha Ravichander, Eduard H. Hovy, Hinrich Schütze, and Yoav Goldberg. 2021. Measuring and Improving Consistency in Pretrained Language Models. Transactions of the Association for Computational Linguistics 9 (2021), 1012–1031. https://api.semanticscholar.org/CorpusID:231740560
- Elgammal et al. (2017) Ahmed Elgammal, Bingchen Liu, Mohamed Elhoseiny, and Marian Mazzone. 2017. CAN: Creative Adversarial Networks, Generating ”Art” by Learning About Styles and Deviating from Style Norms. arXiv:1706.07068 [cs.AI]
- Elgammal and Saleh (2015) A. Elgammal and Babak Saleh. 2015. Quantifying Creativity in Art Networks. ArXiv abs/1506.00711 (2015). https://api.semanticscholar.org/CorpusID:14607648
- Elzohbi and Zhao (2023) Mohamad Elzohbi and Richard Zhao. 2023. Creative Data Generation: A Review Focusing on Text and Poetry. ArXiv abs/2305.08493 (2023). https://api.semanticscholar.org/CorpusID:258686736
- Fajtlowicz (1988) Siemion Fajtlowicz. 1988. On Conjectures of Graffiti. In Graph Theory and Applications, J. Akiyama, Y. Egawa, and H. Enomoto (Eds.). Annals of Discrete Mathematics, Vol. 38. Elsevier, 113–118. https://doi.org/10.1016/S0167-5060(08)70776-3
- Falkenhainer et al. (1989) Brian Falkenhainer, Kenneth D. Forbus, and Dedre Gentner. 1989. The Structure-Mapping Engine: Algorithm and Examples. Artif. Intell. 41 (1989), 1–63. https://api.semanticscholar.org/CorpusID:8751960
- Fan et al. (2018) Angela Fan, Mike Lewis, and Yann Dauphin. 2018. Hierarchical Neural Story Generation. In Annual Meeting of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:44134226
- Farbood and Schöner (2001) Mary Farbood and Bernd Schöner. 2001. Analysis and synthesis of Palestrina-style counterpoint using Markov chains. In ICMC.
- Franceschelli and Musolesi (2021a) Giorgio Franceschelli and Mirco Musolesi. 2021a. Copyright in generative deep learning. Data & Policy 4 (2021). https://api.semanticscholar.org/CorpusID:234777751
- Franceschelli and Musolesi (2021b) Giorgio Franceschelli and Mirco Musolesi. 2021b. Creativity and Machine Learning: A Survey. ArXiv abs/2104.02726 (2021). https://api.semanticscholar.org/CorpusID:233168627
- Franceschelli and Musolesi (2022) Giorgio Franceschelli and Mirco Musolesi. 2022. DeepCreativity: Measuring Creativity with Deep Learning Techniques. Intelligenza Artificiale 16 (2022), 151–163. https://api.semanticscholar.org/CorpusID:246015719
- Franceschelli and Musolesi (2023) Giorgio Franceschelli and Mirco Musolesi. 2023. On the Creativity of Large Language Models. arXiv:2304.00008 [cs.AI] https://arxiv.org/abs/2304.00008
- Franceschelli and Musolesi (2024) Giorgio Franceschelli and Mirco Musolesi. 2024. Creative Beam Search: LLM-as-a-Judge For Improving Response Generation. ArXiv abs/2405.00099 (2024). https://api.semanticscholar.org/CorpusID:269484318
- Frank (2010) S. Frank. 2010. Uncertainty Reduction as a Measure of Cognitive Processing Effort. In CMCL@ACL. https://api.semanticscholar.org/CorpusID:3097133
- França et al. (2016) Celso França, Luís Fabrício Wanderley Góes, Alvaro Amorim, Rodrigo C. O. Rocha, and Alysson Ribeiro Da Silva. 2016. Regent-Dependent Creativity: A Domain Independent Metric for the Assessment of Creative Artifacts. In International Conference on Innovative Computing and Cloud Computing. https://api.semanticscholar.org/CorpusID:7817742
- Gafni et al. (2022) Oran Gafni, Adam Polyak, Oron Ashual, Shelly Sheynin, Devi Parikh, and Yaniv Taigman. 2022. Make-A-Scene: Scene-Based Text-to-Image Generation with Human Priors. ArXiv abs/2203.13131 (2022). https://api.semanticscholar.org/CorpusID:247628171
- Garcon et al. (2021) Antoine Garcon, Julian Vexler, Dmitry Budker, and Stefan Kramer. 2021. Deep neural networks to recover unknown physical parameters from oscillating time series. PLoS ONE 17 (2021). https://api.semanticscholar.org/CorpusID:231572959
- Gatys et al. (2015) Leon A. Gatys, Alexander S. Ecker, and Matthias Bethge. 2015. A Neural Algorithm of Artistic Style. ArXiv abs/1508.06576 (2015). https://api.semanticscholar.org/CorpusID:13914930
- Gaut (2010) Berys Gaut. 2010. The Philosophy of Creativity. Philosophy Compass 5, 12 (2010), 1034–1046. https://doi.org/10.1111/j.1747-9991.2010.00351.x arXiv:https://compass.onlinelibrary.wiley.com/doi/pdf/10.1111/j.1747-9991.2010.00351.x
- Gemini (2024) Team Gemini. 2024. Gemini: A Family of Highly Capable Multimodal Models. arXiv:2312.11805 [cs.CL]
- Gendron et al. (2023) Gaël Gendron, Qiming Bao, M. Witbrock, and Gillian Dobbie. 2023. Large Language Models Are Not Strong Abstract Reasoners. https://api.semanticscholar.org/CorpusID:258988045
- Geng et al. (2023) Zigang Geng, Binxin Yang, Tiankai Hang, Chen Li, Shuyang Gu, Ting Zhang, Jianmin Bao, Zheng Zhang, Han Hu, Dongdong Chen, and Baining Guo. 2023. InstructDiffusion: A Generalist Modeling Interface for Vision Tasks. 2024 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2023), 12709–12720. https://api.semanticscholar.org/CorpusID:261582721
- Gentner (1983) Dedre Gentner. 1983. Structure-Mapping: A Theoretical Framework for Analogy. Cogn. Sci. 7 (1983), 155–170. https://api.semanticscholar.org/CorpusID:5371492
- Gervás (2019) Pablo Gervás. 2019. Exploring Quantitative Evaluations of the Creativity of Automatic Poets. 275–304. https://doi.org/10.1007/978-3-319-43610-4_13
- Ghafarollahi and Buehler (2024) Alireza Ghafarollahi and Markus J. Buehler. 2024. SciAgents: Automating scientific discovery through multi-agent intelligent graph reasoning. https://api.semanticscholar.org/CorpusID:272524655
- Ghazvininejad et al. (2016) Marjan Ghazvininejad, Xing Shi, Yejin Choi, and Kevin Knight. 2016. Generating Topical Poetry. In Proceedings of the 2016 Conference on Empirical Methods in Natural Language Processing, Jian Su, Kevin Duh, and Xavier Carreras (Eds.). Association for Computational Linguistics, Austin, Texas, 1183–1191. https://doi.org/10.18653/v1/D16-1126
- Ginsburg (2018) Jane C. Ginsburg. 2018. People Not Machines: Authorship and What It Means in the Berne Convention. IIC - International Review of Intellectual Property and Competition Law 49 (2018), 131–135. https://api.semanticscholar.org/CorpusID:158580780
- Girotra et al. (2023) Karan Girotra, Lennart Meincke, Christian Terwiesch, and Karl T. Ulrich. 2023. Ideas are Dimes a Dozen: Large Language Models for Idea Generation in Innovation. SSRN Electronic Journal (2023). https://api.semanticscholar.org/CorpusID:260467886
- Gladkova et al. (2016) Anna Gladkova, Aleksandr Drozd, and Satoshi Matsuoka. 2016. Analogy-based detection of morphological and semantic relations with word embeddings: what works and what doesn’t.. In Proceedings of the NAACL Student Research Workshop, Jacob Andreas, Eunsol Choi, and Angeliki Lazaridou (Eds.). Association for Computational Linguistics, San Diego, California, 8–15. https://doi.org/10.18653/v1/N16-2002
- Góes et al. (2023a) Fabrício Góes, Piotr Sawicki, Marek Grze´s, Daniel Brown, and Marco Volpe. 2023a. Is GPT-4 Good Enough to Evaluate Jokes? https://api.semanticscholar.org/CorpusID:265255804
- Góes et al. (2023b) Fabrício Góes, Marco Volpe, Piotr Sawicki, Marek Grze´s, and Jacob Watson. 2023b. Pushing GPT’s Creativity to Its Limits: Alternative Uses and Torrance Tests. https://api.semanticscholar.org/CorpusID:260844252
- Goldfarb-Tarrant et al. (2020) Seraphina Goldfarb-Tarrant, Tuhin Chakrabarty, Ralph Weischedel, and Nanyun Peng. 2020. Content Planning for Neural Story Generation with Aristotelian Rescoring. In the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). 4319–4338.
- Good (1965) I. J. Good. 1965. Speculations Concerning the First Ultraintelligent Machine. Adv. Comput. 6 (1965), 31–88. https://api.semanticscholar.org/CorpusID:17886872
- Goodfellow et al. (2014a) Ian J. Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron Courville, and Yoshua Bengio. 2014a. Generative Adversarial Networks. arXiv:1406.2661 [stat.ML]
- Goodfellow et al. (2014b) Ian J. Goodfellow, Jean Pouget-Abadie, Mehdi Mirza, Bing Xu, David Warde-Farley, Sherjil Ozair, Aaron C. Courville, and Yoshua Bengio. 2014b. Generative adversarial networks. Commun. ACM 63 (2014), 139 – 144. https://api.semanticscholar.org/CorpusID:1033682
- Grynbaum and Mac (2023) M. M. Grynbaum and R. Mac. 2023. The Times Sues OpenAI and Microsoft Over A.I. Use of Copyrighted Work. https://www.nytimes.com/2023/12/27/business/media/new-york-times-open-ai-microsoft-lawsuit.html
- Guadamuz (2017) Andres Guadamuz. 2017. Do Androids Dream of Electric Copyright? Comparative Analysis of Originality in Artificial Intelligence Generated Works. Econometrics: Computer Programs & Software eJournal (2017). https://api.semanticscholar.org/CorpusID:21670175
- Guan and Huang (2020) Jian Guan and Minlie Huang. 2020. UNION: An Unreferenced Metric for Evaluating Open-ended Story Generation. ArXiv abs/2009.07602 (2020). https://api.semanticscholar.org/CorpusID:221739231
- Guilford (1967) J.P. Guilford. 1967. The Nature of Human Intelligence. McGraw-Hill. https://books.google.ch/books?id=T-ZJAAAAMAAJ
- Guimaraes et al. (2017) Gabriel Lima Guimaraes, Benjamín Sánchez-Lengeling, Pedro Luis Cunha Farias, and Alán Aspuru-Guzik. 2017. Objective-Reinforced Generative Adversarial Networks (ORGAN) for Sequence Generation Models. ArXiv abs/1705.10843 (2017). https://api.semanticscholar.org/CorpusID:35911567
- Guimerà et al. (2020) Roger Guimerà, Ignasi Reichardt, Antoni Aguilar-Mogas, Francesco A. Massucci, Manuel Miranda, Jordi Pallarès, and Marta Sales-Pardo. 2020. A Bayesian machine scientist to aid in the solution of challenging scientific problems. Science Advances 6, 5 (2020), eaav6971. https://doi.org/10.1126/sciadv.aav6971 arXiv:https://www.science.org/doi/pdf/10.1126/sciadv.aav6971
- Gunkle and Berlyne (1975) George Gunkle and Daniel E. Berlyne. 1975. Aesthetics and Psychobiology. https://api.semanticscholar.org/CorpusID:144314694
- Gupta et al. (2023) Agrim Gupta, Lijun Yu, Kihyuk Sohn, Xiuye Gu, Meera Hahn, Fei-Fei Li, Irfan Essa, Lu Jiang, and José Lezama. 2023. Photorealistic Video Generation with Diffusion Models. ArXiv abs/2312.06662 (2023). https://api.semanticscholar.org/CorpusID:266163109
- Guzik et al. (2023) Erik E. Guzik, Christian Byrge, and Christian Gilde. 2023. The Originality of Machines: AI Takes the Torrance Test. Journal of Creativity (2023). https://api.semanticscholar.org/CorpusID:261087185
- Hakuk and Reich (2020) Yaron Hakuk and Yoram Reich. 2020. Automated discovery of scientific concepts: Replicating three recent discoveries in mechanics. Advanced Engineering Informatics 44 (2020), 101080. https://doi.org/10.1016/j.aei.2020.101080
- Hale (2006) John Hale. 2006. Uncertainty About the Rest of the Sentence. Cognitive science 30 4 (2006), 643–72. https://api.semanticscholar.org/CorpusID:15922633
- Hans et al. (2024) Abhimanyu Hans, Yuxin Wen, Neel Jain, John Kirchenbauer, Hamid Kazemi, Prajwal Singhania, Siddharth Singh, Gowthami Somepalli, Jonas Geiping, Abhinav Bhatele, and Tom Goldstein. 2024. Be like a Goldfish, Don’t Memorize! Mitigating Memorization in Generative LLMs. ArXiv abs/2406.10209 (2024). https://api.semanticscholar.org/CorpusID:270521774
- Harbison and Haarmann (2014) J. Isaiah Harbison and Henk J. Haarmann. 2014. Automated scoring of originality using semantic representations. Cognitive Science 36 (2014). https://api.semanticscholar.org/CorpusID:2645346
- Hasan et al. (2019) Md Kamrul Hasan, Wasifur Rahman, AmirAli Bagher Zadeh, Jianyuan Zhong, Md Iftekhar Tanveer, Louis-Philippe Morency, and Mohammed (Ehsan) Hoque. 2019. UR-FUNNY: A Multimodal Language Dataset for Understanding Humor. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), Kentaro Inui, Jing Jiang, Vincent Ng, and Xiaojun Wan (Eds.). Association for Computational Linguistics, Hong Kong, China, 2046–2056. https://doi.org/10.18653/v1/D19-1211
- Hassabis (2018) Demis Hassabis. 2018. Creativity and AI. https://www.youtube.com/watch?v=d-bvsJWmqlc&ab_channel=RoyalAcademyofArts
- He et al. (2019) He He, Nanyun Peng, and Percy Liang. 2019. Pun Generation with Surprise. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), Jill Burstein, Christy Doran, and Thamar Solorio (Eds.). Association for Computational Linguistics, Minneapolis, Minnesota, 1734–1744. https://doi.org/10.18653/v1/N19-1172
- He et al. (2023) Qianyu He, Yikai Zhang, Jiaqing Liang, Yuncheng Huang, Yanghua Xiao, and Yunwen Chen. 2023. HAUSER: Towards Holistic and Automatic Evaluation of Simile Generation. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 12557–12572. https://doi.org/10.18653/v1/2023.acl-long.702
- Heinen and Johnson (2017) David Heinen and Dan Johnson. 2017. Semantic Distance: An Automated Measure of Creativity That Is Novel and Appropriate. Psychology of Aesthetics, Creativity, and the Arts 12 (04 2017). https://doi.org/10.1037/aca0000125
- Hendrickx et al. (2013) Iris Hendrickx, Zornitsa Kozareva, Preslav Nakov, Diarmuid Ó Séaghdha, Stan Szpakowicz, and Tony Veale. 2013. SemEval-2013 Task 4: Free Paraphrases of Noun Compounds. In Second Joint Conference on Lexical and Computational Semantics (*SEM), Volume 2: Proceedings of the Seventh International Workshop on Semantic Evaluation (SemEval 2013), Suresh Manandhar and Deniz Yuret (Eds.). Association for Computational Linguistics, Atlanta, Georgia, USA, 138–143. https://aclanthology.org/S13-2025
- Hertz et al. (2022) Amir Hertz, Ron Mokady, Jay M. Tenenbaum, Kfir Aberman, Yael Pritch, and Daniel Cohen-Or. 2022. Prompt-to-Prompt Image Editing with Cross Attention Control. ArXiv abs/2208.01626 (2022). https://api.semanticscholar.org/CorpusID:251252882
- Hessel et al. (2023a) Jack Hessel, Ana Marasovic, Jena D. Hwang, Lillian Lee, Jeff Da, Rowan Zellers, Robert Mankoff, and Yejin Choi. 2023a. Do Androids Laugh at Electric Sheep? Humor “Understanding” Benchmarks from The New Yorker Caption Contest. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Anna Rogers, Jordan Boyd-Graber, and Naoaki Okazaki (Eds.). Association for Computational Linguistics, Toronto, Canada, 688–714. https://doi.org/10.18653/v1/2023.acl-long.41
- Hessel et al. (2023b) Jack Hessel, Ana Marasović, Jena D. Hwang, Lillian Lee, Jeff Da, Rowan Zellers, Robert Mankoff, and Yejin Choi. 2023b. Do Androids Laugh at Electric Sheep? Humor ”Understanding” Benchmarks from The New Yorker Caption Contest. arXiv:2209.06293
- Hiller and Isaacson (1958) Lejaren Hiller and Leonard Isaacson. 1958. Musical composition with a high-speed digital computer. (1958).
- Ho et al. (2022a) Jonathan Ho, William Chan, Chitwan Saharia, Jay Whang, Ruiqi Gao, Alexey A. Gritsenko, Diederik P. Kingma, Ben Poole, Mohammad Norouzi, David J. Fleet, and Tim Salimans. 2022a. Imagen Video: High Definition Video Generation with Diffusion Models. ArXiv abs/2210.02303 (2022). https://api.semanticscholar.org/CorpusID:252715883
- Ho et al. (2020) Jonathan Ho, Ajay Jain, and Pieter Abbeel. 2020. Denoising Diffusion Probabilistic Models. arXiv:2006.11239
- Ho et al. (2022b) Jonathan Ho, Tim Salimans, Alexey Gritsenko, William Chan, Mohammad Norouzi, and David J. Fleet. 2022b. Video Diffusion Models. ArXiv abs/2204.03458 (2022). https://api.semanticscholar.org/CorpusID:248006185
- Hofstadter (2001) Douglas Hofstadter. 2001. Analogy as the Core of Cognition. MIT Press (2001).
- Holtzman et al. (2019) Ari Holtzman, Jan Buys, Li Du, Maxwell Forbes, and Yejin Choi. 2019. The Curious Case of Neural Text Degeneration. ArXiv abs/1904.09751 (2019). https://api.semanticscholar.org/CorpusID:127986954
- Holzinger et al. (2019) Andreas Holzinger, Michael D. Kickmeier-Rust, and Heimo Müller. 2019. KANDINSKY Patterns as IQ-Test for Machine Learning. In International Cross-Domain Conference on Machine Learning and Knowledge Extraction. https://api.semanticscholar.org/CorpusID:201616604
- Hong et al. (2023) Xudong Hong, Asad Sayeed, Khushboo Mehra, Vera Demberg, and Bernt Schiele. 2023. Visual Writing Prompts: Character-Grounded Story Generation with Curated Image Sequences. Transactions of the Association for Computational Linguistics 11 (2023), 565–581. https://doi.org/10.1162/tacl_a_00553
- Horvitz et al. (2024) Zachary Horvitz, Jingru Chen, Rahul Aditya, Harshvardhan Srivastava, Robert West, Zhou Yu, and Kathleen McKeown. 2024. Getting Serious about Humor: Crafting Humor Datasets with Unfunny Large Language Models. ArXiv abs/2403.00794 (2024). https://api.semanticscholar.org/CorpusID:268230695
- Hossain et al. (2019) Nabil Hossain, John Krumm, and Michael Gamon. 2019. “President Vows to Cut <Taxes> Hair”: Dataset and Analysis of Creative Text Editing for Humorous Headlines. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), Jill Burstein, Christy Doran, and Thamar Solorio (Eds.). Association for Computational Linguistics, Minneapolis, Minnesota, 133–142. https://doi.org/10.18653/v1/N19-1012
- Hossain et al. (2020a) Nabil Hossain, John Krumm, Michael Gamon, and Henry Kautz. 2020a. SemEval-2020 Task 7: Assessing Humor in Edited News Headlines. In Proceedings of the Fourteenth Workshop on Semantic Evaluation, Aurelie Herbelot, Xiaodan Zhu, Alexis Palmer, Nathan Schneider, Jonathan May, and Ekaterina Shutova (Eds.). International Committee for Computational Linguistics, Barcelona (online), 746–758. https://doi.org/10.18653/v1/2020.semeval-1.98
- Hossain et al. (2020b) Nabil Hossain, John Krumm, Tanvir Sajed, and Henry Kautz. 2020b. Stimulating Creativity with FunLines: A Case Study of Humor Generation in Headlines. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics: System Demonstrations, Asli Celikyilmaz and Tsung-Hsien Wen (Eds.). Association for Computational Linguistics, Online, 256–262. https://doi.org/10.18653/v1/2020.acl-demos.28
- Hu et al. (2023) Xiaoyang Hu, Shane Storks, Richard L. Lewis, and Joyce Yue Chai. 2023. In-Context Analogical Reasoning with Pre-Trained Language Models. In Annual Meeting of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:258959097
- Huang et al. (2018b) Cheng-Zhi Anna Huang, Ashish Vaswani, Jakob Uszkoreit, Noam M. Shazeer, Ian Simon, Curtis Hawthorne, Andrew M. Dai, Matthew D. Hoffman, Monica Dinculescu, and Douglas Eck. 2018b. Music Transformer: Generating Music with Long-Term Structure. In International Conference on Learning Representations. https://api.semanticscholar.org/CorpusID:54477714
- Huang et al. (2023) Kaiyi Huang, Kaiyue Sun, Enze Xie, Zhenguo Li, and Xihui Liu. 2023. T2I-CompBench: A Comprehensive Benchmark for Open-world Compositional Text-to-image Generation. ArXiv abs/2307.06350 (2023). https://api.semanticscholar.org/CorpusID:259847295
- Huang et al. (2024a) Nisha Huang, Weiming Dong, Yuxin Zhang, Fan Tang, Ronghui Li, Chongyang Ma, Xiu Li, and Changsheng Xu. 2024a. CreativeSynth: Creative Blending and Synthesis of Visual Arts based on Multimodal Diffusion. ArXiv abs/2401.14066 (2024). https://api.semanticscholar.org/CorpusID:267211817
- Huang et al. (2022) Nisha Huang, Fan Tang, Weiming Dong, and Changsheng Xu. 2022. Draw Your Art Dream: Diverse Digital Art Synthesis with Multimodal Guided Diffusion. Proceedings of the 30th ACM International Conference on Multimedia (2022). https://api.semanticscholar.org/CorpusID:252544936
- Huang et al. (2024b) Qian Huang, Jian Vora, Percy Liang, and Jure Leskovec. 2024b. MLAgentBench: Evaluating Language Agents on Machine Learning Experimentation. arXiv:2310.03302 [cs.LG] https://arxiv.org/abs/2310.03302
- Huang et al. (2018a) Xun Huang, Ming-Yu Liu, Serge J. Belongie, and Jan Kautz. 2018a. Multimodal Unsupervised Image-to-Image Translation. In European Conference on Computer Vision. https://api.semanticscholar.org/CorpusID:4883312
- Huang and Yang (2020) Yu-Siang Huang and Yi-Hsuan Yang. 2020. Pop Music Transformer: Beat-based Modeling and Generation of Expressive Pop Piano Compositions. Proceedings of the 28th ACM International Conference on Multimedia (2020). https://api.semanticscholar.org/CorpusID:220919638
- Hubert et al. (2024a) Kent F Hubert, Kim N. Awa, and Darya L. Zabelina. 2024a. The current state of artificial intelligence generative language models is more creative than humans on divergent thinking tasks. Scientific Reports 14 (2024). https://api.semanticscholar.org/CorpusID:267616181
- Hubert et al. (2024b) Thomas Hubert, Rishi Mehta, and Laurent Sartran. 2024b. AlphaProof. https://deepmind.google/discover/blog/ai-solves-imo-problems-at-silver-medal-level/
- Hupkes et al. (2022) Dieuwke Hupkes, Mario Giulianelli, Verna Dankers, Mikel Artetxe, Yanai Elazar, Tiago Pimentel, Christos Christodoulopoulos, Karim Lasri, Naomi Saphra, Arabella J. Sinclair, Dennis Ulmer, Florian Schottmann, Khuyagbaatar Batsuren, Kaiser Sun, Koustuv Sinha, Leila Khalatbari, Maria Ryskina, Rita Frieske, Ryan Cotterell, and Zhijing Jin. 2022. A taxonomy and review of generalization research in NLP. Nature Machine Intelligence 5 (2022), 1161 – 1174. https://api.semanticscholar.org/CorpusID:252735124
- Ifargan et al. (2024) Tal Ifargan, Lukas Hafner, Maor Kern, Ori Alcalay, and Roy Kishony. 2024. Autonomous LLM-driven research from data to human-verifiable research papers. arXiv:2404.17605 [q-bio.OT] https://arxiv.org/abs/2404.17605
- Ippolito et al. (2022) Daphne Ippolito, Florian Tramèr, Milad Nasr, Chiyuan Zhang, Matthew Jagielski, Katherine Lee, Christopher A. Choquette-Choo, and Nicholas Carlini. 2022. Preventing Verbatim Memorization in Language Models Gives a False Sense of Privacy. ArXiv abs/2210.17546 (2022). https://api.semanticscholar.org/CorpusID:253237404
- Ismayilzada et al. (2024) Mete Ismayilzada, Defne Circi, Jonne Sälevä, Hale Sirin, Abdullatif Köksal, Bhuwan Dhingra, Antoine Bosselut, Lonneke van der Plas, and Duygu Ataman. 2024. Evaluating Morphological Compositional Generalization in Large Language Models. arXiv:2410.12656 [cs.CL] https://arxiv.org/abs/2410.12656
- Ismayilzada et al. (2023) Mete Ismayilzada, Debjit Paul, Syrielle Montariol, Mor Geva, and Antoine Bosselut. 2023. CRoW: Benchmarking Commonsense Reasoning in Real-World Tasks. ArXiv abs/2310.15239 (2023). https://api.semanticscholar.org/CorpusID:264438931
- Isola et al. (2016) Phillip Isola, Jun-Yan Zhu, Tinghui Zhou, and Alexei A. Efros. 2016. Image-to-Image Translation with Conditional Adversarial Networks. 2017 IEEE Conference on Computer Vision and Pattern Recognition (CVPR) (2016), 5967–5976. https://api.semanticscholar.org/CorpusID:6200260
- Iten et al. (2020) Raban Iten, Tony Metger, Henrik Wilming, Lídia del Rio, and Renato Renner. 2020. Discovering Physical Concepts with Neural Networks. Phys. Rev. Lett. 124 (Jan 2020), 010508. Issue 1. https://doi.org/10.1103/PhysRevLett.124.010508
- Jacob et al. (2023) Shahar Jacob, Chen Shani, and Dafna Shahaf. 2023. FAME: Flexible, Scalable Analogy Mappings Engine. In Conference on Empirical Methods in Natural Language Processing. https://api.semanticscholar.org/CorpusID:259267034
- Jacobs (2018) Arthur M. Jacobs. 2018. The Gutenberg English Poetry Corpus: Exemplary Quantitative Narrative Analyses. Frontiers in Digital Humanities 5 (2018). https://doi.org/10.3389/fdigh.2018.00005
- Jain et al. (2024) Veedant Jain, Felipe dos Santos Alves Feitosa, and Gabriel Kreiman. 2024. Is AI fun? HumorDB: a curated dataset and benchmark to investigate graphical humor. https://api.semanticscholar.org/CorpusID:270619380
- Jentzsch and Kersting (2023) Sophie Jentzsch and Kristian Kersting. 2023. ChatGPT is fun, but it is not funny! Humor is still challenging Large Language Models. In Proceedings of the 13th Workshop on Computational Approaches to Subjectivity, Sentiment, & Social Media Analysis, Jeremy Barnes, Orphée De Clercq, and Roman Klinger (Eds.). Association for Computational Linguistics, Toronto, Canada, 325–340. https://doi.org/10.18653/v1/2023.wassa-1.29
- Ji et al. (2020) Shulei Ji, Jing Luo, and Xinyu Yang. 2020. A Comprehensive Survey on Deep Music Generation: Multi-level Representations, Algorithms, Evaluations, and Future Directions. ArXiv abs/2011.06801 (2020). https://api.semanticscholar.org/CorpusID:226956064
- Ji et al. (2023) Shulei Ji, Xinyu Yang, and Jing Luo. 2023. A Survey on Deep Learning for Symbolic Music Generation: Representations, Algorithms, Evaluations, and Challenges. ACM Comput. Surv. 56, 1, Article 7 (aug 2023), 39 pages. https://doi.org/10.1145/3597493
- Jiang et al. (2023) Yifan Jiang, Filip Ilievski, and Kaixin Ma. 2023. BRAINTEASER: Lateral Thinking Puzzles for Large Language Models. In Conference on Empirical Methods in Natural Language Processing. https://api.semanticscholar.org/CorpusID:263830212
- Jiayang et al. (2023) Cheng Jiayang, Lin Qiu, Tsz Chan, Tianqing Fang, Weiqi Wang, Chunkit Chan, Dongyu Ru, Qipeng Guo, Hongming Zhang, Yangqiu Song, Yue Zhang, and Zheng Zhang. 2023. StoryAnalogy: Deriving Story-level Analogies from Large Language Models to Unlock Analogical Understanding. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 11518–11537. https://doi.org/10.18653/v1/2023.emnlp-main.706
- Johnson et al. (2022) Dan Richard Johnson, J. Kaufman, Brendan S. Baker, John D. Patterson, Baptiste Barbot, Adam E. Green, Janet G. van Hell, Evan S. Kennedy, Grace F Sullivan, Christa L. Taylor, Thomas Ward, and Roger E. Beaty. 2022. Divergent semantic integration (DSI): Extracting creativity from narratives with distributional semantic modeling. Behavior Research Methods 55 (2022), 3726 – 3759. https://api.semanticscholar.org/CorpusID:252969336
- Johnson et al. (2016) Justin Johnson, Alexandre Alahi, and Li Fei-Fei. 2016. Perceptual Losses for Real-Time Style Transfer and Super-Resolution. ArXiv abs/1603.08155 (2016). https://api.semanticscholar.org/CorpusID:980236
- Jordanous (2012) Anna Jordanous. 2012. A Standardised Procedure for Evaluating Creative Systems: Computational Creativity Evaluation Based on What it is to be Creative. Cognitive Computation 4 (2012), 246 – 279. https://api.semanticscholar.org/CorpusID:311189
- Jordanous (2016) Anna Jordanous. 2016. The longer term value of creativity judgements in computational creativity. https://api.semanticscholar.org/CorpusID:637189
- Jordanous et al. (2015) Anna Jordanous, Daniel Allington, and Byron Dueck. 2015. Measuring cultural value using social network analysis: a case study on valuing electronic musicians. (2015).
- Jumper et al. (2021) John Jumper, Richard Evans, Alexander Pritzel, Tim Green, Michael Figurnov, Olaf Ronneberger, Kathryn Tunyasuvunakool, Russ Bates, Augustin Žídek, Anna Potapenko, Alex Bridgland, Clemens Meyer, Simon A. A. Kohl, Andrew J. Ballard, Andrew Cowie, Bernardino Romera-Paredes, Stanislav Nikolov, Rishub Jain, Jonas Adler, Trevor Back, Stig Petersen, David Reiman, Ellen Clancy, Michal Zielinski, Martin Steinegger, Michalina Pacholska, Tamas Berghammer, Sebastian Bodenstein, David Silver, Oriol Vinyals, Andrew W. Senior, Koray Kavukcuoglu, Pushmeet Kohli, and Demis Hassabis. 2021. Highly accurate protein structure prediction with AlphaFold. Nature 596, 7873 (01 Aug 2021), 583–589. https://doi.org/10.1038/s41586-021-03819-2
- Kandpal et al. (2022) Nikhil Kandpal, Eric Wallace, and Colin Raffel. 2022. Deduplicating Training Data Mitigates Privacy Risks in Language Models. ArXiv abs/2202.06539 (2022). https://api.semanticscholar.org/CorpusID:246823128
- Karampiperis et al. (2014) Pythagoras Karampiperis, Antonis Koukourikos, and Evangelia Koliopoulou. 2014. Towards Machines for Measuring Creativity: The Use of Computational Tools in Storytelling Activities. 2014 IEEE 14th International Conference on Advanced Learning Technologies (2014), 508–512. https://api.semanticscholar.org/CorpusID:17903634
- Karras et al. (2018) Tero Karras, Samuli Laine, and Timo Aila. 2018. A Style-Based Generator Architecture for Generative Adversarial Networks. 2019 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2018), 4396–4405. https://api.semanticscholar.org/CorpusID:54482423
- Karras et al. (2019) Tero Karras, Samuli Laine, Miika Aittala, Janne Hellsten, Jaakko Lehtinen, and Timo Aila. 2019. Analyzing and Improving the Image Quality of StyleGAN. 2020 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2019), 8107–8116. https://api.semanticscholar.org/CorpusID:209202273
- Kaufman and Beghetto (2009) James C. Kaufman and Ronald A. Beghetto. 2009. Beyond Big and Little: The Four C Model of Creativity. Review of General Psychology 13, 1 (2009), 1–12. https://doi.org/10.1037/a0013688 arXiv:https://doi.org/10.1037/a0013688
- Kayatani et al. (2021) Yuta Kayatani, Zekun Yang, Mayu Otani, Noa Garcia, Chenhui Chu, Yuta Nakashima, and Haruo Takemura. 2021. The Laughing Machine: Predicting Humor in Video. 2072–2081. https://doi.org/10.1109/WACV48630.2021.00212
- Kim (2006) Kyung H Kim. 2006. Can We Trust Creativity Tests? A Review of the Torrance Tests of Creative Thinking (TTCT). Creativity Research Journal 18 (2006), 14 – 3. https://api.semanticscholar.org/CorpusID:17636888
- King et al. (2009) Ross D. King, Jem Rowland, Stephen G. Oliver, Michael Young, Wayne Aubrey, Emma Byrne, Maria Liakata, Magdalena Markham, Pinar Pir, Larisa N. Soldatova, Andrew Sparkes, Kenneth E. Whelan, and Amanda Clare. 2009. The Automation of Science. Science 324, 5923 (2009), 85–89. https://doi.org/10.1126/science.1165620 arXiv:https://www.science.org/doi/pdf/10.1126/science.1165620
- Kingma and Welling (2013) Diederik P. Kingma and Max Welling. 2013. Auto-Encoding Variational Bayes. CoRR abs/1312.6114 (2013). https://api.semanticscholar.org/CorpusID:216078090
- Kocoń et al. (2023) Jan Kocoń, Igor Cichecki, Oliwier Kaszyca, Mateusz Kochanek, Dominika Szydło, Joanna Baran, Julita Bielaniewicz, Marcin Gruza, Arkadiusz Janz, Kamil Kanclerz, et al. 2023. ChatGPT: Jack of all trades, master of none. Information Fusion 99 (2023), 101861.
- Koivisto and Grassini (2023) Mika Koivisto and Simone Grassini. 2023. Best humans still outperform artificial intelligence in a creative divergent thinking task. Scientific Reports 13 (2023). https://api.semanticscholar.org/CorpusID:261883326
- Kondratyuk et al. (2023) D. Kondratyuk, Lijun Yu, Xiuye Gu, José Lezama, Jonathan Huang, Rachel Hornung, Hartwig Adam, Hassan Akbari, Yair Alon, Vighnesh Birodkar, Yong Cheng, Ming-Chang Chiu, Josh Dillon, Irfan Essa, Agrim Gupta, Meera Hahn, Anja Hauth, David Hendon, Alonso Martinez, David C. Minnen, David A. Ross, Grant Schindler, Mikhail Sirotenko, Kihyuk Sohn, Krishna Somandepalli, Huisheng Wang, Jimmy Yan, Ming Yang, Xuan Yang, Bryan Seybold, and Lu Jiang. 2023. VideoPoet: A Large Language Model for Zero-Shot Video Generation. ArXiv abs/2312.14125 (2023). https://api.semanticscholar.org/CorpusID:266435847
- Koza (1994) John R. Koza. 1994. Genetic programming as a means for programming computers by natural selection. Statistics and Computing 4 (1994), 87–112. https://api.semanticscholar.org/CorpusID:8750149
- Kramer et al. (2023) Stefan Kramer, Mattia Cerrato, Saso Dzeroski, and Ross D. King. 2023. Automated Scientific Discovery: From Equation Discovery to Autonomous Discovery Systems. ArXiv abs/2305.02251 (2023). https://api.semanticscholar.org/CorpusID:258461620
- Kronfeldner (2009) Maria E. Kronfeldner. 2009. Creativity Naturalized. The Philosophical Quarterly 59, 237 (07 2009), 577–592. https://doi.org/10.1111/j.1467-9213.2009.637.x arXiv:https://academic.oup.com/pq/article-pdf/59/237/577/4299738/pq59-0577.pdf
- Kumar et al. (2024) Sreejan Kumar, Raja Marjieh, Byron Zhang, Declan Campbell, Michael Y. Hu, Umang Bhatt, Brenden Lake, and Thomas L. Griffiths. 2024. Comparing Abstraction in Humans and Large Language Models Using Multimodal Serial Reproduction. ArXiv abs/2402.03618 (2024). https://api.semanticscholar.org/CorpusID:267499989
- Kuznetsova et al. (2013) Polina Kuznetsova, Jianfu Chen, and Yejin Choi. 2013. Understanding and Quantifying Creativity in Lexical Composition. In Proceedings of the 2013 Conference on Empirical Methods in Natural Language Processing, David Yarowsky, Timothy Baldwin, Anna Korhonen, Karen Livescu, and Steven Bethard (Eds.). Association for Computational Linguistics, Seattle, Washington, USA, 1246–1258. https://aclanthology.org/D13-1124
- Laban et al. (2021) Philippe Laban, Luke Dai, and Lucas Bandarkar. 2021. Can Transformer Models Measure Coherence In Text: Re-Thinking the Shuffle Test. ArXiv abs/2107.03448 (2021). https://api.semanticscholar.org/CorpusID:235656966
- Lai and Nissim (2024) Huiyuan Lai and Malvina Nissim. 2024. A Survey on Automatic Generation of Figurative Language: From Rule-based Systems to Large Language Models. Comput. Surveys 56 (2024), 1 – 34. https://api.semanticscholar.org/CorpusID:268804894
- Lamb et al. (2015) Carolyn Lamb, Daniel G. Brown, and Charles L. A. Clarke. 2015. Human Competence in Creativity Evaluation. In International Conference on Innovative Computing and Cloud Computing. https://api.semanticscholar.org/CorpusID:14806090
- Lamb et al. (2018) Carolyn Lamb, Daniel G. Brown, and Charles L. A. Clarke. 2018. Evaluating Computational Creativity: An Interdisciplinary Tutorial. ACM Comput. Surv. 51, 2, Article 28 (Feb. 2018), 34 pages. https://doi.org/10.1145/3167476
- Langley (1977) Pat Langley. 1977. BACON: A Production System That Discovers Empirical Laws. In International Joint Conference on Artificial Intelligence. https://api.semanticscholar.org/CorpusID:2320342
- Langlois et al. (2021) Thomas A. Langlois, Nori Jacoby, Jordan W. Suchow, and Thomas L. Griffiths. 2021. Serial reproduction reveals the geometry of visuospatial representations. Proceedings of the National Academy of Sciences of the United States of America 118 (2021). https://api.semanticscholar.org/CorpusID:232376763
- Lau et al. (2018) Jey Han Lau, Trevor Cohn, Timothy Baldwin, Julian Brooke, and Adam Hammond. 2018. Deep-speare: A joint neural model of poetic language, meter and rhyme. In Annual Meeting of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:49671404
- Lavrenko and Pickens (2003) Victor Lavrenko and Jeremy Pickens. 2003. Polyphonic music modeling with random fields. In Proceedings of the eleventh ACM international conference on Multimedia. 120–129.
- Lebowitz (1983) Michael Lebowitz. 1983. Creating a Story-Telling Universe. In International Joint Conference on Artificial Intelligence. https://api.semanticscholar.org/CorpusID:37655080
- Lebowitz (1984) Michael Lebowitz. 1984. Creating characters in a story-telling universe. Poetics 13, 3 (1984), 171–194. https://doi.org/10.1016/0304-422X(84)90001-9
- LeCun et al. (1989) Y. LeCun, B. Boser, J. S. Denker, D. Henderson, R. E. Howard, W. Hubbard, and L. D. Jackel. 1989. Backpropagation Applied to Handwritten Zip Code Recognition. Neural Computation 1, 4 (1989), 541–551. https://doi.org/10.1162/neco.1989.1.4.541
- Ledig et al. (2016) Christian Ledig, Lucas Theis, Ferenc Huszár, Jose Caballero, Andrew P. Aitken, Alykhan Tejani, Johannes Totz, Zehan Wang, and Wenzhe Shi. 2016. Photo-Realistic Single Image Super-Resolution Using a Generative Adversarial Network. 2017 IEEE Conference on Computer Vision and Pattern Recognition (CVPR) (2016), 105–114. https://api.semanticscholar.org/CorpusID:211227
- Lei et al. (2024) Wen-Ling Lei, Jinting Wang, Fengji Ma, Guanjie Huang, and Li Liu. 2024. A Comprehensive Survey on Human Video Generation: Challenges, Methods, and Insights. ArXiv abs/2407.08428 (2024). https://api.semanticscholar.org/CorpusID:271097523
- Leivada et al. (2022) Evelina Leivada, Elliot Murphy, and Gary Marcus. 2022. DALL-E 2 Fails to Reliably Capture Common Syntactic Processes. ArXiv abs/2210.12889 (2022). https://api.semanticscholar.org/CorpusID:253098841
- Lenat and Brown (1984) Douglas B Lenat and John Seely Brown. 1984. Why AM and EURISKO appear to work. Artificial intelligence 23, 3 (1984), 269–294.
- Lencione et al. (2022) Gabriel R Lencione, Rodrigo F Nogueira, and Paula Y Pasqualini. 2022. Nameling: Creative Neologism Generation with Transfer Learning. International Conference on Computational Creativity (06 2022).
- Lewis and Mitchell (2024) Martha Lewis and Melanie Mitchell. 2024. Using Counterfactual Tasks to Evaluate the Generality of Analogical Reasoning in Large Language Models. ArXiv abs/2402.08955 (2024). https://api.semanticscholar.org/CorpusID:267657861
- Li and Wand (2016) Chuan Li and Michael Wand. 2016. Combining Markov Random Fields and Convolutional Neural Networks for Image Synthesis. 2016 IEEE Conference on Computer Vision and Pattern Recognition (CVPR) (2016), 2479–2486. https://api.semanticscholar.org/CorpusID:6635779
- Li et al. (2024a) Haodong Li, Gelei Deng, Yi Liu, Kailong Wang, Yuekang Li, Tianwei Zhang, Yang Liu, Guoai Xu, Guosheng Xu, and Haoyu Wang. 2024a. Digger: Detecting Copyright Content Mis-usage in Large Language Model Training. arXiv:2401.00676 [cs.CR] https://arxiv.org/abs/2401.00676
- Li et al. (2024b) Ruochen Li, Teerth Patel, Qingyun Wang, Qingyun Wang, and Xinya Du. 2024b. MLR-Copilot: Autonomous Machine Learning Research based on Large Language Models Agents. https://api.semanticscholar.org/CorpusID:271957477
- Li et al. (2023) Xiang Lisa Li, Ari Holtzman, Daniel Fried, Percy Liang, Jason Eisner, Tatsunori Hashimoto, Luke Zettlemoyer, and Mike Lewis. 2023. Contrastive Decoding: Open-ended Text Generation as Optimization. arXiv:2210.15097 [cs.CL]
- Li et al. (2021) Zelong Li, Jianchao Ji, and Yongfeng Zhang. 2021. From Kepler to Newton: Explainable AI for Science Discovery. ArXiv abs/2111.12210 (2021). https://api.semanticscholar.org/CorpusID:244728403
- Lin et al. (2021) Bill Yuchen Lin, Ziyi Wu, Yichi Yang, Dong-Ho Lee, and Xiang Ren. 2021. RiddleSense: Reasoning about Riddle Questions Featuring Linguistic Creativity and Commonsense Knowledge. In Findings. https://api.semanticscholar.org/CorpusID:235731975
- Lindsay (1980) Robert K Lindsay. 1980. Applications of artificial intelligence for organic chemistry: the DENDRAL project. (No Title) (1980).
- Liu et al. (2018a) Bei Liu, Jianlong Fu, Makoto P. Kato, and Masatoshi Yoshikawa. 2018a. Beyond Narrative Description: Generating Poetry from Images by Multi-Adversarial Training. In Proceedings of the 26th ACM international conference on Multimedia (MM ’18). ACM. https://doi.org/10.1145/3240508.3240587
- Liu et al. (2018b) Lizhen Liu, Donghai Zhang, and Wei Song. 2018b. Exploiting Syntactic Structures for Humor Recognition. In Proceedings of the 27th International Conference on Computational Linguistics, Emily M. Bender, Leon Derczynski, and Pierre Isabelle (Eds.). Association for Computational Linguistics, Santa Fe, New Mexico, USA, 1875–1883. https://aclanthology.org/C18-1159
- Liu and Xu (2020) Rongzhi Liu and Fei Xu. 2020. Feeling of Competence Affects Children’s Curiosity and Creativity. Cognitive Science (2020). https://api.semanticscholar.org/CorpusID:220039894
- Liu et al. (2024) Yiren Liu, Si Chen, Haocong Cheng, Mengxia Yu, Xiao Ran, Andrew Mo, Yiliu Tang, and Yun Huang. 2024. How AI Processing Delays Foster Creativity: Exploring Research Question Co-Creation with an LLM-based Agent. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 17, 25 pages. https://doi.org/10.1145/3613904.3642698
- Livingston (2018) Paisley Livingston. 2018. ’Explicating ”Creativity”. In Routledge Handbook on Creativity and Philosophy, Berys Gaut and Matthew Kieran (Eds.). Routledge, 108–123.
- Lorello et al. (2024) Luca Salvatore Lorello, Marco Lippi, and Stefano Melacci. 2024. The KANDY Benchmark: Incremental Neuro-Symbolic Learning and Reasoning with Kandinsky Patterns. ArXiv abs/2402.17431 (2024). https://api.semanticscholar.org/CorpusID:268033655
- Louis and Sutton (2018) Annie Louis and Charles Sutton. 2018. Deep Dungeons and Dragons: Learning Character-Action Interactions from Role-Playing Game Transcripts. In Proceedings of the 2018 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 2 (Short Papers), Marilyn Walker, Heng Ji, and Amanda Stent (Eds.). Association for Computational Linguistics, New Orleans, Louisiana, 708–713. https://doi.org/10.18653/v1/N18-2111
- Lovelace (1843) Lady Lovelace. 1843. Sketch of the Analytical Engine invented by Charles Babbage, Esq. Ada’s Legacy: Cultures of Computing from the Victorian to the Digital Age (1843). https://api.semanticscholar.org/CorpusID:257837181
- Lu et al. (2024a) Chris Lu, Cong Lu, Robert Tjarko Lange, Jakob Foerster, Jeff Clune, and David Ha. 2024a. The AI Scientist: Towards Fully Automated Open-Ended Scientific Discovery. arXiv:2408.06292 [cs.AI] https://arxiv.org/abs/2408.06292
- Lu et al. (2024b) Ximing Lu, Melanie Sclar, Skyler Hallinan, Niloofar Mireshghallah, Jiacheng Liu, Seungju Han, Allyson Ettinger, Liwei Jiang, Khyathi Chandu, Nouha Dziri, and Yejin Choi. 2024b. AI as Humanity’s Salieri: Quantifying Linguistic Creativity of Language Models via Systematic Attribution of Machine Text against Web Text. arXiv:2410.04265 [cs.CL] https://arxiv.org/abs/2410.04265
- Lu et al. (2024c) Yining Lu, Dixuan Wang, Tianjian Li, Dongwei Jiang, and Daniel Khashabi. 2024c. Benchmarking Language Model Creativity: A Case Study on Code Generation. ArXiv abs/2407.09007 (2024). https://api.semanticscholar.org/CorpusID:271162279
- Luo et al. (2023) Zhengxiong Luo, Dayou Chen, Yingya Zhang, Yan Huang, Liangsheng Wang, Yujun Shen, Deli Zhao, Jinren Zhou, and Tien-Ping Tan. 2023. VideoFusion: Decomposed Diffusion Models for High-Quality Video Generation. 2023 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2023), 10209–10218. https://api.semanticscholar.org/CorpusID:257532642
- Maher (2010) Mary Lou Maher. 2010. Evaluating creativity in humans, computers, and collectively intelligent systems. In Network Conference on Creativity and Innovation in Design. https://api.semanticscholar.org/CorpusID:4840039
- Maher and Fisher (2012) Mary Lou Maher and Douglas H Fisher. 2012. Using AI to evaluate creative designs. In DS 73-1 Proceedings of the 2nd International Conference on Design Creativity Volume 1.
- Majumder et al. (2024) Bodhisattwa Prasad Majumder, Harshit Surana, Dhruv Agarwal, Bhavana Dalvi Mishra, Abhijeetsingh Meena, Aryan Prakhar, Tirth Vora, Tushar Khot, Ashish Sabharwal, and Peter Clark. 2024. DiscoveryBench: Towards Data-Driven Discovery with Large Language Models. arXiv:2407.01725 [cs.CL] https://arxiv.org/abs/2407.01725
- Mansimov et al. (2015) Elman Mansimov, Emilio Parisotto, Jimmy Ba, and Ruslan Salakhutdinov. 2015. Generating Images from Captions with Attention. CoRR abs/1511.02793 (2015). https://api.semanticscholar.org/CorpusID:9996719
- Manurung (2004) Hisar Maruli Manurung. 2004. An evolutionary algorithm approach to poetry generation. https://api.semanticscholar.org/CorpusID:170162691
- Manurung et al. (2000) Ruli Manurung, Graeme Ritchie, and Henry Thompson. 2000. A Flexible Integrated Architecture For Generating Poetic Texts. (07 2000).
- Manurung et al. (2012) Ruli Manurung, Graeme D. Ritchie, and Henry S. Thompson. 2012. Using genetic algorithms to create meaningful poetic text. Journal of Experimental & Theoretical Artificial Intelligence 24 (2012), 43 – 64. https://api.semanticscholar.org/CorpusID:4528781
- Marco et al. (2024) Guillermo Marco, Julio Gonzalo, Ram’on del Castillo, and Mar’ia Teresa Mateo Girona. 2024. Pron vs Prompt: Can Large Language Models already Challenge a World-Class Fiction Author at Creative Text Writing? ArXiv abs/2407.01119 (2024). https://api.semanticscholar.org/CorpusID:270870769
- Marcus et al. (2022) Gary Marcus, Ernest Davis, and Scott Aaronson. 2022. A very preliminary analysis of DALL-E 2. ArXiv abs/2204.13807 (2022). https://api.semanticscholar.org/CorpusID:248476147
- Marcus (2020) Gary F. Marcus. 2020. The Next Decade in AI: Four Steps Towards Robust Artificial Intelligence. ArXiv abs/2002.06177 (2020). https://api.semanticscholar.org/CorpusID:211126492
- Marquer et al. (2022) Esteban Marquer, Pierre-Alexandre Murena, and Miguel Couceiro. 2022. Transferring Learned Models of Morphological Analogy. https://api.semanticscholar.org/CorpusID:258688324
- Masterman (1971) Margaret Masterman. 1971. ’Computerized haiku’, in Cybernetics, art and ideas. London Studio Vista.
- McCoy et al. (2021) R. Thomas McCoy, Paul Smolensky, Tal Linzen, Jianfeng Gao, and Asli Celikyilmaz. 2021. How Much Do Language Models Copy From Their Training Data? Evaluating Linguistic Novelty in Text Generation Using RAVEN. Transactions of the Association for Computational Linguistics 11 (2021), 652–670. https://api.semanticscholar.org/CorpusID:244345615
- McCoy et al. (2023) R. Thomas McCoy, Shunyu Yao, Dan Friedman, Matthew Hardy, and Thomas L. Griffiths. 2023. Embers of Autoregression: Understanding Large Language Models Through the Problem They are Trained to Solve. ArXiv abs/2309.13638 (2023). https://api.semanticscholar.org/CorpusID:262464572
- Meaney et al. (2021) J. A. Meaney, Steven Wilson, Luis Chiruzzo, Adam Lopez, and Walid Magdy. 2021. SemEval 2021 Task 7: HaHackathon, Detecting and Rating Humor and Offense. In Proceedings of the 15th International Workshop on Semantic Evaluation (SemEval-2021), Alexis Palmer, Nathan Schneider, Natalie Schluter, Guy Emerson, Aurelie Herbelot, and Xiaodan Zhu (Eds.). Association for Computational Linguistics, Online, 105–119. https://doi.org/10.18653/v1/2021.semeval-1.9
- Mednick (1962) Sarnoff A. Mednick. 1962. The associative basis of the creative process. Psychological review 69 (1962), 220–32. https://api.semanticscholar.org/CorpusID:6702759
- Meehan (1977) James R. Meehan. 1977. TALE-SPIN, An Interactive Program that Writes Stories. In International Joint Conference on Artificial Intelligence. https://api.semanticscholar.org/CorpusID:2372981
- Mehrotra et al. (2024) Pronita Mehrotra, Aishni Parab, and Sumit Gulwani. 2024. Enhancing Creativity in Large Language Models through Associative Thinking Strategies. ArXiv abs/2405.06715 (2024). https://api.semanticscholar.org/CorpusID:269756853
- Meister et al. (2023) Clara Meister, Tiago Pimentel, Gian Wiher, and Ryan Cotterell. 2023. Locally Typical Sampling. Transactions of the Association for Computational Linguistics 11 (2023), 102–121. https://doi.org/10.1162/tacl_a_00536
- Meister et al. (2022) Clara Meister, Gian Wiher, Tiago Pimentel, and Ryan Cotterell. 2022. On the probability–quality paradox in language generation. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), Smaranda Muresan, Preslav Nakov, and Aline Villavicencio (Eds.). Association for Computational Linguistics, Dublin, Ireland, 36–45. https://doi.org/10.18653/v1/2022.acl-short.5
- Mihalcea and Strapparava (2005) Rada Mihalcea and Carlo Strapparava. 2005. Making Computers Laugh: Investigations in Automatic Humor Recognition. In Proceedings of Human Language Technology Conference and Conference on Empirical Methods in Natural Language Processing, Raymond Mooney, Chris Brew, Lee-Feng Chien, and Katrin Kirchhoff (Eds.). Association for Computational Linguistics, Vancouver, British Columbia, Canada, 531–538. https://aclanthology.org/H05-1067
- Mikolov et al. (2013) Tomas Mikolov, Ilya Sutskever, Kai Chen, Greg S Corrado, and Jeff Dean. 2013. Distributed Representations of Words and Phrases and their Compositionality. In Advances in Neural Information Processing Systems, C.J. Burges, L. Bottou, M. Welling, Z. Ghahramani, and K.Q. Weinberger (Eds.), Vol. 26. Curran Associates, Inc. https://proceedings.neurips.cc/paper_files/paper/2013/file/9aa42b31882ec039965f3c4923ce901b-Paper.pdf
- Milic (1970) Louis T. Milic. 1970. The Possible Usefulness of Poetry Generation. https://api.semanticscholar.org/CorpusID:59924092
- Miller (1993) Arthur R. Miller. 1993. COPYRIGHT PROTECTION FOR COMPUTER PROGRAMS, DATABASES, AND COMPUTER-GENERATED WORKS: IS ANYTHING NEW SINCE CONTU? Harvard Law Review 106 (1993), 978–1073. https://api.semanticscholar.org/CorpusID:141462591
- Miller et al. (2017) Tristan Miller, Christian Hempelmann, and Iryna Gurevych. 2017. SemEval-2017 Task 7: Detection and Interpretation of English Puns. In Proceedings of the 11th International Workshop on Semantic Evaluation (SemEval-2017), Steven Bethard, Marine Carpuat, Marianna Apidianaki, Saif M. Mohammad, Daniel Cer, and David Jurgens (Eds.). Association for Computational Linguistics, Vancouver, Canada, 58–68. https://doi.org/10.18653/v1/S17-2005
- Minsky (1982) Marvin Minsky. 1982. Why People Think Computers Can’t. AI Mag. 3 (1982), 3–15. https://api.semanticscholar.org/CorpusID:42565554
- Mirchandani et al. (2023) Suvir Mirchandani, F. Xia, Peter R. Florence, Brian Ichter, Danny Driess, Montse Gonzalez Arenas, Kanishka Rao, Dorsa Sadigh, and Andy Zeng. 2023. Large Language Models as General Pattern Machines. ArXiv abs/2307.04721 (2023). https://api.semanticscholar.org/CorpusID:259501163
- Mirowski et al. (2022) Piotr Wojciech Mirowski, Kory Wallace Mathewson, Jaylen Pittman, and Richard Evans. 2022. Co-Writing Screenplays and Theatre Scripts with Language Models: Evaluation by Industry Professionals. Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (2022). https://api.semanticscholar.org/CorpusID:252596159
- Mirza and Osindero (2014) Mehdi Mirza and Simon Osindero. 2014. Conditional Generative Adversarial Nets. ArXiv abs/1411.1784 (2014). https://api.semanticscholar.org/CorpusID:12803511
- Mitchell (2021) Melanie Mitchell. 2021. Abstraction and analogy‐making in artificial intelligence. Annals of the New York Academy of Sciences 1505, 1 (June 2021), 79–101. https://doi.org/10.1111/nyas.14619
- Mitchell et al. (2023) Melanie Mitchell, Alessandro B. Palmarini, and Arseny Moskvichev. 2023. Comparing Humans, GPT-4, and GPT-4V On Abstraction and Reasoning Tasks. ArXiv abs/2311.09247 (2023). https://api.semanticscholar.org/CorpusID:265220802
- Mittal et al. (2022) Anirudh Mittal, Yufei Tian, and Nanyun Peng. 2022. AmbiPun: Generating Humorous Puns with Ambiguous Context. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Marine Carpuat, Marie-Catherine de Marneffe, and Ivan Vladimir Meza Ruiz (Eds.). Association for Computational Linguistics, Seattle, United States, 1053–1062. https://doi.org/10.18653/v1/2022.naacl-main.77
- Mohammad et al. (2016) Saif M. Mohammad, Ekaterina Shutova, and Peter D. Turney. 2016. Metaphor as a Medium for Emotion: An Empirical Study. In International Workshop on Semantic Evaluation. https://api.semanticscholar.org/CorpusID:989439
- Mohammadi (2024) Behnam Mohammadi. 2024. Creativity Has Left the Chat: The Price of Debiasing Language Models. arXiv:2406.05587 [cs.CL] https://arxiv.org/abs/2406.05587
- Mohler et al. (2016) Michael Mohler, Mary Brunson, Bryan Rink, and Marc Tomlinson. 2016. Introducing the LCC Metaphor Datasets. In Proceedings of the Tenth International Conference on Language Resources and Evaluation (LREC’16), Nicoletta Calzolari, Khalid Choukri, Thierry Declerck, Sara Goggi, Marko Grobelnik, Bente Maegaard, Joseph Mariani, Helene Mazo, Asuncion Moreno, Jan Odijk, and Stelios Piperidis (Eds.). European Language Resources Association (ELRA), Portorož, Slovenia, 4221–4227. https://aclanthology.org/L16-1668
- Moskvichev et al. (2023) Arseny Moskvichev, Victor Vikram Odouard, and Melanie Mitchell. 2023. The ConceptARC Benchmark: Evaluating Understanding and Generalization in the ARC Domain. ArXiv abs/2305.07141 (2023). https://api.semanticscholar.org/CorpusID:258676355
- Murphy et al. (2024) Elliot Murphy, Jill de Villiers, and Sofia Morales. 2024. A Comparative Investigation of Compositional Syntax and Semantics in DALL-E 2. ArXiv abs/2403.12294 (2024). https://api.semanticscholar.org/CorpusID:268531835
- Musker et al. (2024) Sam Musker, Alex Duchnowski, Raphael Milliere, and Ellie Pavlick. 2024. Semantic Structure-Mapping in LLM and Human Analogical Reasoning. https://api.semanticscholar.org/CorpusID:270619799
- Naeini et al. (2023) Saeid Naeini, Raeid Saqur, Mozhgan Saeidi, John Giorgi, and Babak Taati. 2023. Large Language Models are Fixated by Red Herrings: Exploring Creative Problem Solving and Einstellung Effect using the Only Connect Wall Dataset. arXiv:2306.11167
- Nagarajah et al. (2022) Thiloshon Nagarajah, Filip Ilievski, and Jay Pujara. 2022. Understanding Narratives through Dimensions of Analogy. arXiv:2206.07167 [cs.AI] https://arxiv.org/abs/2206.07167
- Nair et al. (2024) Lakshmi Nair, Evana Gizzi, and Jivko Sinapov. 2024. Creative Problem Solving in Large Language and Vision Models - What Would it Take? ArXiv abs/2405.01453 (2024). https://api.semanticscholar.org/CorpusID:269502505
- Nath et al. (2024) Surabhi S. Nath, Peter Dayan, and Claire Stevenson. 2024. Characterising the Creative Process in Humans and Large Language Models. arXiv:2405.00899 [cs.HC] https://arxiv.org/abs/2405.00899
- Netanel (2011) Neil Weinstock Netanel. 2011. Making Sense of Fair Use. https://api.semanticscholar.org/CorpusID:152743703
- Newell et al. (1959) Allen Newell, J. C. Shaw, and Herbert A. Simon. 1959. The Processes of Creative Thinking. https://api.semanticscholar.org/CorpusID:121728459
- Nichol et al. (2021) Alex Nichol, Prafulla Dhariwal, Aditya Ramesh, Pranav Shyam, Pamela Mishkin, Bob McGrew, Ilya Sutskever, and Mark Chen. 2021. GLIDE: Towards Photorealistic Image Generation and Editing with Text-Guided Diffusion Models. In International Conference on Machine Learning. https://api.semanticscholar.org/CorpusID:245335086
- Nijstad et al. (2010) Bernard A. Nijstad, C. D. De Dreu, Eric F. Rietzschel, and Matthijs Baas. 2010. The dual pathway to creativity model: Creative ideation as a function of flexibility and persistence. European Review of Social Psychology 21 (2010), 34 – 77. https://api.semanticscholar.org/CorpusID:145290967
- Odouard and Mitchell (2022) Victor Vikram Odouard and Melanie Mitchell. 2022. Evaluating Understanding on Conceptual Abstraction Benchmarks. ArXiv abs/2206.14187 (2022). https://api.semanticscholar.org/CorpusID:250089291
- Oliveira (2012) Hugo Gonçalo Oliveira. 2012. PoeTryMe : a versatile platform for poetry generation. https://api.semanticscholar.org/CorpusID:51964516
- Oliveira (2017) Hugo Gonçalo Oliveira. 2017. A Survey on Intelligent Poetry Generation: Languages, Features, Techniques, Reutilisation and Evaluation. In International Conference on Natural Language Generation. https://api.semanticscholar.org/CorpusID:32461868
- Olson et al. (2021) Jay A. Olson, Johnny Nahas, Denis Chmoulevitch, Simon J. Cropper, and Margaret E. Webb. 2021. Naming unrelated words predicts creativity. Proceedings of the National Academy of Sciences of the United States of America 118 (2021). https://api.semanticscholar.org/CorpusID:235472442
- OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. arXiv:2303.08774 [cs.CL]
- Opielka et al. (2024) Gustaw Opielka, Hannes Rosenbusch, Veerle Vijverberg, and Claire E. Stevenson. 2024. Do Large Language Models Solve ARC Visual Analogies Like People Do? ArXiv abs/2403.09734 (2024). https://api.semanticscholar.org/CorpusID:268510601
- Oren et al. (2023) Yonatan Oren, Nicole Meister, Niladri Chatterji, Faisal Ladhak, and Tatsunori B. Hashimoto. 2023. Proving Test Set Contamination in Black Box Language Models. arXiv:2310.17623 [cs.CL] https://arxiv.org/abs/2310.17623
- Ormazabal et al. (2022a) Aitor Ormazabal, Mikel Artetxe, Manex Agirrezabal, Aitor Soroa Etxabe, and Eneko Agirre. 2022a. PoeLM: A Meter- and Rhyme-Controllable Language Model for Unsupervised Poetry Generation. In Conference on Empirical Methods in Natural Language Processing. https://api.semanticscholar.org/CorpusID:249018166
- Ormazabal et al. (2022b) Aitor Ormazabal, Mikel Artetxe, Manex Agirrezabal, Aitor Soroa, and Eneko Agirre. 2022b. PoeLM: A Meter- and Rhyme-Controllable Language Model for Unsupervised Poetry Generation. In Findings of the Association for Computational Linguistics: EMNLP 2022, Yoav Goldberg, Zornitsa Kozareva, and Yue Zhang (Eds.). Association for Computational Linguistics, Abu Dhabi, United Arab Emirates, 3655–3670. https://doi.org/10.18653/v1/2022.findings-emnlp.268
- Osborn (1957) Alexander Faickney Osborn. 1957. Applied imagination : principles and procedures of creative problem-solving. https://api.semanticscholar.org/CorpusID:142858886
- Ouyang et al. (2022) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L. Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke E. Miller, Maddie Simens, Amanda Askell, Peter Welinder, Paul Francis Christiano, Jan Leike, and Ryan J. Lowe. 2022. Training language models to follow instructions with human feedback. ArXiv abs/2203.02155 (2022). https://api.semanticscholar.org/CorpusID:246426909
- Padmakumar and He (2023) Vishakh Padmakumar and He He. 2023. Does Writing with Language Models Reduce Content Diversity? ArXiv abs/2309.05196 (2023). https://api.semanticscholar.org/CorpusID:261682154
- Pascual et al. (2021) Damian Pascual, Béni Egressy, Clara Meister, Ryan Cotterell, and Roger Wattenhofer. 2021. A Plug-and-Play Method for Controlled Text Generation. ArXiv abs/2109.09707 (2021). https://api.semanticscholar.org/CorpusID:237571784
- Patashnik et al. (2021) Or Patashnik, Zongze Wu, Eli Shechtman, Daniel Cohen-Or, and Dani Lischinski. 2021. StyleCLIP: Text-Driven Manipulation of StyleGAN Imagery. 2021 IEEE/CVF International Conference on Computer Vision (ICCV) (2021), 2065–2074. https://api.semanticscholar.org/CorpusID:232428282
- Pathak et al. (2016) Deepak Pathak, Philipp Krähenbühl, Jeff Donahue, Trevor Darrell, and Alexei A. Efros. 2016. Context Encoders: Feature Learning by Inpainting. 2016 IEEE Conference on Computer Vision and Pattern Recognition (CVPR) (2016), 2536–2544. https://api.semanticscholar.org/CorpusID:2202933
- Paul (1970) Anthony M. Paul. 1970. Figurative Language. Philosophy and Rhetoric 3, 4 (1970), 225–248.
- Paul and Frank (2021) Debjit Paul and Anette Frank. 2021. COINS: Dynamically Generating COntextualized Inference Rules for Narrative Story Completion. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), Chengqing Zong, Fei Xia, Wenjie Li, and Roberto Navigli (Eds.). Association for Computational Linguistics, Online, 5086–5099. https://doi.org/10.18653/v1/2021.acl-long.395
- Payne (2019) Christine Payne. 2019. MuseNet, 2019. URL https://openai. com/blog/musenet (2019).
- Pease et al. (2002) Alison Pease, Daniel Winterstein, and Simon Colton. 2002. Evaluating Machine Creativity. (10 2002).
- Peeperkorn et al. (2024) Max Peeperkorn, Tom Kouwenhoven, Dan Brown, and Anna Jordanous. 2024. Is Temperature the Creativity Parameter of Large Language Models? arXiv:2405.00492 [cs.CL] https://arxiv.org/abs/2405.00492
- Perkins (2001) D.N. Perkins. 2001. The Eureka Effect: The Art and Logic of Breakthrough Thinking. W. W. Norton, Incorporated. https://books.google.ch/books?id=84HVnfUywZoC
- Petersen and Landajuela (2019) Brenden K. Petersen and Mikel Landajuela. 2019. Deep symbolic regression: Recovering mathematical expressions from data via risk-seeking policy gradients. arXiv: Learning (2019). https://api.semanticscholar.org/CorpusID:212956516
- Petersen and van der Plas (2023) Molly R. Petersen and Lonneke van der Plas. 2023. Can language models learn analogical reasoning? Investigating training objectives and comparisons to human performance. In Conference on Empirical Methods in Natural Language Processing. https://api.semanticscholar.org/CorpusID:263828777
- Peyrard et al. (2021) Maxime Peyrard, Beatriz Borges, Kristina Gligoric, and Robert West. 2021. Laughing Heads: Can Transformers Detect What Makes a Sentence Funny?. In International Joint Conference on Artificial Intelligence. https://api.semanticscholar.org/CorpusID:234778118
- Popescu-Belis et al. (2023) Andrei Popescu-Belis, Àlex R. Atrio, Bastien Bernath, Etienne Boisson, Teo Ferrari, Xavier Theimer-lienhard, and Giorgos Vernikos. 2023. GPoeT: a Language Model Trained for Rhyme Generation on Synthetic Data. Proceedings of the 7th Joint SIGHUM Workshop on Computational Linguistics for Cultural Heritage, Social Sciences, Humanities and Literature (2023). https://api.semanticscholar.org/CorpusID:258486928
- Potash et al. (2017) Peter Potash, Alexey Romanov, and Anna Rumshisky. 2017. SemEval-2017 Task 6: #HashtagWars: Learning a Sense of Humor. In Proceedings of the 11th International Workshop on Semantic Evaluation (SemEval-2017), Steven Bethard, Marine Carpuat, Marianna Apidianaki, Saif M. Mohammad, Daniel Cer, and David Jurgens (Eds.). Association for Computational Linguistics, Vancouver, Canada, 49–57. https://doi.org/10.18653/v1/S17-2004
- Prabhakaran et al. (2013) Ranjani Prabhakaran, Adam E. Green, and Jeremy R. Gray. 2013. Thin slices of creativity: Using single-word utterances to assess creative cognition. Behavior Research Methods 46 (2013), 641 – 659. https://api.semanticscholar.org/CorpusID:2159817
- Qi et al. (2023) Biqing Qi, Kaiyan Zhang, Haoxiang Li, Kai Tian, Sihang Zeng, Zhang-Ren Chen, and Bowen Zhou. 2023. Large Language Models are Zero Shot Hypothesis Proposers. arXiv:2311.05965 [cs.CL] https://arxiv.org/abs/2311.05965
- Qiao et al. (2022) Shuofei Qiao, Yixin Ou, Ningyu Zhang, Xiang Chen, Yunzhi Yao, Shumin Deng, Chuanqi Tan, Fei Huang, and Huajun Chen. 2022. Reasoning with Language Model Prompting: A Survey. ArXiv abs/2212.09597 (2022). https://api.semanticscholar.org/CorpusID:254854219
- Qu et al. (2024) Xingwei Qu, Yuelin Bai, Yinghao Ma, Ziya Zhou, Ka Man Lo, Jiaheng Liu, Ruibin Yuan, Lejun Min, Xueling Liu, Tianyu Zhang, et al. 2024. Mupt: A generative symbolic music pretrained transformer. arXiv preprint arXiv:2404.06393 (2024).
- Raayoni et al. (2021) Gal Raayoni, Shahar Gottlieb, Yahel Manor, George Pisha, Yoav Harris, Uri Mendlovic, Doron Haviv, Yaron Hadad, and Ido Kaminer. 2021. Generating conjectures on fundamental constants with the Ramanujan Machine. Nature 590, 7844 (Feb. 2021), 67–73. https://doi.org/10.1038/s41586-021-03229-4
- Racter (1984) Racter. 1984. The Policeman’s Beard Is Half Constructed. https://api.semanticscholar.org/CorpusID:146585648
- Radev et al. (2016) Dragomir Radev, Amanda Stent, Joel Tetreault, Aasish Pappu, Aikaterini Iliakopoulou, Agustin Chanfreau, Paloma de Juan, Jordi Vallmitjana, Alejandro Jaimes, Rahul Jha, and Robert Mankoff. 2016. Humor in Collective Discourse: Unsupervised Funniness Detection in the New Yorker Cartoon Caption Contest. In Proceedings of the Tenth International Conference on Language Resources and Evaluation (LREC’16), Nicoletta Calzolari, Khalid Choukri, Thierry Declerck, Sara Goggi, Marko Grobelnik, Bente Maegaard, Joseph Mariani, Helene Mazo, Asuncion Moreno, Jan Odijk, and Stelios Piperidis (Eds.). European Language Resources Association (ELRA), Portorož, Slovenia, 475–479. https://aclanthology.org/L16-1076
- Radford et al. (2015) Alec Radford, Luke Metz, and Soumith Chintala. 2015. Unsupervised Representation Learning with Deep Convolutional Generative Adversarial Networks. CoRR abs/1511.06434 (2015). https://api.semanticscholar.org/CorpusID:11758569
- Raffel et al. (2023) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J. Liu. 2023. Exploring the Limits of Transfer Learning with a Unified Text-to-Text Transformer. arXiv:1910.10683 [cs.LG] https://arxiv.org/abs/1910.10683
- Ralston (2005) W.T. Ralston. 2005. Copyright in computer-composed music: Hal meets Handel. Journal of the Copyright Society of the U.S.A 52 (03 2005), 281–284.
- Ramesh et al. (2022) Aditya Ramesh, Prafulla Dhariwal, Alex Nichol, Casey Chu, and Mark Chen. 2022. Hierarchical Text-Conditional Image Generation with CLIP Latents. ArXiv abs/2204.06125 (2022). https://api.semanticscholar.org/CorpusID:248097655
- Ramesh et al. (2021) Aditya Ramesh, Mikhail Pavlov, Gabriel Goh, Scott Gray, Chelsea Voss, Alec Radford, Mark Chen, and Ilya Sutskever. 2021. Zero-Shot Text-to-Image Generation. ArXiv abs/2102.12092 (2021). https://api.semanticscholar.org/CorpusID:232035663
- Rashkin et al. (2020) Hannah Rashkin, Asli Celikyilmaz, Yejin Choi, and Jianfeng Gao. 2020. PlotMachines: Outline-Conditioned Generation with Dynamic Plot State Tracking. ArXiv abs/2004.14967 (2020). https://api.semanticscholar.org/CorpusID:216868683
- Raskin and Attardo (1994) Jonathan D. Raskin and Salvatore Attardo. 1994. Non-literalness and non-bona-fîde in language: An approach to formal and computational treatments of humor. Pragmatics & Cognition 2 (1994), 31–69. https://api.semanticscholar.org/CorpusID:62211347
- Rassin et al. (2022) Royi Rassin, Shauli Ravfogel, and Yoav Goldberg. 2022. DALLE-2 is Seeing Double: Flaws in Word-to-Concept Mapping in Text2Image Models. ArXiv abs/2210.10606 (2022). https://api.semanticscholar.org/CorpusID:252992545
- Raven (1938) John Carlyle Raven. 1938. Raven Progressive Matrices. https://api.semanticscholar.org/CorpusID:61756432
- Ravi et al. (2024) Sahithya Ravi, Patrick Huber, Akshat Shrivastava, Vered Shwartz, and Arash Einolghozati. 2024. Small But Funny: A Feedback-Driven Approach to Humor Distillation. In Proceedings of the 62nd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.). Association for Computational Linguistics, Bangkok, Thailand, 13078–13090. https://aclanthology.org/2024.acl-long.706
- Reed et al. (2016a) Scott E. Reed, Zeynep Akata, Santosh Mohan, Samuel Tenka, Bernt Schiele, and Honglak Lee. 2016a. Learning What and Where to Draw. ArXiv abs/1610.02454 (2016). https://api.semanticscholar.org/CorpusID:1515901
- Reed et al. (2016b) Scott E. Reed, Zeynep Akata, Xinchen Yan, Lajanugen Logeswaran, Bernt Schiele, and Honglak Lee. 2016b. Generative Adversarial Text to Image Synthesis. In International Conference on Machine Learning. https://api.semanticscholar.org/CorpusID:1563370
- Ren and Yang (2017) He Ren and Quan Yang. 2017. Neural Joke Generation. https://api.semanticscholar.org/CorpusID:19647037
- Rhodes (1961) Mel Rhodes. 1961. An Analysis of Creativity. The Phi Delta Kappan 42, 7 (1961), 305–310. http://www.jstor.org/stable/20342603
- Richardson et al. (2020) Elad Richardson, Yuval Alaluf, Or Patashnik, Yotam Nitzan, Yaniv Azar, Stav Shapiro, and Daniel Cohen-Or. 2020. Encoding in Style: a StyleGAN Encoder for Image-to-Image Translation. 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2020), 2287–2296. https://api.semanticscholar.org/CorpusID:220936362
- Riley and Gabora (2019) Sean N. Riley and Liane Gabora. 2019. Evidence that Threatening Situations Enhance Creativity. arXiv:1308.4245 [q-bio.NC] https://arxiv.org/abs/1308.4245
- Ritchie (2007) Graeme D. Ritchie. 2007. Some Empirical Criteria for Attributing Creativity to a Computer Program. Minds and Machines 17 (2007), 67–99. https://api.semanticscholar.org/CorpusID:7947377
- Roberts et al. (2018) Adam Roberts, Jesse Engel, Colin Raffel, Curtis Hawthorne, and Douglas Eck. 2018. A Hierarchical Latent Vector Model for Learning Long-Term Structure in Music. ArXiv abs/1803.05428 (2018). https://api.semanticscholar.org/CorpusID:3891811
- Roberts and Kreuz (1994) Richard M Roberts and Roger J Kreuz. 1994. Why do people use figurative language? Psychological science 5, 3 (1994), 159–163.
- Rombach et al. (2021) Robin Rombach, A. Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2021. High-Resolution Image Synthesis with Latent Diffusion Models. 2022 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2021), 10674–10685. https://api.semanticscholar.org/CorpusID:245335280
- Rowe and Partridge (1993) Jon Rowe and Derek Partridge. 1993. Creativity: a survey of AI approaches. Artificial Intelligence Review 7 (1993), 43–70. https://api.semanticscholar.org/CorpusID:13433112
- Ruiz et al. (2022) Nataniel Ruiz, Yuanzhen Li, Varun Jampani, Yael Pritch, Michael Rubinstein, and Kfir Aberman. 2022. DreamBooth: Fine Tuning Text-to-Image Diffusion Models for Subject-Driven Generation. 2023 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2022), 22500–22510. https://api.semanticscholar.org/CorpusID:251800180
- Runco and Jaeger (2012) Mark A. Runco and Garrett J. Jaeger. 2012. The Standard Definition of Creativity. Creativity Research Journal 24, 1 (2012), 92–96. https://doi.org/10.1080/10400419.2012.650092 arXiv:https://doi.org/10.1080/10400419.2012.650092
- Rzevski et al. (1987) George Rzevski, Pat Langley, Herbert A. Simon, Gary L. Bradshaw, and Jan M. Zytkow. 1987. Scientific discovery: compulalional explorations of the creative process. https://api.semanticscholar.org/CorpusID:62762320
- Saharia et al. (2022) Chitwan Saharia, William Chan, Saurabh Saxena, Lala Li, Jay Whang, Emily L. Denton, Seyed Kamyar Seyed Ghasemipour, Burcu Karagol Ayan, Seyedeh Sara Mahdavi, Raphael Gontijo Lopes, Tim Salimans, Jonathan Ho, David J. Fleet, and Mohammad Norouzi. 2022. Photorealistic Text-to-Image Diffusion Models with Deep Language Understanding. ArXiv abs/2205.11487 (2022). https://api.semanticscholar.org/CorpusID:248986576
- Samadarshi et al. (2024) Prisha Samadarshi, Mariam Mustafa, Anushka Kulkarni, Raven Rothkopf, Tuhin Chakrabarty, and Smaranda Muresan. 2024. Connecting the Dots: Evaluating Abstract Reasoning Capabilities of LLMs Using the New York Times Connections Word Game. arXiv:2406.11012 [cs.CL] https://arxiv.org/abs/2406.11012
- Sampson (2017) Geoffrey Sampson. 2017. The Linguistics Delusion. Equinox Publishing. https://www.equinoxpub.com/home/linguistics-delusion/
- Samuelson (1986) Pamela Samuelson. 1986. Allocating Ownership Rights in Computer-Generated Works. University of Pittsburgh Law Review 47 (1986), 1185. https://api.semanticscholar.org/CorpusID:142911255
- Santoro et al. (2018) Adam Santoro, Felix Hill, David G. T. Barrett, Ari S. Morcos, and Timothy P. Lillicrap. 2018. Measuring abstract reasoning in neural networks. ArXiv abs/1807.04225 (2018). https://api.semanticscholar.org/CorpusID:49665167
- Saphra and Wiegreffe (2024) Naomi Saphra and Sarah Wiegreffe. 2024. Mechanistic? arXiv:2410.09087 [cs.AI] https://arxiv.org/abs/2410.09087
- Sarmento et al. (2024) Pedro Sarmento, John H. Loth, and Mathieu Barthet. 2024. Between the AI and Me: Analysing Listeners’ Perspectives on AI- and Human-Composed Progressive Metal Music. ArXiv abs/2407.21615 (2024). https://api.semanticscholar.org/CorpusID:271571547
- Savage (2012) Neil Savage. 2012. Automating scientific discovery. Commun. ACM 55, 5 (2012), 9–11.
- Schick et al. (2022) Timo Schick, Jane Dwivedi-Yu, Zhengbao Jiang, Fabio Petroni, Patrick Lewis, Gautier Izacard, Qingfei You, Christoforos Nalmpantis, Edouard Grave, and Sebastian Riedel. 2022. PEER: A Collaborative Language Model. ArXiv abs/2208.11663 (2022). https://api.semanticscholar.org/CorpusID:251765117
- Schmidt and Lipson (2009) Michael Schmidt and Hod Lipson. 2009. Distilling free-form natural laws from experimental data. science 324, 5923 (2009), 81–85.
- Shahaf et al. (2015) Dafna Shahaf, Eric Horvitz, and Robert Mankoff. 2015. Inside Jokes: Identifying Humorous Cartoon Captions. Proceedings of the 21th ACM SIGKDD International Conference on Knowledge Discovery and Data Mining (2015). https://api.semanticscholar.org/CorpusID:14570747
- Shen et al. (2024) Tiancheng Shen, Jun Hao Liew, Long Mai, Lu Qi, Jiashi Feng, and Jiaya Jia. 2024. Empowering Visual Creativity: A Vision-Language Assistant to Image Editing Recommendations. ArXiv abs/2406.00121 (2024). https://api.semanticscholar.org/CorpusID:270218211
- Shi et al. (2024) Weijia Shi, Anirudh Ajith, Mengzhou Xia, Yangsibo Huang, Daogao Liu, Terra Blevins, Danqi Chen, and Luke Zettlemoyer. 2024. Detecting Pretraining Data from Large Language Models. arXiv:2310.16789 [cs.CL] https://arxiv.org/abs/2310.16789
- Si et al. (2024) Chenglei Si, Diyi Yang, and Tatsunori Hashimoto. 2024. Can LLMs Generate Novel Research Ideas? A Large-Scale Human Study with 100+ NLP Researchers. arXiv:2409.04109 [cs.CL] https://arxiv.org/abs/2409.04109
- Simonton (2012) Dean Keith Simonton. 2012. Taking the U.S. Patent Office Criteria Seriously: A Quantitative Three-Criterion Creativity Definition and Its Implications. Creativity Research Journal 24, 2-3 (2012), 97–106. https://doi.org/10.1080/10400419.2012.676974 arXiv:https://doi.org/10.1080/10400419.2012.676974
- Singer et al. (2022) Uriel Singer, Adam Polyak, Thomas Hayes, Xiaoyue Yin, Jie An, Songyang Zhang, Qiyuan Hu, Harry Yang, Oron Ashual, Oran Gafni, Devi Parikh, Sonal Gupta, and Yaniv Taigman. 2022. Make-A-Video: Text-to-Video Generation without Text-Video Data. ArXiv abs/2209.14792 (2022). https://api.semanticscholar.org/CorpusID:252595919
- Skarlinski et al. (2024) Michael D. Skarlinski, Sam Cox, Jon M. Laurent, James D. Braza, Michaela Hinks, Michael J. Hammerling, Manvitha Ponnapati, Samuel G. Rodriques, and Andrew D. White. 2024. Language agents achieve superhuman synthesis of scientific knowledge. preprint (2024). https://paper.wikicrow.ai
- Sobel (2017) Benjamin Sobel. 2017. Artificial Intelligence’s Fair Use Crisis. Columbia Journal of Law and the Arts 41 (2017), 45–97. https://api.semanticscholar.org/CorpusID:115500744
- Somepalli et al. (2022) Gowthami Somepalli, Vasu Singla, Micah Goldblum, Jonas Geiping, and Tom Goldstein. 2022. Diffusion Art or Digital Forgery? Investigating Data Replication in Diffusion Models. 2023 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2022), 6048–6058. https://api.semanticscholar.org/CorpusID:254366634
- Sourati et al. (2023) Zhivar Sourati, Filip Ilievski, and Pia Sommerauer. 2023. ARN: Analogical Reasoning on Narratives. https://api.semanticscholar.org/CorpusID:263605669
- Stein (1953) Morris I Stein. 1953. Creativity and culture. The journal of psychology 36, 2 (1953), 311–322.
- Stevenson et al. (2022) Claire Stevenson, Iris Smal, Matthijs Baas, Raoul Grasman, and Han van der Maas. 2022. Putting GPT-3’s Creativity to the (Alternative Uses) Test. arXiv:2206.08932 [cs.AI]
- Stock and Strapparava (2005) Oliviero Stock and Carlo Strapparava. 2005. HAHAcronym: A Computational Humor System. In Proceedings of the ACL Interactive Poster and Demonstration Sessions, Masaaki Nagata and Ted Pedersen (Eds.). Association for Computational Linguistics, Ann Arbor, Michigan, 113–116. https://doi.org/10.3115/1225753.1225782
- Stowe et al. (2021) Kevin Stowe, Tuhin Chakrabarty, Nanyun Peng, Smaranda Muresan, and Iryna Gurevych. 2021. Metaphor Generation with Conceptual Mappings. ArXiv abs/2106.01228 (2021). https://api.semanticscholar.org/CorpusID:235294009
- Stowe et al. (2020) Kevin Stowe, Leonardo Ribeiro, and Iryna Gurevych. 2020. Metaphoric Paraphrase Generation. ArXiv abs/2002.12854 (2020). https://api.semanticscholar.org/CorpusID:211572789
- Sultan and Shahaf (2022) Oren Sultan and Dafna Shahaf. 2022. Life is a Circus and We are the Clowns: Automatically Finding Analogies between Situations and Processes. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, Yoav Goldberg, Zornitsa Kozareva, and Yue Zhang (Eds.). Association for Computational Linguistics, Abu Dhabi, United Arab Emirates, 3547–3562. https://doi.org/10.18653/v1/2022.emnlp-main.232
- Summers-Stay et al. (2023) Douglas Summers-Stay, Stephanie M. Lukin, and Clare R. Voss. 2023. Brainstorm, then Select: a Generative Language Model Improves Its Creativity Score. https://api.semanticscholar.org/CorpusID:259305709
- Sun et al. (2023) Jiao Sun, Yufei Tian, Wangchunshu Zhou, Nan Xu, Qian Hu, Rahul Gupta, John Wieting, Nanyun Peng, and Xuezhe Ma. 2023. Evaluating Large Language Models on Controlled Generation Tasks. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 3155–3168. https://doi.org/10.18653/v1/2023.emnlp-main.190
- Suzuki et al. (2018) Wendy A. Suzuki, Mónica I. Feliú-Mójer, Uri Hasson, Rachel Yehuda, and Jean Mary Zarate. 2018. Dialogues: The Science and Power of Storytelling. Journal of Neuroscience 38, 44 (2018), 9468–9470. https://doi.org/10.1523/JNEUROSCI.1942-18.2018 arXiv:https://www.jneurosci.org/content/38/44/9468.full.pdf
- Swanson et al. (2021) Benjamin Swanson, Kory Wallace Mathewson, Ben Pietrzak, Sherol Chen, and Monica Dinalescu. 2021. Story Centaur: Large Language Model Few Shot Learning as a Creative Writing Tool. In Conference of the European Chapter of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:233365238
- Sybrandt et al. (2020) Justin Sybrandt, Ilya Tyagin, Michael Shtutman, and Ilya Safro. 2020. AGATHA: Automatic Graph Mining And Transformer based Hypothesis Generation Approach. In Proceedings of the 29th ACM International Conference on Information & Knowledge Management (Virtual Event, Ireland) (CIKM ’20). Association for Computing Machinery, New York, NY, USA, 2757–2764. https://doi.org/10.1145/3340531.3412684
- Tam et al. (2022) Derek Tam, Anisha Mascarenhas, Shiyue Zhang, Sarah Kwan, Mohit Bansal, and Colin Raffel. 2022. Evaluating the Factual Consistency of Large Language Models Through News Summarization. In Annual Meeting of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:253523092
- Tambwekar et al. (2018) Pradyumna Tambwekar, Murtaza Dhuliawala, Lara J. Martin, Animesh Mehta, Brent Harrison, and Mark O. Riedl. 2018. Controllable Neural Story Plot Generation via Reward Shaping. In International Joint Conference on Artificial Intelligence. https://api.semanticscholar.org/CorpusID:199465680
- Tang et al. (2022) Leonard Tang, Alexander Cai, Steve Li, and Jason Wang. 2022. The Naughtyformer: A Transformer Understands Offensive Humor. arXiv:2211.14369 [cs.CL] https://arxiv.org/abs/2211.14369
- Taylor and Mazlack (2004) Julia M Taylor and Lawrence J Mazlack. 2004. Computationally recognizing wordplay in jokes. In Proceedings of the Annual Meeting of the Cognitive Science Society, Vol. 26.
- Thrush et al. (2022) Tristan Thrush, Ryan Jiang, Max Bartolo, Amanpreet Singh, Adina Williams, Douwe Kiela, and Candace Ross. 2022. Winoground: Probing Vision and Language Models for Visio-Linguistic Compositionality. 2022 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2022), 5228–5238. https://api.semanticscholar.org/CorpusID:248006414
- Tian et al. (2024) Yufei Tian, Tenghao Huang, Miri Liu, Derek Jiang, Alexander Spangher, Muhao Chen, Jonathan May, and Nanyun Peng. 2024. Are Large Language Models Capable of Generating Human-Level Narratives? ArXiv abs/2407.13248 (2024). https://api.semanticscholar.org/CorpusID:271270813
- Tian et al. (2021) Yufei Tian, Arvind krishna Sridhar, and Nanyun Peng. 2021. HypoGen: Hyperbole Generation with Commonsense and Counterfactual Knowledge. ArXiv abs/2109.05097 (2021). https://api.semanticscholar.org/CorpusID:237491428
- Tian and Peng (2022) Yufei Tian and Nanyun Peng. 2022. Zero-shot Sonnet Generation with Discourse-level Planning and Aesthetics Features. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Marine Carpuat, Marie-Catherine de Marneffe, and Ivan Vladimir Meza Ruiz (Eds.). Association for Computational Linguistics, Seattle, United States, 3587–3597. https://doi.org/10.18653/v1/2022.naacl-main.262
- Tian et al. (2023) Yufei Tian, Abhilasha Ravichander, Lianhui Qin, Ronan Joseph Le Bras, Raja Marjieh, Nanyun Peng, Yejin Choi, Thomas L. Griffiths, and Faeze Brahman. 2023. MacGyver: Are Large Language Models Creative Problem Solvers? ArXiv abs/2311.09682 (2023). https://api.semanticscholar.org/CorpusID:265221054
- Todd (1989) Peter M Todd. 1989. A connectionist approach to algorithmic composition. Computer Music Journal 13, 4 (1989), 27–43.
- Todorovski (1997) Ljup Co Todorovski. 1997. Declarative Bias in Equation Discovery. https://api.semanticscholar.org/CorpusID:7320694
- Torrance (1974) E.P. Torrance. 1974. Torrance Tests of Creative Thinking: Verbal Tests, Forms A and B, Figural Tests, Forms A and B. Xerox. https://books.google.ch/books?id=_4dUYAAACAAJ
- Trinh et al. (2024) Trieu H. Trinh, Yuhuai Wu, Quoc V. Le, He He, and Thang Luong. 2024. Solving olympiad geometry without human demonstrations. Nature 625, 7995 (01 Jan 2024), 476–482. https://doi.org/10.1038/s41586-023-06747-5
- Tulyakov et al. (2017) S. Tulyakov, Ming-Yu Liu, Xiaodong Yang, and Jan Kautz. 2017. MoCoGAN: Decomposing Motion and Content for Video Generation. 2018 IEEE/CVF Conference on Computer Vision and Pattern Recognition (2017), 1526–1535. https://api.semanticscholar.org/CorpusID:4475365
- Turing (1950) A. M. Turing. 1950. I.—COMPUTING MACHINERY AND INTELLIGENCE. Mind LIX, 236 (10 1950), 433–460. https://doi.org/10.1093/mind/LIX.236.433 arXiv:https://academic.oup.com/mind/article-pdf/LIX/236/433/30123314/lix-236-433.pdf
- Turner (1994) Scott R. Turner. 1994. The Creative Process: A Computer Model of Storytelling and Creativity. https://api.semanticscholar.org/CorpusID:260575817
- Turney (2008) Peter D. Turney. 2008. The Latent Relation Mapping Engine: Algorithm and Experiments. J. Artif. Intell. Res. 33 (2008), 615–655. https://api.semanticscholar.org/CorpusID:7112602
- Udrescu et al. (2020) S. M. Udrescu, Andrew Yong-Yi Tan, Jiahai Feng, Orisvaldo Neto, Tailin Wu, and Max Tegmark. 2020. AI Feynman 2.0: Pareto-optimal symbolic regression exploiting graph modularity. ArXiv abs/2006.10782 (2020). https://api.semanticscholar.org/CorpusID:219955930
- Ulam (1958) Stanisław Ulam. 1958. John von Neumann 1903-1957. Bull. Amer. Math. Soc. 64 (1958), 1–49. https://api.semanticscholar.org/CorpusID:122374268
- Ushio et al. (2021) Asahi Ushio, Luis Espinosa Anke, Steven Schockaert, and Jose Camacho-Collados. 2021. BERT is to NLP what AlexNet is to CV: Can Pre-Trained Language Models Identify Analogies?. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), Chengqing Zong, Fei Xia, Wenjie Li, and Roberto Navigli (Eds.). Association for Computational Linguistics, Online, 3609–3624. https://doi.org/10.18653/v1/2021.acl-long.280
- Van de Cruys (2020) Tim Van de Cruys. 2020. Automatic Poetry Generation from Prosaic Text. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, Dan Jurafsky, Joyce Chai, Natalie Schluter, and Joel Tetreault (Eds.). Association for Computational Linguistics, Online, 2471–2480. https://doi.org/10.18653/v1/2020.acl-main.223
- van den Oord et al. (2016b) Aäron van den Oord, Nal Kalchbrenner, Lasse Espeholt, Koray Kavukcuoglu, Oriol Vinyals, and Alex Graves. 2016b. Conditional Image Generation with PixelCNN Decoders. ArXiv abs/1606.05328 (2016). https://api.semanticscholar.org/CorpusID:14989939
- van den Oord et al. (2016a) Aäron van den Oord, Nal Kalchbrenner, and Koray Kavukcuoglu. 2016a. Pixel Recurrent Neural Networks. In International Conference on Machine Learning. https://api.semanticscholar.org/CorpusID:8142135
- Van Hee et al. (2018) Cynthia Van Hee, Els Lefever, and Véronique Hoste. 2018. SemEval-2018 Task 3: Irony Detection in English Tweets. In Proceedings of the 12th International Workshop on Semantic Evaluation, Marianna Apidianaki, Saif M. Mohammad, Jonathan May, Ekaterina Shutova, Steven Bethard, and Marine Carpuat (Eds.). Association for Computational Linguistics, New Orleans, Louisiana, 39–50. https://doi.org/10.18653/v1/S18-1005
- Vaswani et al. (2017) Ashish Vaswani, Noam M. Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N. Gomez, Lukasz Kaiser, and Illia Polosukhin. 2017. Attention is All you Need. In Neural Information Processing Systems. https://api.semanticscholar.org/CorpusID:13756489
- Veale (2015) Tony Veale. 2015. Game of Tropes: Exploring the Placebo Effect in Computational Creativity. In International Conference on Innovative Computing and Cloud Computing. https://api.semanticscholar.org/CorpusID:3966712
- Veale and Cardoso (2019) T. Veale and F.A. Cardoso. 2019. Computational Creativity: The Philosophy and Engineering of Autonomously Creative Systems. Springer. https://books.google.ch/books?id=EBY1ygEACAAJ
- Veale et al. (2016) T. Veale, E. Shutova, and B. Klebanov. 2016. Metaphor: A Computational Perspective. Morgan & Claypool Publishers. https://books.google.ch/books?id=_ripCwAAQBAJ
- Vecchi et al. (2017) Eva M. Vecchi, Marco Marelli, Roberto Zamparelli, and Marco Baroni. 2017. Spicy Adjectives and Nominal Donkeys: Capturing Semantic Deviance Using Compositionality in Distributional Spaces. Cognitive Science 41, 1 (2017), 102–136. https://doi.org/10.1111/cogs.12330 arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1111/cogs.12330
- Vinchon et al. (2023) Florent Vinchon, Todd Lubart, Sabrina Bartolotta, Valentin Gironnay, Marion Botella, Samira Bourgeois-Bougrine, Jean-Marie Burkhardt, Nathalie Bonnardel, Giovanni Emanuele Corazza, Vlad Petre Glăveanu, Michael Hanchett Hanson, Zorana Ivcevic, Maciej Karwowski, J. Kaufman, Takeshi Okada, Roni Reiter‐Palmon, and Andrea Gaggioli. 2023. Artificial Intelligence & Creativity: A Manifesto for Collaboration. The Journal of Creative Behavior (2023). https://api.semanticscholar.org/CorpusID:259480496
- Vondrick et al. (2016) Carl Vondrick, Hamed Pirsiavash, and Antonio Torralba. 2016. Generating Videos with Scene Dynamics. In Neural Information Processing Systems. https://api.semanticscholar.org/CorpusID:9933254
- Wallas (1926) G. Wallas. 1926. The Art of Thought. Solis Press. https://books.google.ch/books?id=JR6boAEACAAJ
- Waltz and Buchanan (2009) David Waltz and Bruce G. Buchanan. 2009. Automating Science. Science 324, 5923 (2009), 43–44. https://doi.org/10.1126/science.1172781 arXiv:https://www.science.org/doi/pdf/10.1126/science.1172781
- Wang et al. (2024b) Haonan Wang, James Zou, Michael Mozer, Anirudh Goyal, Alex Lamb, Linjun Zhang, Weijie J Su, Zhun Deng, Michael Qizhe Xie, Hannah Brown, and Kenji Kawaguchi. 2024b. Can AI Be as Creative as Humans? arXiv:2401.01623 [cs.AI] https://arxiv.org/abs/2401.01623
- Wang et al. (2023) Qingyun Wang, Doug Downey, Heng Ji, and Tom Hope. 2023. SciMON: Scientific Inspiration Machines Optimized for Novelty. In Annual Meeting of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:258841365
- Wang et al. (2024a) Qingyun Wang, Doug Downey, Heng Ji, and Tom Hope. 2024a. SciMON: Scientific Inspiration Machines Optimized for Novelty. arXiv:2305.14259 [cs.CL] https://arxiv.org/abs/2305.14259
- Wang et al. (2019) Qingyun Wang, Lifu Huang, Zhiying Jiang, Kevin Knight, Heng Ji, Mohit Bansal, and Yi Luan. 2019. PaperRobot: Incremental Draft Generation of Scientific Ideas. ArXiv abs/1905.07870 (2019). https://api.semanticscholar.org/CorpusID:159040684
- Webb et al. (2017) Margaret Webb, Daniel Little, and Simon Cropper. 2017. Once more with feeling: Normative data for the aha experience in insight and noninsight problems. Behavior Research Methods 50 (10 2017). https://doi.org/10.3758/s13428-017-0972-9
- Webb et al. (2022) Taylor W. Webb, Keith J. Holyoak, and Hongjing Lu. 2022. Emergent analogical reasoning in large language models. Nature Human Behaviour 7 (2022), 1526 – 1541. https://api.semanticscholar.org/CorpusID:254854575
- Wei et al. (2022) Jason Wei, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, Sebastian Borgeaud, Dani Yogatama, Maarten Bosma, Denny Zhou, Donald Metzler, Ed H. Chi, Tatsunori Hashimoto, Oriol Vinyals, Percy Liang, Jeff Dean, and William Fedus. 2022. Emergent Abilities of Large Language Models. arXiv:2206.07682 [cs.CL]
- Weisberg (1986) R.W. Weisberg. 1986. Creativity: Genius and Other Myths. W.H. Freeman. https://books.google.ch/books?id=xiCLQgAACAAJ
- Weisberg (2015a) Robert Weisberg. 2015a. On the Usefulness of “Value” in the Definition of Creativity. Creativity Research Journal 27 (06 2015), 111–124. https://doi.org/10.1080/10400419.2015.1030320
- Weisberg (2015b) Robert W. Weisberg. 2015b. On the Usefulness of “Value” in the Definition of Creativity. Creativity Research Journal 27, 2 (2015), 111–124. https://doi.org/10.1080/10400419.2015.1030320 arXiv:https://doi.org/10.1080/10400419.2015.1030320
- Weissweiler et al. (2023) Leonie Weissweiler, Valentin Hofmann, Anjali Kantharuban, Anna Cai, Ritam Dutt, Amey Hengle, Anubha Kabra, Atharva Kulkarni, Abhishek Vijayakumar, Haofei Yu, Hinrich Schuetze, Kemal Oflazer, and David Mortensen. 2023. Counting the Bugs in ChatGPT’s Wugs: A Multilingual Investigation into the Morphological Capabilities of a Large Language Model. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 6508–6524. https://doi.org/10.18653/v1/2023.emnlp-main.401
- Weller and Seppi (2019) Orion Weller and Kevin Seppi. 2019. Humor Detection: A Transformer Gets the Last Laugh. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), Kentaro Inui, Jing Jiang, Vincent Ng, and Xiaojun Wan (Eds.). Association for Computational Linguistics, Hong Kong, China, 3621–3625. https://doi.org/10.18653/v1/D19-1372
- Wijesiriwardene et al. (2023) Thilini Wijesiriwardene, Ruwan Wickramarachchi, Bimal Gajera, Shreeyash Mukul Gowaikar, Chandan Gupta, Aman Chadha, Aishwarya N. Reganti, Amit P. Sheth, and Amitava Das. 2023. ANALOGICAL - A Novel Benchmark for Long Text Analogy Evaluation in Large Language Models. ArXiv abs/2305.05050 (2023). https://api.semanticscholar.org/CorpusID:258947824
- Winters and Delobelle (2020) Thomas Winters and Pieter Delobelle. 2020. Dutch humor detection by generating negative examples. arXiv preprint arXiv:2010.13652 (2020).
- Wisniewski (1997) Edward J Wisniewski. 1997. When concepts combine. Psychonomic bulletin & review 4 (1997), 167–183.
- Wu and Tegmark (2019) Tailin Wu and Max Tegmark. 2019. Toward an artificial intelligence physicist for unsupervised learning. Physical Review E 100, 3 (2019), 033311.
- Xie et al. (2023b) Binzhu Xie, Sicheng Zhang, Zitang Zhou, Bo Li, Yuanhan Zhang, Jack Hessel, Jingkang Yang, and Ziwei Liu. 2023b. FunQA: Towards Surprising Video Comprehension. ArXiv abs/2306.14899 (2023). https://api.semanticscholar.org/CorpusID:259262297
- Xie et al. (2021) Yubo Xie, Junze Li, and Pearl Pu. 2021. Uncertainty and Surprisal Jointly Deliver the Punchline: Exploiting Incongruity-Based Features for Humor Recognition. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 2: Short Papers), Chengqing Zong, Fei Xia, Wenjie Li, and Roberto Navigli (Eds.). Association for Computational Linguistics, Online, 33–39. https://doi.org/10.18653/v1/2021.acl-short.6
- Xie et al. (2023a) Zhuohan Xie, Miao Li, Trevor Cohn, and Jey Lau. 2023a. DeltaScore: Fine-Grained Story Evaluation with Perturbations. In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.). Association for Computational Linguistics, Singapore, 5317–5331. https://doi.org/10.18653/v1/2023.findings-emnlp.353
- Xing et al. (2023) Zhen Xing, Qijun Feng, Haoran Chen, Qi Dai, Hang-Rui Hu, Hang Xu, Zuxuan Wu, and Yu-Gang Jiang. 2023. A Survey on Video Diffusion Models. ArXiv abs/2310.10647 (2023). https://api.semanticscholar.org/CorpusID:264172934
- Xu et al. (2022) Yudong Xu, Elias Boutros Khalil, and Scott Sanner. 2022. Graphs, Constraints, and Search for the Abstraction and Reasoning Corpus. In AAAI Conference on Artificial Intelligence. https://api.semanticscholar.org/CorpusID:252968055
- Xu et al. (2023) Yudong Xu, Wenhao Li, Pashootan Vaezipoor, Scott Sanner, and Elias Boutros Khalil. 2023. LLMs and the Abstraction and Reasoning Corpus: Successes, Failures, and the Importance of Object-based Representations. ArXiv abs/2305.18354 (2023). https://api.semanticscholar.org/CorpusID:258968016
- y Pérez and Sharples (2001) Rafael Pérez y Pérez and Mike Sharples. 2001. MEXICA: A computer model of a cognitive account of creative writing. Journal of Experimental & Theoretical Artificial Intelligence 13 (2001), 119 – 139. https://api.semanticscholar.org/CorpusID:18676334
- Yan et al. (2021) Wilson Yan, Yunzhi Zhang, P. Abbeel, and A. Srinivas. 2021. VideoGPT: Video Generation using VQ-VAE and Transformers. ArXiv abs/2104.10157 (2021). https://api.semanticscholar.org/CorpusID:233307257
- Yan et al. (2015) Xinchen Yan, Jimei Yang, Kihyuk Sohn, and Honglak Lee. 2015. Attribute2Image: Conditional Image Generation from Visual Attributes. In European Conference on Computer Vision. https://api.semanticscholar.org/CorpusID:7577075
- Yang et al. (2015) Diyi Yang, Alon Lavie, Chris Dyer, and Eduard H. Hovy. 2015. Humor Recognition and Humor Anchor Extraction. In Conference on Empirical Methods in Natural Language Processing. https://api.semanticscholar.org/CorpusID:11128248
- Yang et al. (2022) Kevin Yang, Yuandong Tian, Nanyun Peng, and Dan Klein. 2022. Re3: Generating Longer Stories With Recursive Reprompting and Revision. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, Yoav Goldberg, Zornitsa Kozareva, and Yue Zhang (Eds.). Association for Computational Linguistics, Abu Dhabi, United Arab Emirates, 4393–4479. https://doi.org/10.18653/v1/2022.emnlp-main.296
- Yang et al. (2017) Li-Chia Yang, Szu-Yu Chou, and Yi-Hsuan Yang. 2017. MidiNet: A Convolutional Generative Adversarial Network for Symbolic-Domain Music Generation. ArXiv abs/1703.10847 (2017). https://api.semanticscholar.org/CorpusID:2002865
- Yang and Lerch (2018) Li-Chia Yang and Alexander Lerch. 2018. On the evaluation of generative models in music. Neural Computing and Applications 32 (2018), 4773 – 4784. https://api.semanticscholar.org/CorpusID:53258271
- Yang et al. (2023) Zonglin Yang, Xinya Du, Junxian Li, Jie Zheng, Soujanya Poria, and E. Cambria. 2023. Large Language Models for Automated Open-domain Scientific Hypotheses Discovery. In Annual Meeting of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:261557055
- Yao et al. (2019) Lili Yao, Nanyun Peng, Weischedel Ralph, Kevin Knight, Dongyan Zhao, and Rui Yan. 2019. Plan-And-Write: Towards Better Automatic Storytelling. In The Thirty-Third AAAI Conference on Artificial Intelligence (AAAI-19).
- Yasunaga et al. (2023) Michihiro Yasunaga, Xinyun Chen, Yujia Li, Panupong Pasupat, Jure Leskovec, Percy Liang, Ed Huai hsin Chi, and Denny Zhou. 2023. Large Language Models as Analogical Reasoners. ArXiv abs/2310.01714 (2023). https://api.semanticscholar.org/CorpusID:263608847
- Yin et al. (2021) Zong Yin, Federico Reuben, Susan Stepney, and Tom Collins. 2021. ”A Good Algorithm Does Not Steal - It Imitates”: The Originality Report as a Means of Measuring When a Music Generation Algorithm Copies Too Much. In EvoMUSART. https://api.semanticscholar.org/CorpusID:233236232
- Yu et al. (2016) Lantao Yu, Weinan Zhang, Jun Wang, and Yong Yu. 2016. SeqGAN: Sequence Generative Adversarial Nets with Policy Gradient. In AAAI Conference on Artificial Intelligence. https://api.semanticscholar.org/CorpusID:3439214
- Yu et al. (2018) Zhiwei Yu, Jiwei Tan, and Xiaojun Wan. 2018. A Neural Approach to Pun Generation. In Annual Meeting of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:51874992
- Yuan et al. (2022) Ann Yuan, Andy Coenen, Emily Reif, and Daphne Ippolito. 2022. Wordcraft: Story Writing With Large Language Models. Proceedings of the 27th International Conference on Intelligent User Interfaces (2022). https://api.semanticscholar.org/CorpusID:247585187
- Yuan et al. (2024) Ruibin Yuan, Hanfeng Lin, Yi Wang, Zeyue Tian, Shangda Wu, Tianhao Shen, Ge Zhang, Yuhang Wu, Cong Liu, Ziya Zhou, et al. 2024. Chatmusician: Understanding and generating music intrinsically with llm. arXiv preprint arXiv:2402.16153 (2024).
- Yuan et al. (2023) Siyu Yuan, Jiangjie Chen, Xuyang Ge, Yanghua Xiao, and Deqing Yang. 2023. Beneath Surface Similarity: Large Language Models Make Reasonable Scientific Analogies after Structure Abduction. In Conference on Empirical Methods in Natural Language Processing. https://api.semanticscholar.org/CorpusID:258833317
- Zambaldi et al. (2024) Vinicius Zambaldi, David La, Alexander E. Chu, Harshnira Patani, Amy E. Danson, Tristan O. C. Kwan, Thomas Frerix, Rosalia G. Schneider, David Saxton, Ashok Thillaisundaram, Zachary Wu, Isabel Moraes, Oskar Lange, Eliseo Papa, Gabriella Stanton, Victor Martin, Sukhdeep Singh, Lai H. Wong, Russ Bates, Simon A. Kohl, Josh Abramson, Andrew W. Senior, Yilmaz Alguel, Mary Y. Wu, Irene M. Aspalter, Katie Bentley, David L. V. Bauer, Peter Cherepanov, Demis Hassabis, Pushmeet Kohli, Rob Fergus, and Jue Wang. 2024. De novo design of high-affinity protein binders with AlphaProteo. arXiv:2409.08022 [q-bio.BM] https://arxiv.org/abs/2409.08022
- Zarei et al. (2024) Arman Zarei, Keivan Rezaei, Samyadeep Basu, Mehrdad Saberi, Mazda Moayeri, Priyatham Kattakinda, and Soheil Feizi. 2024. Understanding and Mitigating Compositional Issues in Text-to-Image Generative Models. ArXiv abs/2406.07844 (2024). https://api.semanticscholar.org/CorpusID:270391565
- Zawada (2006) Britta Zawada. 2006. Linguistic creativity from a cognitive perspective. Southern African Linguistics and Applied Language Studies 24, 2 (2006), 235–254. https://doi.org/10.2989/16073610609486419 arXiv:https://doi.org/10.2989/16073610609486419
- Zhang et al. (2019a) Chi Zhang, Feng Gao, Baoxiong Jia, Yixin Zhu, and Song-Chun Zhu. 2019a. RAVEN: A Dataset for Relational and Analogical Visual REasoNing. 2019 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2019), 5312–5322. https://api.semanticscholar.org/CorpusID:71148268
- Zhang et al. (2021) Chi Zhang, Baoxiong Jia, Mark Edmonds, Song-Chun Zhu, and Yixin Zhu. 2021. ACRE: Abstract Causal REasoning Beyond Covariation. 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2021), 10638–10648. https://api.semanticscholar.org/CorpusID:232380074
- Zhang et al. (2019b) Dongyu Zhang, Heting Zhang, Xikai Liu, Hongfei Lin, and Feng Xia. 2019b. Telling the Whole Story: A Manually Annotated Chinese Dataset for the Analysis of Humor in Jokes. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), Kentaro Inui, Jing Jiang, Vincent Ng, and Xiaojun Wan (Eds.). Association for Computational Linguistics, Hong Kong, China, 6402–6407. https://doi.org/10.18653/v1/D19-1673
- Zhang et al. (2016b) Han Zhang, Tao Xu, Hongsheng Li, Shaoting Zhang, Xiaogang Wang, Xiaolei Huang, and Dimitris N. Metaxas. 2016b. StackGAN: Text to Photo-Realistic Image Synthesis with Stacked Generative Adversarial Networks. 2017 IEEE International Conference on Computer Vision (ICCV) (2016), 5908–5916. https://api.semanticscholar.org/CorpusID:1277217
- Zhang et al. (2023) Muru Zhang, Ofir Press, William Merrill, Alisa Liu, and Noah A. Smith. 2023. How Language Model Hallucinations Can Snowball. ArXiv abs/2305.13534 (2023). https://api.semanticscholar.org/CorpusID:258841857
- Zhang et al. (2016a) Richard Zhang, Phillip Isola, and Alexei A. Efros. 2016a. Colorful Image Colorization. In European Conference on Computer Vision. https://api.semanticscholar.org/CorpusID:50698
- Zhang and Liu (2014) Renxian Zhang and Naishi Liu. 2014. Recognizing Humor on Twitter. In Proceedings of the 23rd ACM International Conference on Conference on Information and Knowledge Management (Shanghai, China) (CIKM ’14). Association for Computing Machinery, New York, NY, USA, 889–898. https://doi.org/10.1145/2661829.2661997
- Zhang and Lapata (2014) Xingxing Zhang and Mirella Lapata. 2014. Chinese Poetry Generation with Recurrent Neural Networks. In Conference on Empirical Methods in Natural Language Processing. https://api.semanticscholar.org/CorpusID:12964363
- Zhao et al. (2023a) Wenting Zhao, Justin T Chiu, Jena D. Hwang, Faeze Brahman, Jack Hessel, Sanjiban Choudhury, Yejin Choi, Xiang Lorraine Li, and Alane Suhr. 2023a. UNcommonsense Reasoning: Abductive Reasoning about Uncommon Situations. ArXiv abs/2311.08469 (2023). https://api.semanticscholar.org/CorpusID:265213087
- Zhao et al. (2023b) Wei Zhao, Michael Strube, and Steffen Eger. 2023b. DiscoScore: Evaluating Text Generation with BERT and Discourse Coherence. In Proceedings of the 17th Conference of the European Chapter of the Association for Computational Linguistics, Andreas Vlachos and Isabelle Augenstein (Eds.). Association for Computational Linguistics, Dubrovnik, Croatia, 3865–3883. https://doi.org/10.18653/v1/2023.eacl-main.278
- Zhao et al. (2023c) Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, Yifan Du, Chen Yang, Yushuo Chen, Zhipeng Chen, Jinhao Jiang, Ruiyang Ren, Yifan Li, Xinyu Tang, Zikang Liu, Peiyu Liu, Jian-Yun Nie, and Ji-Rong Wen. 2023c. A Survey of Large Language Models. arXiv:2303.18223 [cs.CL]
- Zhao et al. (2022) Xuandong Zhao, Lei Li, and Yu-Xiang Wang. 2022. Provably Confidential Language Modelling. arXiv:2205.01863 [cs.CL] https://arxiv.org/abs/2205.01863
- Zhao et al. (2024) Yunpu Zhao, Rui Zhang, Wenyi Li, Di Huang, Jiaming Guo, Shaohui Peng, Yifan Hao, Yuanbo Wen, Xingui Hu, Zidong Du, Qi Guo, Ling Li, and Yunji Chen. 2024. Assessing and Understanding Creativity in Large Language Models. ArXiv abs/2401.12491 (2024). https://api.semanticscholar.org/CorpusID:267094860
- Zhong et al. (2023) Shan Zhong, Zhongzhan Huang, Shanghua Gao, Wushao Wen, Liang Lin, Marinka Zitnik, and Pan Zhou. 2023. Let’s Think Outside the Box: Exploring Leap-of-Thought in Large Language Models with Creative Humor Generation. 2024 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) (2023), 13246–13257. https://api.semanticscholar.org/CorpusID:265659076
- Zhu and de Melo (2020) Xunjie Zhu and Gerard de Melo. 2020. Sentence Analogies: Linguistic Regularities in Sentence Embeddings. In Proceedings of the 28th International Conference on Computational Linguistics, Donia Scott, Nuria Bel, and Chengqing Zong (Eds.). International Committee on Computational Linguistics, Barcelona, Spain (Online), 3389–3400. https://doi.org/10.18653/v1/2020.coling-main.300
- Ziser et al. (2020) Yftah Ziser, Elad Kravi, and David Carmel. 2020. Humor Detection in Product Question Answering Systems. In Proceedings of the 43rd International ACM SIGIR Conference on Research and Development in Information Retrieval (Virtual Event, China) (SIGIR ’20). Association for Computing Machinery, New York, NY, USA, 519–528. https://doi.org/10.1145/3397271.3401077

### v3

@@
 Hence, it is not entirely clear whether the seemingly extraordinary outputs of these models are the result of a truly creative inner process and robust generalization or the result of powerful interpolation and strong memorization skills (McCoy et al., 2023; Hupkes et al., 2022; Carlini et al., 2022; Bender et al., 2021; Marcus, 2020; Bender and Koller, 2020).
-Recent works have shown that language models fail at real-world commonsense reasoning and compositionality tasks (Ismayilzada et al., 2023; Dziri et al., 2023), occasionally copy large amounts of text from their training data (McCoy et al., 2021), significantly lag behind experts in creative writing (Chakrabarty et al., 2023a), produce less diverse content (Anderson et al., 2024a; Padmakumar and He, 2023), struggle with creative problem-solving (Tian et al., 2023) and abstract reasoning (Gendron et al., 2023; Mitchell et al., 2023) and suffer from factual inconsistency and hallucination issues (Banerjee et al., 2024; Elazar et al., 2021).
+Recent works have shown that language models fail at real-world commonsense reasoning and compositionality tasks (Ismayilzada et al., 2023; Dziri et al., 2023), occasionally copy large amounts of text from their training data (Lu et al., 2024b; McCoy et al., 2021), significantly lag behind humans in creative writing (Ismayilzada et al., 2024b; Chakrabarty et al., 2023a), produce less diverse content (Anderson et al., 2024a; Padmakumar and He, 2023), struggle with creative problem-solving (Huang et al., 2024b; Tian et al., 2023) and abstract reasoning (Gendron et al., 2023; Mitchell et al., 2023) and suffer from factual inconsistency and hallucination issues (Banerjee et al., 2024; Elazar et al., 2021).
 While previous works have reviewed developments on some aspects of AI creativity (Amin and Burghardt, 2020; Lai and Nissim, 2024; Oliveira, 2017; Elzohbi and Zhao, 2023; Franceschelli and Musolesi, 2021b; Rowe and Partridge, 1993), a more holistic and broader review of the field is necessary to understand its rapid advancements.
 In this work, we provide the general AI audience with a timely summary of the state of the creative capabilities of the latest AI systems.
-While creativity is a broad concept that can be explored in a wide range of areas, in this survey, we focus on four main areas where machine creativity has been extensively investigated: Linguistic Creativity (§3), Creative Problem-Solving (§4), Artistic Creativity (§5) and Scientific Creativity (§6).
+While creativity is a broad concept that can be explored in a wide range of areas, in this survey, we focus on four main areas where machine creativity has been extensively investigated: Linguistic Creativity (§3.1), Creative Problem-Solving (§3.2), Artistic Creativity (§3.3) and Scientific Creativity (§3.4).
 These areas capture four crucial pillars of creative thinking in humans: linguistic creativity enables us to manipulate language in novel ways for effective communication of ideas; creative problem-solving helps us find efficient solutions by thinking out-of-the-box; artistic creativity allows for the expression of emotions, ideas, and aesthetics through various media; and scientific creativity drives innovation by enabling the formulation of new hypotheses, theories, and discoveries.
@@
 2. Creativity
-2.1. Definitions
+2.1. Defining Creativity
 While creativity as a concept seems intuitively easy to understand on the surface, there is still no consensus on what constitutes true creativity.
@@
 The novelty criterion is typically self-explanatory to the point that people equate it to creativity in everyday life.
-This is generally known as intrapersonal or personal creativity (often denoted as P-creativity), i.e.
-the product is novel within the frame of a person’s life (Boden, 2004; Stein, 1953; Weisberg, 1986).
-Researchers distinguish it from the interpersonal or historical creativity (often denoted as H-creativity), i.e.
-the product is novel with respect to the entire history of people such as Einstein’s general relativity theory.
-Many theorists have argued that novelty is insufficient for creativity, and value dimension is needed to filter out original nonsense, such as something generated by a truly random process.
+However, many theorists have argued that novelty is insufficient for creativity, and value dimension is needed to filter out original nonsense, such as something generated by a truly random process.
 While value is generally understood as something inherently “good” for the respective audience, there appears to be such a thing as malevolent or “dark” creativity.
@@
 Moreover, sometimes novelty is itself the value created by the artist because no one has done it before, particularly, in visual arts.
-Hence, some researchers have recently argued to drop the value criterion altogether from the definition of creativity (Brandt, 2021; Weisberg, 2015b).
-Despite its wide adoption, the sufficiency of the novelty and value conditions for creativity has also been challenged (Gaut, 2010; Arnheim, 2001; Weisberg, 2015a).
+Hence, some researchers have recently argued to drop the value criterion altogether from the definition of creativity (Brandt, 2021; Weisberg, 2015a).
+Despite its wide adoption, the sufficiency of the novelty and value conditions for creativity has also been challenged (Gaut, 2010; Arnheim, 2001; Weisberg, 2015b).
 It has been argued that agency, a capacity to have beliefs, desires or intentional states, is a required attribute of creativity.
@@
 Gaut (2010) illustrates this with an example where a person walking in a studio accidentally knocks over a set of paints, which spill onto a canvas and happen to create a beautiful and original painting.
-Weisberg (2015b) has gone even further to suggest that creativity is simply intentional novelty.
+Weisberg (2015a) has gone even further to suggest that creativity is simply intentional novelty.
 While the creative process should be intentionally initiated, others have argued that the creative process should involve an element of spontaneity (Kronfeldner, 2009).
@@
 In this survey, we will also take this extended definition as our working definition throughout the paper.
-2.2. Types
-While creativity manifests itself in various forms across domains, even within a particular domain, different types of creativity can be distinguished based on the difficulty level of the inherent creative process involved.
-Four-C model of creativity is one such popular theory that differentiates between four types of creativity corresponding to four levels of difficulty involved in producing creative artifacts (Kaufman and Beghetto, 2009).
+2.2. Types of Creativity
+While creativity manifests itself in various forms across domains, even within a particular domain, different types of creativity can be distinguished based on its timing or target audience and the difficulty level of the inherent creative process involved.
+A person might come up with a creative idea that is new to him/her but already invented by someone else in history.
+This is generally known as intrapersonal or personal (a.k.a psychological) creativity (often denoted as P-creativity), i.e.
+the product is novel within the frame of a person’s life (Boden, 2004; Stein, 1953; Weisberg, 1986).
+Researchers distinguish it from the interpersonal or historical creativity (often denoted as H-creativity), i.e.
+the product is novel with respect to the entire history of people such as Einstein’s general relativity theory.
+Four-C model of creativity, on the other hand, differentiates between four types of creativity corresponding to four levels of difficulty involved in producing creative artifacts (Kaufman and Beghetto, 2009).
 The first major type of creativity is known as little-c creativity which is what we find in everyday life as solutions to minor problems.
@@
 Inventing chess itself or any major scientific invention is a perfect example of this type of creativity.
+This type of creativity typically requires transformation of the existing conceptual space and is also known as transformational creativity (Boden, 2004).
 2.3. Evaluation
@@
 For example, in story generation, humans are typically asked to rate a generated story on aspects such as interestingness, coherence, relevance, humanlikeness and etc.
-(Goldfarb-Tarrant et al., 2020; Rashkin et al., 2020; Yang et al., 2022).
+(Yang and Jin, 2024; Goldfarb-Tarrant et al., 2020; Rashkin et al., 2020; Yang et al., 2022).
 In other tasks where the goal is to produce multiple responses such as common psychometric creativity tests Alternative Uses Task (AUT) (Guilford, 1967) and Torrance Tests of Creative Thinking (TTCT) (Torrance, 1974), evaluation is centered around four dimensions of creativity: fluency (the total number of meaningful, and relevant ideas generated in response to the stimulus), flexibility (the number of different categories of relevant responses), originality (the uniqueness or rarity of responses) and elaboration (the amount of detail in the responses).
@@
 For example, in open-ended story generation, a minimally useful story can be defined as a relevant, coherent, and meaningful story.
-In this sense, automated metrics measuring the overall quality of a story can be leveraged (Xie et al., 2023a; Guan and Huang, 2020; Chen et al., 2022b), however, it is often challenging to measure coherence (Zhao et al., 2023b; Laban et al., 2021).
+In this sense, automated metrics measuring the overall quality of a story can be leveraged (Xie et al., 2023a, b; Guan and Huang, 2020; Chen et al., 2022b), however, it is often challenging to measure coherence (Zhao et al., 2023b; Laban et al., 2021).
 Another more general evaluation of utility has been suggested by Franceschelli and Musolesi (2022) based on the discriminator score in GANs.
+Since in GANs, the discriminator learns the distribution of the real (and valuable) data, its score can directly be used as a proxy metric to measure value.
 Surprise
@@
 Linguistic
-Creativity (§3)
-Humor (§3.1)
+Creativity (§3.1)
+Humor (§3.1.1)
 Datasets
@@
 Methods
-Humorous content (Ravi et al., 2024; Hessel et al., 2023b; Xie et al., 2023b; Arora et al., 2022; Peyrard et al., 2021; Kayatani et al., 2021),
+Humorous content (Ravi et al., 2024; Hessel et al., 2023b; Xie et al., 2023c; Arora et al., 2022; Peyrard et al., 2021; Kayatani et al., 2021),
 (Annamoradnejad and Zoghi, 2020; Ziser et al., 2020; Amin and Burghardt, 2020; Chen and Soo, 2018; Liu et al., 2018b; Bertero and Fung, 2016),
@@
 Figurative
-Language (§3.2)
+Language (§3.1.2)
 Datasets
@@
 Lexical Innovation
-(§3.3)
+(§3.1.3)
 Novel compound (Coil and Shwartz, 2023; Dhar and van der Plas, 2019; Vecchi et al., 2017; Kuznetsova et al., 2013; Hendrickx et al., 2013; Butnariu et al., 2009),
 Neologism (Lencione et al., 2022; Das and Ghosh, 2017),
-Morphological Productivity (Ismayilzada et al., 2024; Anh et al., 2024; Weissweiler et al., 2023)
+Morphological Productivity (Ismayilzada et al., 2024a; Anh et al., 2024; Weissweiler et al., 2023)
 , lc-leaf, text width=48em
@@
 Problem
-Solving (§4)
+Solving (§3.2)
 Convergent
-Thinking (§4.1)
+Thinking (§3.2.1)
 NYT Connections (Samadarshi et al., 2024),
@@
 Divergent
-Thinking (§4.2)
+Thinking (§3.2.2)
 BrainTeaser (Jiang et al., 2023),
@@
 Artistic
-Creativity (§5)
+Creativity (§3.3)
 Story
-Generation (§5.1)
+Generation (§3.3.1)
 Datasets
@@
 , ac-leaf, text width=41em
-Poetry (§5.2)
+Poetry (§3.3.2)
 Datasets
@@
 Visual Creativity
-(§5.3)
+(§3.3.3)
 Image Editing
@@
 Musical Creativity
-(§5.4)
+(§3.3.4)
 MusicGen (Copet et al., 2023), MusicLM (Agostinelli et al., 2023), Jukebox (Dhariwal et al., 2020),
@@
 Scientific
-Creativity (§6)
+Creativity (§3.4)
 Equation discovery
@@
 , sc-leaf, text width=50em
-3. Linguistic Creativity
+3. Domains of Creativity
+Creativity is a multifaceted concept that spans across various domains, each harnessing its unique form of imaginative thought and innovation.
+In this section, we will review the state of creativity in AI across four major domains where machine creativity is most extensively explored: linguistics, art, science, and problem-solving.
+3.1. Linguistic Creativity
 The creative aspect of language in linguistics has been discussed since the early days (Chomsky, 1965).
@@
 In this sense, Sampson (2017) suggests distinguishing between F-creativity (fixed) and E-creativity (extending), where F-creativity refers to the Chomskian interpretation of linguistic creativity (a.k.a productivity in morphology) and E-creativity corresponds to the real linguistic innovation such as metaphors, jokes, neologisms, etc.
-Some recent works have explored the F-creativity of large language models and found that this task is challenging in general and even harder in more morphological complex languages (Ismayilzada et al., 2024; Anh et al., 2024; Weissweiler et al., 2023).
+Some recent works have explored the F-creativity of large language models and found that this task is challenging in general and even harder in more morphological complex languages (Ismayilzada et al., 2024a; Anh et al., 2024; Weissweiler et al., 2023).
 Most past works however have focused on studying the E-creativity of AI systems which we review in the following sections.
-3.1. Humor
-Humor is one of the most common ways humans creatively use language to express their ideas and feelings.
+3.1.1. Humor
+Humor is one of the most common ways in which humans creatively use language to express their ideas and feelings.
 Early works to model humor focused on hand-crafted linguistic templates and wordplay (Stock and Strapparava, 2005; Taylor and Mazlack, 2004; Raskin and Attardo, 1994).
@@
 Several datasets have also been proposed to benchmark the humor capacity of LLMs in several languages including English (Horvitz et al., 2024; Jain et al., 2024; Tang et al., 2022; Meaney et al., 2021; Hossain et al., 2020b, 2019; Miller et al., 2017), Chinese (Zhang et al., 2019b), Italian (Buscaldi and Rosso, 2007), Spanish (Castro et al., 2017), Dutch (Winters and Delobelle, 2020) and Russian (Blinov et al., 2019).
-Computational humor has also been explored in multimodal settings involving images, audio, and video in addition to text (Hessel et al., 2023a; Xie et al., 2023b; Hasan et al., 2019; Bertero and Fung, 2016; Radev et al., 2016; Shahaf et al., 2015).
+Computational humor has also been explored in multimodal settings involving images, audio, and video in addition to text (Hessel et al., 2023a; Xie et al., 2023c; Hasan et al., 2019; Bertero and Fung, 2016; Radev et al., 2016; Shahaf et al., 2015).
 While the latest methods particularly LLMs show an impressive ability to generate and detect humorous content, recent work has also shown that these models still fail to reliably understand humor (Hessel et al., 2023a; Góes et al., 2023a; Kocoń et al., 2023; Borji, 2023a) and generated jokes typically lack diversity (Jentzsch and Kersting, 2023) which has been attributed to training on less diverse humor datasets (Baranov et al., 2023).
@@
 We refer the reader to Amin and Burghardt (2020) for an in-depth survey on computational humor.
-3.2. Figurative Language
+3.1.2. Figurative Language
 Figurative language is a term in language studies encompassing various figures of speech like hyperbole, similes and metaphor (Veale et al., 2016; Roberts and Kreuz, 1994; Paul, 1970).
@@
 (2020) for an in-depth survey on the automatic generation and detection of figurative language.
-3.3. Lexical Innovation
+3.1.3. Lexical Innovation
 Understanding and generating novel words or word compounds is a challenging linguistic task that often requires creativity, commonsense knowledge, and an ability to generalize over seen concepts (Costello and Keane, 2000; Wisniewski, 1997).
 Similar noun compounds might have different meanings based on our common understanding.
-For example, knowing that “chocolate croissant” means a “croissant filled with chocolate” does not necessarily imply that “chocolate bunny” would mean “a bunny filled with chocolate”, but rather a bar of “chocolate in the shape of a bunny”.
+For example, knowing that “chocolate croissant” means a “croissant filled with chocolate” does not necessarily imply that “chocolate bunny” would mean “a bunny filled with chocolate”, but rather a piece of “chocolate in the shape of a bunny”.
 Several works have evaluated and analyzed language models on the task of interpreting and predicting the emergence of these noun compounds and found that models generally show a moderate performance (Coil and Shwartz, 2023; Dhar and van der Plas, 2019; Kuznetsova et al., 2013).
@@
 newly coined words or phrases) (Das and Ghosh, 2017; Lencione et al., 2022).
-On the other hand, previous works have also shown that large language models can fail at linguistic generalization tasks such as morphologically deriving new words from nonce roots (Ismayilzada et al., 2024; Weissweiler et al., 2023) and can occasionally duplicate large amounts of text from its training data (McCoy et al., 2021).
+On the other hand, previous works have also shown that large language models can fail at linguistic generalization tasks such as morphologically deriving new words from nonce roots (Ismayilzada et al., 2024a; Weissweiler et al., 2023) and can occasionally duplicate large amounts of text from its training data (McCoy et al., 2021).
 Similarly, a recent work explores the linguistic creativity of both large language models and humans by reconstructing their text output from the existing text snippets on the web and finds that the seemingly remarkable creativity of model outputs may be in large part attributable to the remarkable creativity of human-written texts on the web (Lu et al., 2024b).
-4. Creative Problem-Solving
+3.2. Creative Problem-Solving
 Creative problem-solving is the mental process of searching and coming up with creative solutions to a given problem (Duncker and Lees, 1948).
@@
 In addition, creatively solving a problem is usually characterized by two kinds of thinking, namely, convergent and divergent thinking, and involves deep abstraction and analogy-making abilities.
-4.1. Convergent Thinking
+3.2.1. Convergent Thinking
 Convergent thinking models creativity in terms of an ability to produce a single optimal solution for a given problem (Guilford, 1967).
@@
 misleading stimuli or distractors) and evaluates large language models such as GPT-4 on these problems.
-They report poor model performance and show that models are highly susceptible to distractors in the input and manifest a form of fixation effect (a.k.a functional fixedness or Einstellung effect).
+They report poor model performance and show that models are highly susceptible to distractors in the input and manifest a form of fixation effect (a.k.a functional fixedness or Einstellung effect) (Wiley, 1998; Smith and Blankenship, 1991; Barber, 1960).
 This type of cognitive bias forces the model to fixate on its past knowledge and prevents it from thinking “out-of-the-box”.
@@
 Very recently, large language models such as GPT-4o have been evaluated on the popular New York Times game Connections and have been found to struggle with associating encyclopedic and linguistic knowledge at an abstract level (Samadarshi et al., 2024).
-Another study has revealed that language models also fall short of demonstrating human-like creativity in code generation (Lu et al., 2024c).
-4.2. Divergent Thinking
+Another study investigating both convergent and divergent creativity of language models has revealed that language models also fall short of demonstrating human-like convergent creativity in code generation (Lu et al., 2024c).
+3.2.2. Divergent Thinking
 Divergent thinking requires one to conceptualize multiple, often seemingly disconnected ideas (Guilford, 1967).
@@
 Chen and Ding (2023) uses this test to evaluate the creativity of large language models and finds that models outperform humans.
-While language models perform strongly on AUT-like divergent thinking tasks, they, however, struggle when these tasks require some form of lateral thinking or “thinking out-of-the-box”.
+While language models perform strongly on AUT-like divergent thinking tasks, they, however, struggle when these tasks require some form of lateral thinking or “thinking out-of-the-box” (Huang et al., 2024b).
 For example, recent works have found that defying default commonsense associations and modeling unexpected or unlikely situations are challenging for large language models (Jiang et al., 2023; Tian et al., 2023; Zhao et al., 2023a).
@@
 (2023) that involves unconventional use of everyday objects.
-4.3. Abstraction and Analogy-Making
+3.2.3. Abstraction and Analogy-Making
 Conceptual abstraction and analogy-making lie at the core of human cognition and intelligence (Chollet, 2019; Mitchell, 2021; Hofstadter, 2001).
@@
 In other words, abstraction and analogy-making serve as foundational building blocks for creative thinking.
-4.3.1. Abstraction
+Abstraction
 As the cornerstone of human intelligence, abstraction, and abstract reasoning are typically evaluated using visual IQ tests in humans.
-Popular examples of these tests are RAVEN progressive matrices (Raven, 1938), Bongard problems (Bongard, 1970) and recently introduced Kandinsky Patterns (Holzinger et al., 2019), Abstraction and Reasoning Corpus (ARC) (Chollet, 2019) and its variations (Moskvichev et al., 2023).
+Popular examples of these tests are the RAVEN progressive matrices (Raven, 1938), Bongard problems (Bongard, 1970) and the recently introduced Kandinsky Patterns (Holzinger et al., 2019), the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019) and its variations (Moskvichev et al., 2023).
 These tests require the participants to identify and complete an abstract visual pattern based on given examples.
@@
 The problems in this corpus are quite hard to solve to the extent that this task has been recognized as the de facto benchmark for measuring progress towards Artificial General Intelligence (AGI) and a public competition with a grand prize of $ has recently been launched222https://arcprize.org/.
-At the time of writing this paper, the highest score is far from from the passing threshold of (human-level).
-4.3.2. Analogy-Making
+At the time of writing this paper, the highest score is far from the passing threshold of (human-level).
+Analogy-Making
 In its basic form, analogy-making is the ability to identify a relation between two concepts and apply it to a new concept.
@@
 structured representations of both the entities and their relations (Turney, 2008; Gentner, 1983; Falkenhainer et al., 1989).
-Later, word embedding models based on neural networks were shown to exhibit analogy-making abilities at the word level (Jacob et al., 2023; Petersen and van der Plas, 2023; Marquer et al., 2022; Ushio et al., 2021; Gladkova et al., 2016; Mikolov et al., 2013).
-However, word-level analogies are generally simple and do not reflect the typical analogical reasoning humans perform in everyday life about complex situations.
+Later, word embedding models based on neural networks were shown to exhibit analogy-making abilities at the word level and most works focused on a limited set of analogy types based on a handful of relations that are often of a morphological nature.
+(Marquer et al., 2022; Gladkova et al., 2016; Mikolov et al., 2013).
+These do not encompass the typical analogical reasoning humans perform in everyday life about complex situations.
+More recent work has focused on including a multitude of relations and datasets that are used to test analogical reasoning in humans.
+(Petersen and van der Plas, 2023; Ushio et al., 2021; Jacob et al., 2023)
 While some works have argued for emergent analogical reasoning abilities of large language models (Webb et al., 2022; Yasunaga et al., 2023; Hu et al., 2023), other works have shown that these models lack the robustness and generality exhibited by humans when it comes to long text analogies (Wijesiriwardene et al., 2023; Zhu and de Melo, 2020), scientific analogies (Yuan et al., 2023; Czinczoll et al., 2022), story analogies (Jiayang et al., 2023; Sourati et al., 2023; Sultan and Shahaf, 2022; Nagarajah et al., 2022), visual analogies (Opielka et al., 2024; Bitton et al., 2022; Zhang et al., 2019a) and complex analogical reasoning (Musker et al., 2024; Lewis and Mitchell, 2024).
 Figure 4(b) illustrates an example from the analogical reasoning over narratives benchmark (Sourati et al., 2023).
-5. Artistic Creativity
+3.3. Artistic Creativity
 Artistic creativity is the ability to produce original, imaginative, and expressive works in various art forms, such as creative writing, poetry, visual arts, music, dance, theater, and more.
 In this section, we will focus on the advancements made in AI to produce creative stories, poetry, visual, and musical content automatically and also point out the remaining challenges.
-5.1. Story Generation
+3.3.1. Story Generation
 Storytelling is at the heart of human communication, a powerful tool for connecting and conveying ideas effectively (Suzuki et al., 2018).
@@
 In addition, long-term factual inconsistency and hallucinations still remain as major issues in language model generated texts (Banerjee et al., 2024; Elazar et al., 2021; Tam et al., 2022; Zhang et al., 2023).
-Language models have also been evaluated on their ability to produce and judge creative content as a professional writer (Chakrabarty et al., 2023b; Marco et al., 2024).
+Language models have also been evaluated on their ability to produce and judge creative content as a professional writer (Chakrabarty et al., 2023b; Marco et al., 2024; Gómez-Rodríguez and Williams, 2023).
 Chakrabarty et al.
@@
 Additionally, (Tian et al., 2024) finds that LLM-generated stories are positively homogenous and typically lack suspense and tension.
+LLMs have also been shown to produce more complex, but less creative stories than average humans (Ismayilzada et al., 2024b).
 To complement the shortcomings of LLMs in creative content generation, recently several works have developed frameworks to use these models as creative assistants for humans and these collaborative systems have shown strong performance across domains and editing tasks (Yuan et al., 2022; Schick et al., 2022; Mirowski et al., 2022; Chakrabarty et al., 2023c; Swanson et al., 2021).
@@
 Particularly, adapting language models with human feedback (Ouyang et al., 2022) has been found to be a main contributing factor in diversity reduction (Mohammadi, 2024; Padmakumar and He, 2023; Bai et al., 2022).
-5.2. Poetry
+3.3.2. Poetry
 Poetry is a form of literary expression that uses rhythmic and often condensed creative language to evoke emotions, convey ideas, or tell stories.
@@
 We refer the reader to (Oliveira, 2017; Elzohbi and Zhao, 2023) for an in-depth survey on automatic poetry generation.
-5.3. Visual Creativity
+3.3.3. Visual Creativity
 Humans have been producing visual content to convey emotions, concepts, and narratives since ancient times, from cave paintings and hieroglyphics to classical and Renaissance art masterpieces.
@@
 Similarly, video generation models often suffer from a lack of reliable spatial reasoning, appearance inconsistency, temporal inalignment, body deformation and occlusion issues (Lei et al., 2024; Brooks et al., 2024).
-5.4. Musical Creativity
+3.3.4. Musical Creativity
 Music is another major artistic medium that allows individuals to express emotions, ideas, and cultural narratives through sound, often transcending language barriers to connect people across diverse backgrounds and experiences.
@@
 Recent studies also show that humans exhibit a preference for human compositions over AI compositions and they report something “off” about the latter such as a lack of sense of coherence or consistency, odd note choices, unnecessary complexity, repetition, uninterestingness, and failure to come to a resolution (Sarmento et al., 2024).
-6. Scientific Creativity
+3.4. Scientific Creativity
 Scientific creativity refers to the ability to generate novel ideas, approaches, or solutions within the realm of science, often leading to new discoveries, theories, or technologies.
@@
 Recent methods, however, often employ Bayesian statistics (Guimerà et al., 2020) and neural networks (Chen et al., 2022a; Garcon et al., 2021; Udrescu et al., 2020; Cranmer et al., 2020; Petersen and Landajuela, 2019).
-Another line of work has focused on automating the discovery of other scientific knowledge such as generating new mathematical conjectures or theories (Raayoni et al., 2021; Wu and Tegmark, 2019; Chen et al., 2016; Buchberger et al., 2006; Fajtlowicz, 1988), automatically proving theorems (Hubert et al., 2024b; Trinh et al., 2024), discovering new concepts (Hakuk and Reich, 2020; Iten et al., 2020; Lenat and Brown, 1984) and predicting molecular structures (Abramson et al., 2024; Zambaldi et al., 2024; Jumper et al., 2021; Lindsay, 1980) among others.
+Another line of work has focused on automating the discovery of other scientific knowledge such as generating new mathematical conjectures or theories (Raayoni et al., 2021; Wu and Tegmark, 2019; Chen et al., 2016; Buchberger et al., 2006; Fajtlowicz, 1988), automatically proving theorems (Hubert et al., 2024b; Trinh et al., 2024), discovering new concepts (Hakuk and Reich, 2020; Iten et al., 2020; Lenat and Brown, 1984) and predicting new molecular structures (Abramson et al., 2024; Zambaldi et al., 2024; Jumper et al., 2021; Lindsay, 1980) among others.
 A notable example in this area is the recent AlphaFold model (Jumper et al., 2021) that can predict millions of intricate 3D protein structures which has the potential to significantly accelerate research in biology.
@@
 Their qualitative analysis also reveals some common failure modes such as vague implementation details, misuse of datasets, inappropriate baselines, unrealistic assumptions, and overall poorly-motivated ideas.
-Similarly, another study benchmarking the machine learning experimentation capabilities of LLMs reports hallucinations and poor planning as some of the major issues with these models (Huang et al., 2024b).
+Similarly, another study benchmarking the machine learning experimentation capabilities of LLMs reports hallucinations and poor planning as some of the major issues with these models (Huang et al., 2024c).
 Another important aspect of scientific discovery is the explainability (Li et al., 2021) which helps humans prevent or better prepare for a possible future technological singularity (Good, 1965; Ulam, 1958).
@@
 Humans typically communicate an emotion or a deeper meaning through creative products and their creative process is characterized by spontaneous “Aha” or “Eureka” moments coupled with deliberate decisions made at each step of the way 666https://www.newyorker.com/culture/the-weekend-essay/why-ai-isnt-going-to-make-art.
-However, current AI systems lack agency and are typically trained to generate the most likely output leaving no room for any intentional or spontaneous action (Franceschelli and Musolesi, 2023).
+However, current AI systems lack agency and are typically trained to generate the most likely output leaving no room for any intentional or spontaneous action (Peeperkorn et al., 2023; Franceschelli and Musolesi, 2023).
 Therefore, a holistic evaluation of machine creativity should involve consideration of all these different dimensions that characterize human creativity.
@@
 Acknowledgements.
-We gratefully acknowledge the support of Swiss National Science Foundation (grant 205121_207437: C - LING).References
+We gratefully acknowledge the support of Swiss National Science Foundation (grant 205121_207437: C - LING). We also thank Angelika Romanou for her help with creating figures for this paper.References
 - (1)
@@
 https://api.semanticscholar.org/CorpusID:232417054
-- Arnheim (2001) Rudolf Arnheim. 2001. What it Means to be Creative. British Journal of Aesthetics 41 (01 2001). https://doi.org/10.1093/bjaesthetics/41.1.24
+- Arnheim (2001) Rudolf Arnheim. 2001. What it means to be creative. British Journal of Aesthetics 41 (2001), 24–25. https://api.semanticscholar.org/CorpusID:191484364
 - Arora et al.
@@
 https://doi.org/10.18653/v1/2023.emnlp-main.845
-- Barron (1955) Frank Barron. 1955. The disposition toward originality. The Journal of Abnormal and Social Psychology 51, 3 (1955), 478.
+- Barber (1960) Theodore Xenophon Barber.
+1960.
+Rigidity of Behavior.
+A Variational Approach to the Effect of Einstellung.Abraham S.
+Luchins , Edith Hirsch Luchins.
+The Quarterly Review of Biology 35 (1960), 255–255.
+https://api.semanticscholar.org/CorpusID:85022531
+- Barron (1955) Frank Barron. 1955. The disposition toward originality. Journal of abnormal psychology 51 3 (1955), 478–85. https://api.semanticscholar.org/CorpusID:30102436
 - Beaty and Johnson (2020) Roger E.
@@
 - Elgammal et al.
-(2017) Ahmed Elgammal, Bingchen Liu, Mohamed Elhoseiny, and Marian Mazzone.
+(2017) A.
+Elgammal, Bingchen Liu, Mohamed Elhoseiny, and Marian Mazzone.
 2017.
 CAN: Creative Adversarial Networks, Generating ”Art” by Learning About Styles and Deviating from Style Norms.
-arXiv:1706.07068 [cs.AI]
+In International Conference on Innovative Computing and Cloud Computing.
+https://api.semanticscholar.org/CorpusID:24986117
 - Elgammal and Saleh (2015) A. Elgammal and Babak Saleh. 2015. Quantifying Creativity in Art Networks. ArXiv abs/1506.00711 (2015). https://api.semanticscholar.org/CorpusID:14607648
@@
 4319–4338.
+- Gómez-Rodríguez and Williams (2023) Carlos Gómez-Rodríguez and Paul Williams.
+2023.
+A Confederacy of Models: a Comprehensive Evaluation of LLMs on Creative Writing.
+In Findings of the Association for Computational Linguistics: EMNLP 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.).
+Association for Computational Linguistics, Singapore, 14504–14528.
+https://doi.org/10.18653/v1/2023.findings-emnlp.966
 - Good (1965) I. J. Good. 1965. Speculations Concerning the First Ultraintelligent Machine. Adv. Comput. 6 (1965), 31–88. https://api.semanticscholar.org/CorpusID:17886872
@@
 - Huang et al.
-(2024b) Qian Huang, Jian Vora, Percy Liang, and Jure Leskovec.
-2024b.
+(2024c) Qian Huang, Jian Vora, Percy Liang, and Jure Leskovec.
+2024c.
 MLAgentBench: Evaluating Language Agents on Machine Learning Experimentation.
 arXiv:2310.03302 [cs.LG] https://arxiv.org/abs/2310.03302
+- Huang et al.
+(2024b) Shulin Huang, Shirong Ma, Yinghui Li, Mengzuo Huang, Wuhe Zou, Weidong Zhang, and Haitao Zheng.
+2024b.
+LatEval: An Interactive LLMs Evaluation Benchmark with Incomplete Information from Lateral Thinking Puzzles.
+In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), Nicoletta Calzolari, Min-Yen Kan, Veronique Hoste, Alessandro Lenci, Sakriani Sakti, and Nianwen Xue (Eds.).
+ELRA and ICCL, Torino, Italia, 10186–10197.
+https://aclanthology.org/2024.lrec-main.889
 - Huang et al.
@@
 - Ismayilzada et al.
-(2024) Mete Ismayilzada, Defne Circi, Jonne Sälevä, Hale Sirin, Abdullatif Köksal, Bhuwan Dhingra, Antoine Bosselut, Lonneke van der Plas, and Duygu Ataman.
-2024.
+(2024a) Mete Ismayilzada, Defne Circi, Jonne Sälevä, Hale Sirin, Abdullatif Köksal, Bhuwan Dhingra, Antoine Bosselut, Lonneke van der Plas, and Duygu Ataman.
+2024a.
 Evaluating Morphological Compositional Generalization in Large Language Models.
@@
 https://api.semanticscholar.org/CorpusID:264438931
+- Ismayilzada et al.
+(2024b) Mete Ismayilzada, Claire Stevenson, and Lonneke van der Plas.
+2024b.
+Evaluating Creative Short Story Generation in Humans and Large Language Models.
+arXiv:2411.02316 [cs.CL] https://arxiv.org/abs/2411.02316
 - Isola et al.
@@
 - Pease et al. (2002) Alison Pease, Daniel Winterstein, and Simon Colton. 2002. Evaluating Machine Creativity. (10 2002).
+- Peeperkorn et al. (2023) Max Peeperkorn, Dan Brown, and Anna Jordanous. 2023. On characterizations of large language models and creativity evaluation. (2023).
 - Peeperkorn et al.
@@
 https://paper.wikicrow.ai
+- Smith and Blankenship (1991) Steven M.
+Smith and Steven E.
+Blankenship.
+1991.
+Incubation and the persistence of fixation in problem solving.
+The American journal of psychology 104 1 (1991), 61–87.
+https://api.semanticscholar.org/CorpusID:10359632
 - Sobel (2017) Benjamin Sobel. 2017. Artificial Intelligence’s Fair Use Crisis. Columbia Journal of Law and the Arts 41 (2017), 45–97. https://api.semanticscholar.org/CorpusID:115500744
@@
 - Sourati et al. (2023) Zhivar Sourati, Filip Ilievski, and Pia Sommerauer. 2023. ARN: Analogical Reasoning on Narratives. https://api.semanticscholar.org/CorpusID:263605669
-- Stein (1953) Morris I Stein. 1953. Creativity and culture. The journal of psychology 36, 2 (1953), 311–322.
+- Stein (1953) Morris Isaac Stein. 1953. Creativity and Culture. Creativity in Art, Religion, and Culture (1953). https://api.semanticscholar.org/CorpusID:144504562
 - Stevenson et al. (2022) Claire Stevenson, Iris Smal, Matthijs Baas, Raoul Grasman, and Han van der Maas. 2022. Putting GPT-3’s Creativity to the (Alternative Uses) Test. arXiv:2206.08932 [cs.AI]
@@
 - Weisberg (1986) R.W. Weisberg. 1986. Creativity: Genius and Other Myths. W.H. Freeman. https://books.google.ch/books?id=xiCLQgAACAAJ
-- Weisberg (2015a) Robert Weisberg.
+- Weisberg (2015a) Robert W.
+Weisberg.
 2015a.
 On the Usefulness of “Value” in the Definition of Creativity.
-Creativity Research Journal 27 (06 2015), 111–124.
-https://doi.org/10.1080/10400419.2015.1030320
+Creativity Research Journal 27, 2 (2015), 111–124.
+https://doi.org/10.1080/10400419.2015.1030320 arXiv:https://doi.org/10.1080/10400419.2015.1030320
 - Weisberg (2015b) Robert W.
@@
 On the Usefulness of “Value” in the Definition of Creativity.
-Creativity Research Journal 27, 2 (2015), 111–124.
-https://doi.org/10.1080/10400419.2015.1030320 arXiv:https://doi.org/10.1080/10400419.2015.1030320
+Creativity Research Journal 27 (2015), 111 – 124.
+https://api.semanticscholar.org/CorpusID:146831140
 - Weissweiler et al.
@@
 https://api.semanticscholar.org/CorpusID:258947824
+- Wiley (1998) Jennifer Wiley.
+1998.
+Expertise as mental set: The effects of domain knowledge in creative problem solving.
+Memory & Cognition 26 (1998), 716–730.
+https://api.semanticscholar.org/CorpusID:45775509
 - Winters and Delobelle (2020) Thomas Winters and Pieter Delobelle. 2020. Dutch humor detection by generating negative examples. arXiv preprint arXiv:2010.13652 (2020).
@@
 - Xie et al.
-(2023b) Binzhu Xie, Sicheng Zhang, Zitang Zhou, Bo Li, Yuanhan Zhang, Jack Hessel, Jingkang Yang, and Ziwei Liu.
-2023b.
+(2023c) Binzhu Xie, Sicheng Zhang, Zitang Zhou, Bo Li, Yuanhan Zhang, Jack Hessel, Jingkang Yang, and Ziwei Liu.
+2023c.
 FunQA: Towards Surprising Video Comprehension.
@@
 - Xie et al.
-(2023a) Zhuohan Xie, Miao Li, Trevor Cohn, and Jey Lau.
+(2023a) Zhuohan Xie, Trevor Cohn, and Jey Han Lau.
 2023a.
+The Next Chapter: A Study of Large Language Models in Storytelling.
+In Proceedings of the 16th International Natural Language Generation Conference, C.
+Maria Keet, Hung-Yi Lee, and Sina Zarrieß (Eds.).
+Association for Computational Linguistics, Prague, Czechia, 323–351.
+https://doi.org/10.18653/v1/2023.inlg-main.23
+- Xie et al.
+(2023b) Zhuohan Xie, Miao Li, Trevor Cohn, and Jey Lau.
+2023b.
 DeltaScore: Fine-Grained Story Evaluation with Perturbations.
@@
 https://api.semanticscholar.org/CorpusID:7577075
+- Yang and Jin (2024) Dingyi Yang and Qin Jin. 2024. What makes a good story and how can we measure it? a comprehensive survey of story evaluation. arXiv preprint arXiv:2408.14622 (2024).
 - Yang et al.